A Complete Bibliography of Computer Animation and Virtual Worlds

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 April 2022 Version 1.27

Title word cross-reference

+ [563]. 2 [777, 752, 159, 303, 246, 720, 282]. 3 [622, 113, 406, 291, 672, 819, 837, 908, 742, 349, 174, 201, 721, 65, 356, 491, 232, 255, 293, 549, 752, 535, 149, 220, 393, 475, 192, 826, 806, 530, 51, 134, 298, 328, 412, 992, 159, 757, 277, 525, 625, 366, 453, 401, 854, 759, 396, 830, 831, 308, 221, 124, 664, 22, 860, 142, 610, 765, 920, 867, 766, 395]. ^{∘} [925]. d [699]. k [760, 699]. l_{1} [398]. m [143]. N [273, 143].

-D [742]. -nearest [760]. -sparse [398]. -to- [143]. -way [273].

0 [102]. 0-471-36089-9 [102].

14 [314].

2 [314]. 2006 [175]. 25.5-6 [601]. 2D [273]. 2nd [102].

3-axis [327]. 3D [598, 986].

9 [102].

AA [558]. AA-FVDM [558]. ablation [54]. absorption [523]. abstracted [520]. abstraction [116, 133]. accelerated [506, 403]. Accelerating [745]. acceleration [22]. accelerator [960]. accelerometers [274]. accessibility [444]. accessing [713]. accident [558]. accident-avoidance [558]. ACCLMesh [684]. account [50]. accuracy [667]. Accurate [806, 51, 532]. achieving [327]. ACM [175]. acoustic [79]. acquisition [392]. across [604]. action [902, 915, 752, 578]. actions [137]. active [145]. activity [906]. adaptation [214, 203, 227]. adapting [734]. Adaptive [644, 593, 59, 320, 367, 470, 428, 209, 371, 797, 385, 621, 634, 300, 397, 799, 930, 69, 609, 865, 305, 643, 934, 965]. Adding [235]. additive [961]. ADF [731]. adjustment [300, 324, 787]. Advanced [629, 930]. advances [162]. Advected [281]. advection [621, 186, 912, 934]. advection-reaction-diffusion [186]. adversarial [842, 991, 993]. advice [598]. AFF [967]. AFF-Dehazing [967]. affect [338]. Affection [35]. Affective [467, 834, 529, 27, 950]. Africa [944]. against [485, 663]. Age [747]. Age-related [747]. Agent [234, 852, 263, 657, 839, 359, 987, 851, 239, 342, 437, 510, 861]. Agent-based [234, 342]. agents [547, 907, 598, 208, 214, 126, 741, 467, 32, 145, 307, 4, 2, 674, 633, 5, 527, 850, 597, 526, 163, 360, 727, 780, 583, 342, 337, 284, 740]. AGRMTS [990]. aided [167, 154]. aircraft [990]. airways [38]. ALET [740]. algorithm [605, 577, 223, 298, 447, 865, 400, 900]. algorithms [351, 42]. alignment [135, 765]. all-in-one [587]. ambient [237, 614]. amidst [658]. among [680]. Analysis [3, 217, 820, 122, 347, 516, 421, 719, 334, 772, 168]. analytic [936]. Analytical [219, 611]. Analyzing [984]. Anatomical [433, 269]. anatomy [82, 983]. ancient [55, 154]. And-Or [450]. aneurysm [694]. angle [285]. angle-split [285]. angles [602]. Angular [302]. animatable [349, 203]. animated [960, 438, 721, 638]. Animating [89, 6, 73, 335, 933, 178, 828, 446, 437, 912, 4, 325, 292, 623, 558, 585, 509, 84, 913]. Animation [333, 117, 472, 642, 310, 295, 354, 273, 445, 118, 835, 120, 515, 836, 948, 226, 709, 424, 358, 70, 10, 464, 392, 671, 915, 552, 3, 479, 416, 92, 15, 232, 749, 779, 225, 116, 789, 848, 99, 459, 475, 153, 192, 524, 382, 179, 231, 228, 596, 898, 48, 312, 813, 277, 447, 256, 200, 334, 37, 197, 74, 321, 148, 456, 242, 409, 124, 584, 609, 613, 72, 687, 315, 238, 425, 567, 595, 783, 636, 193, 452, 196, 190, 514, 764, 539, 81, 499]. animations [838, 86, 356, 491, 618, 454, 957]. Anime [303]. anisotropy [186]. annotation [421, 578]. ant [799, 866]. anterior [497]. anthropometrics [36]. Anticipated [658]. anticipation [388]. Anticipatory [504, 811]. ants [743]. anxious [361]. app [925]. appearance [321]. Application [781, 504, 540, 977]. Applications [218, 9, 45, 108, 431, 460, 714, 326, 804]. applied [786, 433]. approach [805, 402, 43, 623, 845, 512, 749, 825, 975, 61, 164, 347, 971, 588, 577, 134, 439, 330, 80, 112, 591, 575, 499]. approaches [498, 620]. Approximate [680]. Approximating [188]. approximation [472, 859]. arbitrary [186, 51, 602]. architectural [625]. architecture [141, 116, 290, 313]. arrangement [955]. art [498, 845, 629, 554, 664, 63]. arthropod [499]. arthroscopic [497]. articulated [406, 230, 447]. articulated-character [447]. articulation [3]. artificial [32]. artisanal [409]. artistic [618]. artists [64]. artworks [878]. As-consistent-As-possible [131]. as-rigid-as-possible [777]. Asch [803]. ASFNet [965]. aspects [839]. assembly [685, 862]. Assessing [714]. assessment [268]. assistance [977]. Assistant [598, 436]. assisted [641, 289]. Asymmetric [547, 978]. asynchronous [680, 401]. attachment [526]. Attention [967, 276, 415, 360, 337, 940]. Attention-based [967]. attributes [437]. Audio [835, 30, 662]. Audio-driven [835]. audio-visual [30]. Augmented [214, 56, 310, 723, 432, 978, 152, 325, 378, 347, 216, 54, 970, 360, 213, 891, 977, 44, 365, 804]. authoring [776, 838, 618, 789, 861]. auto [94]. autoencoder [904]. automata [33]. Automated [77, 821, 41, 631, 715]. Automatic [907, 349, 174, 277, 324, 759, 130, 81, 395, 552, 301, 698, 51, 955, 447, 873, 315, 539]. automatically [37]. Autonomous [121, 762, 208, 33, 32, 145, 415, 597, 191]. autonomy [307, 314]. avatar [634, 753, 813, 570, 718]. avatar-mediated [753]. avatars [742, 349, 757, 23, 395]. avoidance [650, 558, 511]. aware [789, 265, 620, 854, 720, 964, 282, 923, 457, 452, 241]. awareness [258, 215]. axis [327].

B [918, 917]. B-spline [918, 917]. background [152, 408, 928]. balance [89, 811]. ball [819]. ball-and-socket [819]. Bar [196]. Bar-net [196]. Bas [955]. Bas-relief [955]. based [56, 805, 960, 141, 548, 472, 98, 354, 644, 176, 684, 942, 818, 195, 26, 406, 120, 363, 405, 724, 902, 938, 963, 33, 686, 529, 706, 736, 887, 961, 675, 797, 464, 19, 838, 97, 477, 725, 479, 185, 986, 201, 703, 422, 90, 842, 956, 183, 521, 369, 746, 43, 289, 531, 796, 871, 88, 747, 654, 549, 517, 73, 658, 975, 116, 194, 752, 61, 220, 265, 614, 754, 954, 82, 378, 287, 629, 403, 755, 941, 266, 104, 322, 886, 179, 347, 473, 756, 872, 677, 234, 528, 566, 227, 51, 712, 223, 247, 312, 286, 360, 812, 827, 397, 640, 638, 447, 25, 80]. based [324, 511, 853, 540, 112, 780, 772, 781, 758, 342, 168, 483, 791, 401, 74, 370, 321, 372, 437, 510, 576, 541, 76, 759, 787, 377, 958, 24, 246, 428, 694, 808, 613, 795, 68, 649, 865, 891, 922, 344, 567, 46, 990, 859, 940, 923, 455, 345, 523, 78, 343, 383, 542, 182, 142, 564, 667, 764, 920, 919, 929, 294, 407, 867, 900, 665, 967, 993, 140, 365, 766, 81, 571, 630, 357, 440]. Basis [619]. Bayesian [424]. be [710]. beam [195, 292]. Behavior [34, 631, 805, 504, 691, 820, 261, 599, 927, 234, 864, 511, 603, 715]. Behavioral [33, 310, 10, 750]. behaviors [371, 897, 586, 423, 276, 585, 413, 316, 541, 861]. Behaviour [104]. Behaviour-based [104]. behaviours [669]. behind [615]. Believable [166, 733, 541]. benchmark [316]. bending [776, 253]. between [208, 987, 753, 382, 23, 582, 773, 461, 243]. big [709]. big-object [709]. bilinear [765]. billiards [366]. BIM [971]. binary [140]. binomial [937]. Biologically [866]. biomechanical [703, 747]. biomechanics [188]. blend [948]. blend-trees [948]. blending [477, 721, 341, 580, 13, 11, 667]. blendshape [458]. blendshapes [231]. Blob [88]. Blob-based [88]. blood [641, 525]. blooming [651]. Board [441]. bodies [790, 230, 225, 628]. body [9, 675, 416, 606, 219, 933, 806, 286, 812, 420, 696, 976, 396]. body-fixed [9]. boids [121]. Boltzmann [140]. Book [102, 920]. books [432]. boosting [291]. boundaries [58]. Boundary [651, 964]. boundary-aware [964]. Boundary-dominant [651]. bounded [658, 540]. bounding [484, 701]. box [850]. BPNN [990]. braid [952]. braided [952]. brain [821, 823]. branches [292]. breast [255, 168]. Bringing [119]. brittle [515, 870, 386, 463]. browser [685]. brush [469]. brush-pressure [469]. Brush2Model [766]. brushes [242, 766]. Brushwork [913]. bubble [652, 808]. bubbles [483]. buckling [253]. budget [906]. Building [968, 971, 867, 625, 173]. buildup [459]. Bulging [581]. Bulging-free [581]. Burdea [102].

C [102]. CAD [688]. cage [624, 455]. cages [593]. calibrating [173]. Call [218, 561]. calligraphic [469]. camera [723, 532, 309, 378, 695, 925, 324, 854, 396, 400]. cantilever [292]. canvas [676]. canvases [325]. capture [879, 262, 118, 189, 375, 956, 637, 752, 231, 577, 51, 312, 112, 546, 816, 570, 74, 76, 398, 937, 39]. capture-based [112]. captured [445, 459, 901]. Capturing [9]. car [733]. cardiovascular [575]. Caricature [181]. cartoon [92, 232, 335, 382, 148, 279, 425, 209, 193, 243, 343, 241]. carving [469]. Cascaded [858]. casting [265, 555]. casual [898]. casually [901]. catching [630]. Catheterization [683]. Catmull [621]. caused [808]. cave [873]. CAVW [435]. CCA [549]. CCA-based [549]. cells [105]. cells-and-portals [105]. chalk [664]. Chameleon [541]. changes [306]. changing [444, 553, 587]. Character [719, 388, 226, 686, 424, 358, 949, 10, 552, 491, 654, 892, 959, 847, 296, 153, 290, 228, 227, 188, 447, 200, 865, 315, 567, 130, 762, 196, 514]. Characteristic [394, 248]. Characteristics [456]. characterization [27]. characterized [632]. Characters [968, 77, 233, 548, 438, 835, 777, 33, 495, 599, 415, 277, 781, 332, 582, 773, 720, 702, 687, 129, 429, 632, 490]. checking [582]. cheer [936]. Chemical [289]. Chi [26]. Chinese [153, 629, 554, 913, 288]. Choreographing [336, 274]. chroma [304]. cities [768]. city [364, 216]. city/region [364]. cityscapes [62]. classical [825]. Classification [529, 985]. classroom [977]. client [663]. client-server [663]. clinical [9, 944]. clips [241]. clocks [154]. cloning [25, 80, 148]. close [837]. close-up [837]. cloth [67, 460, 252, 592, 791, 69, 636, 253]. cloth-like [67]. clothes [531]. cloud [644, 919, 929]. Cloudy [945, 692]. clumping [779]. cluster [171]. clustering [506, 157]. CNC [327]. CNN [840, 923]. co [3]. co-articulation [3]. coarse [339]. CODE [778]. cognition [698]. cognitive [371, 10, 423, 290]. cohabitant [35]. Coherence [265, 185, 454, 750, 217]. coherent [454, 849, 236]. Coiffet [102]. collaboration [57, 159]. collaborations [680]. Collaborative [597, 104]. collective [743]. colliding [543]. collision [471, 299, 67, 650, 692, 892, 223, 505, 511, 582, 823, 151, 974, 68, 172]. collisions [773]. colonoscopy [265]. colony [799, 866]. color [344]. coloring [94]. colorization [991]. Combined [308]. Combining [48, 667, 379, 310, 202]. combustion [387]. Comic [424]. commodity [806, 570]. communication [572, 215]. communicative [135, 597]. comparative [258, 760]. comparing [439]. comparison [691]. Compatible [275]. Compensating [934]. Competitive [287]. complete [701]. completion [240]. complex [226, 708, 633, 58, 971, 647]. complexity [298]. component [719]. compositing [704, 131]. composition [820, 573, 453]. comprehensive [834]. Compressed [875]. Compression [530, 113, 721, 134, 298, 29, 308, 407]. computation [823]. Computational [771, 845, 825, 290, 540, 758]. computed [38]. Computer [167, 218, 926, 154, 423, 885, 163, 541, 337, 694]. Computer-aided [154]. computer-based [694]. concatenation [51, 302]. condensation [930]. condensed [22]. Conditional [358, 842]. conference [634]. confidence [324]. confinement [385, 934]. conflicts [680]. conformity [803]. connected [849, 916]. connectivity [431, 807, 157]. considering [206]. consistent [761, 131]. constrained [836, 887, 251, 330, 679, 238]. Constraint [247, 363, 620, 631]. constraint-aware [620]. Constraint-based [247]. constraints [670, 788, 416, 225, 848, 534]. construction [349, 701]. Constructive [749]. contact [691]. content [942]. contents [713]. context [201, 563, 633, 698, 437, 631]. context-dependent [631]. Continuous [696, 471, 172]. Contour [614, 41]. Contour-based [614]. control [741, 709, 309, 788, 986, 634, 219, 815, 150, 287, 770, 503, 328, 75, 719, 112, 811, 781, 758, 321, 409, 209, 44, 142, 610, 643, 798, 294, 630]. control-space [112]. Controllable [86, 653, 595, 387]. controlled [536, 777]. controller [890, 813, 283]. controllers [548]. Controlling [15]. conversational [547, 907, 126, 31, 4, 2, 5, 702]. conversations [437]. conversers [466]. conversion [602]. Convolution [766, 517, 919, 991]. convolutional [839]. cooperative [503]. coordinate [765, 919]. coordinates [422, 453]. coordinating [702]. coordination [936, 807]. core [443, 688]. correct [256]. corrected [295]. correlates [911]. Correlative [334]. correspondence [899]. corridor [165]. Corruptible [674]. cosmetic [36, 521]. Cosserat [195]. Cosserat-beam-based [195]. cost [925]. counselor [748]. Coupling [513, 861, 790, 619, 641, 731, 535, 780, 592, 461, 591]. Cover [882, 841, 843, 800, 893, 792, 894]. CPU [512]. craniofacial [403]. create [356]. creation [654, 652]. creature [129]. critical [70, 299, 254, 306]. Cross [413, 97, 904]. cross-modal [904]. cross-parameterization [97]. Cross-platform [413]. Crowd [189, 778, 468, 984, 117, 645, 657, 506, 744, 465, 698, 905, 927, 769, 669, 872, 898, 234, 852, 827, 133, 511, 583, 510, 435, 760, 646, 979, 440]. Crowd-optimized [778]. crowd-sourcing [698]. crowded [84]. crowds [805, 492, 938, 276, 439, 509, 144, 390, 787, 436]. cruciate [497]. crystallization [746]. CSLML [288]. CT [724]. cubic [86, 564]. cues [137, 324]. culling [471, 217, 171]. cultural [839, 43, 878]. current [498]. curvature [684]. curvature-based [684]. curvatures [293]. Curve [129]. curved [220]. curves [248, 917]. curvilinear [58]. cut [192, 712, 412, 867]. cut-out [412]. cutting [484, 565, 928, 640, 22, 591]. cyberspace [62]. cyberworld [116]. cycle [37].

D [535, 159, 303, 854, 622, 723, 113, 406, 291, 672, 777, 819, 837, 908, 742, 349, 174, 201, 721, 65, 356, 491, 232, 255, 293, 549, 752, 149, 220, 393, 475, 192, 826, 806, 530, 51, 134, 298, 328, 412, 992, 159, 757, 277, 525, 625, 366, 453, 401, 759, 396, 830, 831, 308, 221, 124, 246, 720, 664, 22, 282, 860, 142, 610, 765, 920, 867, 903, 766, 395, 571]. D- [159]. D-RGB [752]. damage [692]. damping [69]. Dance [697, 985, 399, 936, 136, 301, 153, 72]. dancing [77, 399]. Data [853, 687, 707, 879, 262, 869, 724, 706, 8, 745, 956, 668, 280, 654, 563, 459, 231, 588, 821, 577, 51, 312, 578, 313, 456, 221, 896, 937, 653, 209, 190]. Data-driven [853, 687, 707, 654, 588, 456, 896, 190]. database [969, 870]. datasets [266]. DBSC [120]. DBSC-based [120]. Dealing [306]. debris [609]. decision [467, 526, 254, 306]. decision-making [467, 526, 254, 306]. decomposition [556, 99, 624]. decoration [561]. decoupled [503]. dedicated [423]. deducing [566]. Deep [904, 964, 948, 910, 852, 858, 940, 991, 903, 957]. Deferred [95]. defined [786]. definition [334]. Deformable [583, 790, 108, 59, 70, 139, 299, 476, 484, 731, 247, 16, 518, 679, 41, 974, 68, 172, 331, 628]. Deformation [164, 300, 881, 262, 777, 788, 71, 376, 138, 183, 280, 293, 430, 650, 789, 147, 296, 82, 565, 348, 525, 592, 823, 246, 720, 130, 591, 457, 248, 182, 478, 486, 590]. deformations [21, 170, 340, 654, 345]. deformed [928]. deforming [471, 610]. Dehazing [967]. Dense [941, 492, 532, 938, 891, 40]. Densely [916]. Density [689, 744, 927, 789, 164, 440]. density-aware [789]. density-based [440]. Density-enhanced [689]. density-flow [744]. dependent [230, 631]. deployable [819]. depth [573, 475, 614, 806, 772, 570, 365]. depth-of-field [573]. derivatives [543]. descent [765]. description [208, 184]. descriptors [529, 840, 604]. desiccation [926]. Design [987, 816, 63, 485, 174, 778, 842, 215, 958, 699]. Designing [548, 819, 205, 925]. desk [57]. desk-sized [57]. Detail [788, 457, 474, 322, 677, 772, 574]. Detail-aware [457]. Detail-feature-preserving [474]. Detail-preserving [788]. detailed [654, 660, 687]. details [642, 853, 615, 643, 798, 903]. detection [201, 650, 756, 223, 151, 974, 68, 172, 903]. detection- [756]. Determining [733]. developable [540]. Development [290, 327, 34, 770]. devices [367, 770, 397]. diagram [839]. difference [558]. different [642]. differentiable [922]. differential [629]. differentiated [856]. differently [659]. diffraction [859]. diffusion [186, 401, 523]. digital [942, 676, 409]. dimension [64, 811]. dimensional [249, 202, 540, 123]. direct [412]. Directable [225]. directed [309]. Directing [787]. direction [927]. directly [565]. Disassembling [826]. disclosure [361, 748]. discontinuities [197]. discourse [4, 34]. discrete [920]. disk [917]. disocclusion [849]. display [961, 397, 110, 766, 173]. display-based [766]. displays [57, 873]. dissection [693]. dissipation [621]. dissolution [796]. dissonance [633]. distance [291, 853]. distant [992]. Distributed [741, 807, 222, 663, 258, 718, 878]. distribution [469, 468, 313, 215]. Divergence [251, 962]. Divergence-constrained [251]. divergence-free [962]. diviseme [51]. divisor [517]. DKD [699]. dolls [35]. domain [478]. dominant [651]. Draft [531]. Draft-space [531]. Dramaturgical [633]. drawing [846, 524]. drawings [335]. drawn [425]. driven [79, 835, 593, 3, 280, 654, 848, 274, 588, 853, 23, 197, 456, 718, 687, 896, 425, 707, 452, 196, 190, 929, 357]. drivers [733]. drop [464]. Dual [180, 555, 478, 581].

Dual-domain [478]. Dual-space [555]. DualSPHysics [871]. ductile [594]. during [691, 987, 744, 905, 906, 696, 934]. dyadic [437]. Dynamic [187, 618, 87, 248, 917, 176, 834, 195, 170, 155, 506, 89, 136, 484, 495, 680, 444, 223, 134, 620, 210, 864, 390, 83, 95, 306, 787, 921, 433, 308, 244, 39, 699, 875, 429, 610, 798, 253, 489]. dynamical [780]. Dynamically [237]. dynamicist [727]. dynamics [118, 675, 292, 278, 230, 640, 509, 758, 401, 679, 808, 632].

Easy [392, 44]. EasyHome [561]. ECG [37]. ecosystems [451, 517]. edge [520]. edge-guided [520]. Editing [28, 869, 19, 986, 556, 563, 177, 534, 123, 294]. edition [958]. editor [317, 162]. Editorial [419, 449, 441, 442, 569, 601, 627, 682, 737, 329, 545, 257, 60, 947, 1, 7, 12, 49, 53, 103, 114, 115, 156, 169, 211, 212, 224, 264, 311, 318, 362, 368, 411, 374, 427, 481, 488, 494, 501, 538, 551, 560, 608, 717, 272, 381, 323, 435, 508, 617, 728, 730, 774, 782, 785, 794, 802, 855, 822, 829, 874, 877, 884, 889, 924, 932, 939, 973, 982, 989, 895]. Editors [175]. edn [102]. education [970]. effect [951, 813]. Effective [638, 624]. Effects [158, 307, 138, 692, 750, 927, 100, 446, 567]. Efficient [438, 108, 700, 232, 654, 451, 58, 105, 378, 677, 992, 701, 313, 436, 624, 400, 407, 665, 443, 892, 826, 223, 647, 958, 69, 587, 688, 912]. Effort [118]. eGaLab [64]. EigenFires [516]. elastic [513, 225, 534, 370, 678]. elasticity [628]. electrical [37]. electrodermal [906]. element [823, 22, 591]. elicit [361]. Elva [63]. Embedding [908, 509, 664]. embodied [547, 907, 126, 467, 307, 949, 4, 216, 63, 499, 284]. embodiment [906]. embroidery [736]. emergency [993]. emergent [979]. EMG [280]. EMG-data [280]. EmoDescriptor [985]. emotion [869, 2, 313, 696, 389, 979]. emotion-related [313]. Emotional [968, 127, 233, 985, 835, 467, 526, 336]. emotions [547, 337]. empirically [284]. empower [356]. Emulating [229]. enabled [187, 575]. endoscopy [404]. Endovascular [683, 694]. Energized [693]. Energy [796]. Energy-based [796]. engagement [467, 284]. engines [691]. Enhanced [690, 94, 120, 689, 966]. enhancement [706, 454, 941, 772]. Enhancing [968]. enrichment [619]. Entropy [74]. Entropy-based [74]. Environment [740, 645, 208, 267, 32, 944, 844, 496, 658, 754, 886, 503, 159, 780, 313, 873, 95, 804, 862]. Environmental [133]. Environments [984, 598, 834, 818, 908, 659, 237, 222, 778, 612, 522, 606, 495, 367, 680, 704, 850, 444, 223, 48, 620, 330, 589, 810, 158, 553, 413, 258, 625, 23, 557, 662, 766, 489]. equation [629]. erosion [111]. Essential [735]. estimation [723, 291, 424, 378, 347, 812, 315, 628]. Euler [602]. evacuation [645, 586, 744, 873, 436, 993]. evader [658]. evaluating [316]. Evaluation [606, 820, 987, 454, 663, 814, 697, 816, 63]. event [450, 979]. evolution [970, 929]. Example [886, 25, 808, 455, 943, 178, 357]. Example-based [886, 25, 808, 455, 357]. examples [177, 164]. Experience [968, 634, 754]. experiences [216, 950]. experiment [803, 856]. experiments [873, 314]. Expert [944]. Explicit [416, 633]. exploding [933]. Exploitation [770]. exploration [942, 563, 821]. exploratory [685]. Exploring [970, 911, 905]. explosion [933]. ExpressGesture [969]. expression [915, 127, 695, 80, 953, 125, 338]. expressions [547, 467, 31, 336, 337]. Expressive [30, 969, 288]. Extended [226, 640]. Extending [670, 85, 918]. externalized [342]. extraction [375, 491, 821, 638, 74, 631, 41, 937]. Eye [906, 65]. Eyelid [332]. fabrication [845]. fabrics [384]. FABRIK [670]. face [36, 127, 986, 201, 293, 393, 625, 765, 920, 395]. FaceCaps [953]. faces [335, 425, 142]. Facetons [625]. Facial [137, 915, 149, 231, 321, 519, 547, 572, 467, 31, 4, 392, 3, 479, 349, 293, 470, 391, 695, 341, 80, 204, 336, 394, 433, 337, 953, 452]. Facilitating [215]. factorization [961]. faithful [252]. falling [128]. Fast [10, 471, 552, 280, 609, 713, 297, 110, 198, 731, 252, 151, 699, 387]. faster [605]. FASTSWARM [896]. Feature [477, 376, 383, 985, 120, 840, 80, 533, 474, 967]. Feature-based [477, 383, 80]. features [79, 291, 902, 32, 27, 65, 623, 293, 692]. Feedback [758, 905, 847, 503, 662]. feel [905]. fellowship [409]. FEM [139]. few [848, 954]. few-shot [954]. fidelity [751, 110]. field [469, 961, 573, 292, 520, 555, 158, 46]. fields [787, 923, 40]. Filament [755, 923]. Filament-aware [923]. Filament-based [755]. filters [471]. findings [511]. fine [840, 994]. fine-grained [840, 994]. finite [422, 823, 22, 591]. Fire [516, 758, 993]. firefighters [943]. firefly [975]. Fireworks [283, 887]. first [162]. fish [828]. fitting [937]. fixations [906]. fixed [9]. Flexible [515, 556, 980, 292, 207]. flicking [446]. flight [396]. FLIP [772]. floating [832]. Flock [584, 238]. flocking [741]. flocks [727]. flow [641, 885, 744, 842, 871, 164, 677, 518, 609, 40]. flower [651]. flows [543]. Fluid [945, 542, 295, 354, 790, 963, 706, 788, 619, 885, 15, 251, 612, 86, 512, 596, 772, 758, 17, 930, 161, 372, 14, 305, 707, 643, 798, 912]. fluid-solid [161]. fluidity [523]. fluids [513, 644, 85, 731, 16, 864, 250, 615, 462, 929]. fluorescence [50]. flying [836, 896]. Foam [945]. Foams [945]. focus [563]. Folding [899]. foliage [319]. footprint [357]. footprint-driven [357]. Footstep [390]. force [935, 401, 339, 934]. forehead [613]. forest [319, 853, 110]. forest-based [853]. Foreword [66]. formation [872, 456]. formations [466]. formulation [515]. fraction [354]. fractional [543]. fracture [515, 594, 870, 386, 463]. fracturing [300]. frame [320]. frames [309, 721]. framework [818, 869, 26, 820, 8, 165, 266, 391, 823, 254, 891, 896, 688, 294, 993, 284]. free [21, 581, 420, 962]. freehand [576]. freezing [746]. frequency [54]. friction [370]. friendly [405]. full [558, 420]. full-body [420]. Functionally [56]. functions [517]. fundamental [839]. Furstyling [285]. fusion [902, 8, 840, 783, 965, 903, 967]. future [215]. Fuzzy [371, 359, 416]. FVDM [558]. fwd [992].

gait [502, 747]. gallery [63]. Game [109, 830, 831, 467, 951, 399, 733, 366, 687]. Game-ready [830, 831]. gameplay [536]. Games [968, 422, 588, 786, 541, 940]. gaming [797, 317]. GAN [991]. Garment [637, 759]. gaseous [85]. Gaussian [579]. gaze [805, 276, 814, 762]. GCN [916]. gender [951]. general [165, 351]. Generalized [259]. generated [871]. Generating [657, 143, 812, 203]. Generation [233, 310, 897, 948, 454, 100, 167, 907, 684, 152, 969, 183, 408, 301, 414, 528, 588, 13, 983, 128, 631, 624, 567, 130, 763, 667, 665, 81]. generative [842, 991, 993]. generator [389]. Generic [2, 208, 869]. genetic [359, 577, 42]. genetic-fuzzy [359]. genus [143]. genus- [143]. geodesic [245]. geometric [881, 291, 201, 15, 278, 623]. geometrical [726]. Geometry [23, 29, 638, 20, 407]. Geometry-driven [23]. gestual [992]. gesture [907, 969, 135, 770, 328, 702, 990]. gestures [776]. Giant [652]. GIS [364]. Give [307]. glass [409]. global [340, 827]. global/local [340]. glossy [48]. gloves [150]. GPU [405, 641, 384, 512, 265, 689, 403, 266, 210, 823, 377, 68, 22, 14, 688]. GPU-assisted [641]. GPU-based [265, 266, 68]. GPU-friendly [405]. GPUs [171]. Gradient [183]. Gradient-based [183]. grading [531]. graffiti [676]. grained [840, 994]. granular [182]. Graph [712, 406, 450, 966, 343]. graph-based [406]. Graph-cut [712]. Graphics [218, 926, 885]. graphs [379]. Grasp [249]. grass [87]. grazing [880]. grey [890]. grid [267, 928, 930, 609, 282]. grids [58]. Grigore [102]. grounded [606]. group [805, 741, 820, 566, 810]. group-based [805]. groups [897, 509, 236]. growth [451, 24]. Guest [317, 162, 175]. Guidance [646, 658]. guide [54, 63]. guided [472, 270, 520, 302, 221, 991]. guidewires [402].

Hair [148, 779, 210, 952, 830, 831, 687, 923]. hairstyle [396]. Hand [242, 776, 910, 416, 369, 147, 203, 812, 816, 425]. hand-drawn [425]. hand-held [369]. hand-object [910]. Hand-painted [242]. handheld [532]. handler [851]. Handling [365, 299, 67, 779, 892, 505, 250, 823]. hands [719]. HAO [923]. HAO-CNN [923]. Haptic [267, 363, 505, 185, 402, 320, 119, 847, 503, 640, 366, 575, 639]. Haptic-constraint [363]. haptic-enabled [575]. HapticFlow [19]. haptics [19, 119, 223, 221]. Hardware [506]. Hardware-accelerated [506]. hazard [871]. haze [756]. head [942, 79, 724, 887, 65, 629, 873, 766, 571]. head-mounted [887, 873, 766]. healthy [659]. heart [256, 37]. heat [852]. height [555]. held [369]. herd [838]. herds [880]. Heterogeneous [789, 522, 512, 367, 704, 585, 591]. hidden [697]. Hierarchical [872, 582, 590, 463, 94, 417, 867]. hierarchies [701]. hierarchy [484, 116]. High [751, 165, 756, 912]. High-fidelity [751]. high-order [912]. high-quality [165]. high-visibility [756]. hip [269, 39]. HMD [862]. home [561]. horse [502]. hot [85]. HRP [314]. HRP-2 [314]. hug [307]. hugging [949]. Human [980, 271, 549, 847, 203, 414, 576, 976, 42, 579, 482, 9, 202, 672, 529, 949, 27, 423, 820, 671, 127, 556, 745, 748, 750, 927, 147, 82, 806, 234, 577, 330, 525, 533, 321, 229, 714, 337, 715, 72, 398, 937, 129, 653, 400, 860, 667, 957, 81, 571]. human-motion [714]. Human-virtual [847, 949]. Humanoid [975, 825, 890, 635, 799]. humanoids [314]. humans [237, 361, 312, 83, 541, 631, 125]. humor [748]. Hybrid [902, 649, 764, 598, 985, 880, 660, 287, 330, 640, 799, 930, 823, 22, 331, 591, 574]. Hybrid-based [649]. Hydraulic [111]. hydrodynamics [513, 515, 885, 623, 543, 615, 930, 962, 567]. hydrodynamics-based [567]. hydrodynamics-fluid [930].

ice [692]. iCutter [412]. ICWall [173]. identification [604, 952, 900]. illumination [347, 705, 95, 519]. illusions [664]. Image [736, 882, 838, 93, 841, 843, 270, 220, 322, 800, 893, 639, 792, 894, 31, 408, 61, 475, 941, 886, 520, 704, 712, 540, 964, 857, 920, 967]. Image-based [736, 838, 220, 322, 61, 540]. Image-inspired [639]. images [752, 475, 565, 756, 830, 831, 857, 42]. imitation [541, 993]. immersion [369]. Immersive [927, 420, 905, 803, 754, 821, 769]. immiscible [354, 250]. impaired [659]. Implementation [113, 890]. Implicit [69, 416, 930]. implicits [562]. impostors [438]. improve [146, 667]. Improved [962, 881, 385, 634, 14]. improvement [476]. improves [492]. Impulse [176, 286]. Impulse-based [176, 286]. in-game [951]. including [76, 613]. incompressible [885, 512, 17]. inconsistency [350, 157]. incorporates [458]. Incorporating [731, 805]. incrementally [116]. individual [371]. indoor [886, 971]. induced [834]. inertial [160]. Inextensible [370]. inference [450]. influence [748, 847]. Information [373, 380, 410, 418, 426, 487, 434, 448, 480, 493, 500, 507, 537, 544, 550, 626, 793, 801, 809, 817, 824, 833, 863, 868, 876, 883, 888, 931, 946, 972, 981, 988, 559, 568, 600, 607, 616, 655, 656, 661, 666, 673, 681, 711, 716, 722, 729, 738, 767, 775, 784, 622, 27, 700, 99, 971, 158, 215, 365, 507]. information-rich [158]. information-theoretic [622]. informed [208]. Inhomogeneous [348]. ink [913]. insect [896]. insects [836]. inserting [429]. inspired [845, 866, 763, 514, 639]. instruction [856]. instructions [26]. instructors [951]. integrated [145, 364, 159]. Integrating [141, 526]. integration [69]. intelligence [825]. intelligent [797, 825, 290, 413, 242, 490]. intensity [858]. interactant [361]. Interacting [951]. Interaction [769, 850, 16, 546, 208, 790, 648, 978, 949, 369, 847, 166, 753, 770, 382, 104, 503, 458, 328, 216, 992, 286, 23, 483, 554, 372, 921, 631, 243, 44, 639, 994]. interactions [987, 320, 367, 828, 420, 161]. Interactive [304, 723, 202, 170, 776, 709, 885, 90, 138, 384, 650, 676, 892, 959, 563, 459, 534, 565, 528, 312, 580, 397, 207, 303, 128, 678, 679, 461, 977, 571, 284, 263, 45, 388, 835, 363, 399, 618, 751, 153, 266, 705, 712, 37, 76, 958, 631, 694, 172, 22, 688, 63, 762, 763, 862, 490]. interactivity [492]. Interconnecting [663]. Interest [718, 222]. Interest-driven [718]. interface [214, 897, 686, 606, 356, 166, 898, 719, 366]. interfaces [174]. interfacial [197]. interpolation [405, 86, 194, 519, 46, 957]. intersection [605, 151]. interview [987]. intrinsic [68]. Introducing [536, 451]. Introduction [175, 317, 162]. intuitive [328, 142]. invasive [402]. inverse [219, 73, 188, 455, 667]. Investigating [839]. investigation [38]. Inviscid [17]. involving [223]. iridal [751]. Iris [954]. ISBN [102]. isometric [606]. Issue [218, 373, 380, 410, 418, 426, 487, 434, 448, 480, 493, 500, 507, 537, 544, 601, 550, 559, 568, 600, 607, 616, 626, 655, 656, 661, 666, 673, 681, 711, 716, 722, 729, 738, 767, 775, 784, 793, 801, 809, 817, 824, 833, 863, 868, 876, 883, 888, 931, 946, 972, 981, 988, 502, 545, 800, 411, 481, 488, 494, 501, 538, 551, 560, 608, 717, 323, 728, 730, 774, 782, 785, 794, 802, 982, 989, 442, 317, 271, 947, 162, 175, 435, 617, 822, 829]. issue [877, 884, 889, 932, 939, 973, 874, 924]. Italian [154].

Jacobian [667]. Jacobian-based [667]. jets [739]. jitter [814]. job [987]. joint [269, 438, 910, 334, 39]. joints [819]. juggler [504, 710]. Just [742]. Just-in-time [742].

kaleidoscope [116]. kernel [549]. Key [124, 26, 721]. key-frames [721]. key-postures [26]. Key-styling [124]. keyframe [375, 491, 937]. keyframing [226]. keying [304]. keypoints [245]. Kinect [672]. Kinect^{T M} [536]. kinematics [219, 73, 188, 332, 455, 667]. kinetics [289]. kinetics-assisted [289]. knowledge [83, 878].

LabanDancer [72]. laboratory [856]. Lagrange [370]. Lagrangian [621]. Lambertian [322]. Landslide [871]. Landslide-generated [871]. lane [647, 587]. lane-changing [587]. Language [263, 897, 898, 397, 288, 78]. laparoscopic [735]. Laplacian [180, 535, 348]. Large [945, 237, 688, 110, 574]. Large-Scale [945, 688, 110, 574]. laryngoplasty [270]. latency [146]. latent [915]. lattice [403, 428, 140]. lattice-based [428]. layer [573, 296, 393, 772]. layers [821]. layout [174, 955, 158]. layouts [816]. leaf [90, 24]. leafy [761]. learnable [10]. learned [263, 173]. Learning [740, 201, 958, 957, 263, 879, 948, 910, 954, 526, 852, 541, 124, 857, 940, 977, 994, 993, 630]. Learning-based [201, 958]. least [251, 789, 720]. leaves [292]. Lessons [173, 263]. Let [710]. level [960, 562, 558, 677, 851, 574]. level-of-detail [574]. levels [83]. LiDAR [854]. life [32, 762]. lifting [596]. ligament [497]. light [723, 961, 844, 751, 825, 705, 48, 428, 967]. lighting [279]. lightning [763]. lightweight [971, 713]. like [67, 611]. limb [797]. Limbless [795]. limited [976]. line [199, 223, 13]. line-based [223]. linear [176, 938, 139, 116]. lion [153]. lip [232]. lip-synch [232]. liquid [739, 708, 88, 660, 197, 462, 514]. liquids [18]. listening [261, 987]. Literature [878, 970]. Live [532]. LMA [118]. LMA-Effort [118]. Local [986, 340, 604, 621, 86, 296, 341]. localization [844, 854, 891]. Locally [777, 391]. location [468]. locomotion [325, 495, 150, 199, 421, 13, 702]. locomotive [703]. locomotor [780]. logic [342]. lookup [378]. loss [934]. low [249, 202, 844, 906, 298, 925, 851, 123, 967]. low-budget [906]. low-cost [925]. low-dimensional [249, 202, 123]. low-level [851]. low-light [844, 967]. luminance [520].

macaque [821]. machine [422, 327]. machining [327]. Macroscopic [592]. magnetic [546]. maintenance [807, 369, 990]. making [467, 526, 254, 306]. management [208, 222, 104, 413]. Managing [236]. maneuvering [741]. manifold [879]. manipulation [709, 837, 850, 640, 282, 858, 345]. manipulative [244]. mannequins [112]. map [165, 925, 852]. Mapping [62, 844, 85, 891]. maps [167, 371, 417, 854, 259]. marionette [629]. marker [485, 420, 816]. markerless [41]. markers [398]. Markov [697]. markup [288]. massive [818]. master [913]. matching [406, 969, 403, 300, 94, 246, 245]. material [138, 182]. materials [732, 678, 331]. MAVE [754]. Maze [754]. Maze-based [754]. me [307, 905]. mean [453]. mean-value [453]. measure [29]. measurement [9]. measures [622]. measuring [700]. mechanism [826]. media [768, 428]. medialization [270]. mediated [753]. Medical [218, 563, 909]. medicine [404]. members [506]. mental [771]. Mesh [21, 97, 99, 177, 684, 113, 445, 790, 593, 19, 350, 476, 376, 138, 654, 848, 143, 96, 530, 134, 298, 29, 326, 875, 294, 486, 590, 665, 489]. Mesh-free [21]. meshes [464, 180, 562, 17, 308, 631, 713, 339, 478]. Meshless [386, 71, 484, 928, 591]. metaballs [363]. Metabots [663]. metamorphosis [96, 40]. Metaverse [786]. method [724, 185, 165, 85, 503, 955, 252, 930, 377, 715, 591, 707, 765, 140, 912]. methodology [268]. methods [8, 258]. metric [759]. microscopic [592]. Microsoft [536]. midpoint [69]. Minimal [136]. minimally [402]. minimizing [309]. minimum [259]. minimum-norm [259]. mirror [43]. miscible [354, 250, 140]. missing [398]. mission [528]. mission-based [528]. mixed [304, 648, 910, 935, 660, 786, 609, 804]. Mixing [55]. mixtures [140]. Mobile [873, 658, 213, 977]. Mobile-powered [873]. Mobility [659]. mocap [399]. modal [320, 904]. model [670, 45, 938, 59, 880, 910, 619, 935, 292, 956, 65, 871, 599, 652, 663, 199, 726, 916, 558, 585, 458, 928, 647, 210, 864, 239, 324, 511, 540, 780, 983, 84, 401, 161, 433, 830, 831, 587, 22, 990, 783, 542, 762, 632, 765, 482, 862, 571]. model-based [324, 571]. model-reduced [619]. modeler [168]. Modeling [586, 27, 155, 82, 382, 727, 18, 243, 979, 732, 363, 736, 10, 423, 139, 299, 90, 136, 451, 517, 650, 633, 116, 220, 192, 971, 234, 421, 952, 87, 260, 922, 764, 920]. Modelling [415, 195, 629, 611, 766]. models [249, 141, 881, 202, 672, 908, 986, 721, 376, 840, 654, 697, 203, 351, 414, 51, 247, 640, 625, 714, 41, 688]. modular [116]. momentum [73, 302, 182]. momentum-based [73, 182]. monocular [854, 891, 857, 966, 860, 42]. monotonic [86]. mood [632]. morphable [986]. Morphing [46, 98, 273, 97, 180, 88, 194, 143, 584, 761, 20, 344, 142]. mosaic [454, 689]. Motion [106, 52, 879, 227, 757, 75, 123, 39, 314, 78, 117, 77, 9, 548, 536, 79, 202, 26, 118, 120, 980, 529, 189, 671, 122, 375, 703, 556, 745, 956, 369, 747, 301, 549, 731, 914, 959, 73, 752, 535, 99, 459, 685, 534, 916, 287, 274, 231, 414, 503, 577, 51, 247, 312, 163, 341, 516, 580, 719, 812, 330, 25, 13, 207, 112, 810, 334, 533, 578, 816, 302, 74, 148, 229, 576, 76, 714, 124, 315, 398, 937, 400, 860, 514, 667, 42, 579, 957, 482, 81, 630, 357, 379]. motions [776, 184, 204, 128]. motivated [301]. motivations [861]. motor [122]. mounted [887, 873, 766]. MOUT [191, 254, 306]. movement [942, 927, 503, 100, 795, 78]. movement-based [942]. Movements [984, 985, 132]. moving [251, 43, 905, 658, 789, 720, 198]. MRI [269, 821, 39]. Muddy [642]. Multi [960, 443, 67, 393, 810, 244, 310, 32, 416, 320, 959, 335, 562, 159, 772, 823, 389]. multi-body [416]. multi-character [959]. Multi-core [443]. multi-GPU [823]. Multi-group [810]. Multi-layer [393, 772]. Multi-level [960, 562]. multi-modal [320]. multi-object [310]. multi-perceptive [32]. Multi-resolution [67, 244]. multi-sensorial [32]. multi-stimulus [389]. multi-user [159]. multi-view [335]. Multiagent [993, 948, 861]. multiblock [58]. multicameras [604]. multicharacter [950]. multifactor [414]. Multifinger [753]. multigrid [512, 252]. Multilayer [840, 942]. multilegged [495]. Multimodal [125, 141, 869, 8, 5, 904]. multiphase [871]. multiphysical [764]. multiplayer [770, 104, 503]. Multiple [61, 890, 799, 462, 937, 628]. multiplicity [657]. multipliers [370]. Multiresolution [319]. Multiscale [375, 965]. Multitask [857]. multiuser [557]. multivariate [579]. Multiview [849, 899, 941]. Multiview-coherent [849]. mummification [926]. muscle [82, 130]. musculotendon [726]. museums [878]. music [301]. Myriad [157].

NAO [825]. narratives [861]. Natural [79, 496, 603, 50, 897, 926, 898]. navigating [144]. Navigation [799, 684, 606, 825, 975, 364, 890, 620, 852, 827, 390, 191, 787, 489]. nearest [760]. needle [221]. neighbor [718]. neighbor-organizing [718]. neighbour [760]. net [196, 914]. network [902, 842, 846, 344, 858, 965, 967, 957]. network-based [344]. Networked [503, 313]. networking [704]. networks [548, 839, 991]. neural [839, 911]. No [102, 314]. noise [496, 28]. nominal [184]. non [322, 47, 519]. non-Lambertian [322]. non-photorealistic [47]. non-skin [519]. noncollinear [471]. nonholonomic [741, 807]. nonorthotropic [881]. nonrigid [806]. norm [259]. normalized [867]. notation [72, 78]. Novel [772, 447, 715, 891, 575]. novices [356]. number [976]. numerical [621]. numerous [451]. NVEs [113].

object [310, 948, 709, 910, 484, 901, 473, 132, 903, 198]. objects [208, 536, 406, 108, 139, 476, 750, 220, 382, 444, 328, 992, 413, 615, 679, 974, 664, 591, 243, 131]. obscurances [48]. observations [703, 747]. obstacles [658]. occluded [992]. occlusion [910, 614, 217, 171]. occlusions [485, 365]. ocean [107, 382, 808, 832]. octree [668, 701]. odometry [854]. offline [908]. On-line [13, 199]. One [900, 396, 587]. One-shot [900]. Online [703, 429, 541, 878, 561]. ontology [127]. open [733]. optical [39]. Optimal [837, 580]. optimisation [799]. optimising [359]. Optimization [566, 645, 910, 815, 849, 520, 955, 890, 260, 646, 859, 665]. Optimization-based [566]. Optimized [117, 139, 491, 438, 721, 778, 231, 453]. optimizing [612]. order [912]. ordinary [629]. Organizing [184, 718]. orientation [309]. oriented [594, 34, 862]. out-of-core [688]. outdoor [9, 347]. overactuated [781]. overview [400]. Oz [126].

P2P [718]. packets [859]. packing [826]. pages [102]. Paint [409, 45, 91, 523]. Paint-on-glass [409]. painted [242]. Painterly [498]. painting [913]. palpation [187]. panel [44]. panorama [564]. Panoramas [52]. Paper [502, 192]. paper-cut [192]. Papers [218, 561, 162]. paradigms [986]. Parallel [963, 612, 171, 798, 512, 351, 443]. Parallel-optimizing [612]. parameterization [97, 430, 143, 326]. parameters [118]. Parametric [445, 244]. parametrization [698]. parametrized [531]. parsimonious [780]. part [986]. part-based [986]. participating [428]. Particle [790, 464, 483, 372, 513, 515, 963, 885, 668, 623, 731, 543, 615, 930, 161, 795, 646, 962, 567, 764]. Particle-based [464, 483, 372, 963, 795, 764]. Particle-mesh [790]. particles [594]. partition [562]. partitioning [745, 105]. PASCAL [325]. passive [557]. patch [194, 899]. patch-based [194]. Patches [296]. path [165, 522, 289, 235, 496, 61, 685, 971, 827, 133, 456, 646, 400, 379]. path-based [289, 61]. pathfinding [443]. patient [565, 221]. patient-specific [565]. Pattern [945, 516]. patterns [734, 353]. pavement [664]. PC [157]. PCMD [632]. PDE [724, 19]. PDE-based [724, 19]. pedestrian [648, 353, 239, 509]. pedestrians [84]. peer [818, 258, 157]. peer-to-peer [818, 258, 157]. Pencil [524, 846]. people [307, 41]. per-joint [438]. Perception [773, 145, 951, 83, 229]. perceptions [813]. perceptive [32]. Perceptual [291, 671, 814, 689, 360, 727]. Perceptually [466, 340, 301]. Performance [274, 425, 445, 797, 3, 663, 697, 391, 566, 158, 519]. performance-captured [445]. Performance-driven [274, 425, 3]. performance-sensitive [797]. performances [696, 539]. peripheral [911]. Person [604, 978, 900]. personalities [805]. personality [137, 33, 2, 585, 733, 632]. personality-characterized [632]. personalized [83, 632, 395]. Perspective [241, 786, 727, 259]. Perspective-aware [241]. perspectives [557]. Phaneros [818]. phase [462]. phenomena [791]. Philippe [102]. photo [664]. photograph [637]. photon [85]. Photorealistic [460, 417, 47]. PhyCoVIS [936]. Physical [706, 908, 936, 119, 750, 597, 458, 928, 23, 591]. Physical-based [706]. Physically [98, 521, 179, 473, 24, 613, 763, 141, 472, 252, 345]. Physically-based [24, 613]. physics [548, 691, 325]. physics-based [548]. Physiological [271, 834, 911]. Physiologically [256]. physiology [82]. piano [539]. picture [132]. piecewise [918]. PIK [202]. pilot [925]. pipeline [983]. pipelines [552]. Pipelining [704]. Piying [554]. planar [43]. planners [379]. Planning [620, 490, 165, 522, 235, 496, 348, 872, 971, 444, 827, 133, 810, 400]. planning-based [872]. plant [24]. plants [776]. plastic [255]. platform [126, 797, 364, 413]. plausible [466]. play [554]. player [813]. players [733]. plume [749]. Point [919, 98, 225, 223, 20]. point-constraints [225]. point-sampled [98, 20]. Points [686, 976]. Points-based [686]. Poisson [512]. polarized [938]. policies [657]. polite [527]. Politeness [492]. polycube [97]. polycube-based [97]. polygonal [351, 614, 583, 339]. polynomial [918]. Pompeii [55]. Populating [768, 589, 237]. Porous [518, 732]. portals [105]. Pose [430, 291, 177, 378, 812, 958, 860]. pose-editing [177]. poses [653]. posing [686, 892]. Position [675, 479, 791, 644, 723, 534, 640, 929]. Position-based [675, 479, 791, 640, 929]. positional [848]. positioning [759]. possible [777, 131]. postproduction [878]. postural [987]. posture [980, 27, 219, 815]. postures [26, 27]. potential [15]. power [832]. powered [873]. Practical [403, 463]. precise [837]. Precomputing [190]. predator [828]. Predictable [691]. Predicting [870, 398, 662]. prediction [871, 916]. Predictive [721, 145]. preference [468]. prefetching [146]. preparation [603]. presence [106, 369, 750, 754]. preservation [629, 486, 590]. preserved [47]. preserving [59, 788, 456, 474]. Pressure [295, 469]. prey [828]. primitives [122, 625]. principal [719]. priority [962]. proactive [850]. probabilistic [249, 33, 477, 553]. Procedural [495, 517, 153, 260, 952]. procedure [497]. process [823, 913]. processes [764, 579]. professionals [943]. profiles [127]. programming [33, 81]. Progressive [96, 705, 200, 713, 875]. project [72]. projection [76, 707]. propagation [938, 638]. propagation-based [938]. property [865]. prosodic [79]. Provision [369]. proximity [351, 337]. proxy [700]. Pseudo [292, 138]. Pseudo-dynamics [292]. PSO [990]. PSO-BPNN [990]. Psychological [84, 511]. public [215]. pulmonary [38]. puncture [221]. puppetry [770].

quadrangulation [275]. quality [165, 205]. quantitative [439]. Quartz [191]. quasi [166]. quasi-tangible [166]. quaternion [581].

race [456]. racer [456]. racing [733]. radio [54]. radio-frequency [54]. raindrops [417, 461]. random [853]. randomized [712]. range [262]. Rapid [570, 515]. rapidly [746]. rates [320, 751]. ray [960, 265, 555, 701]. ray-tracing [960]. RBF [344]. re [604, 900]. re-identification [604, 900]. reaction [186]. reactive [423, 73, 850]. ready [983, 830, 831]. Real [432, 594, 399, 152, 573, 71, 92, 385, 278, 502, 708, 65, 522, 484, 150, 346, 695, 928, 210, 640, 144, 578, 37, 91, 206, 683, 315, 567, 305, 331, 591, 783, 860, 193, 909, 338, 440, 108, 880, 464, 392, 165, 641, 703, 107, 460, 451, 517, 223, 188, 55, 13, 417, 95, 435, 124, 14, 865, 896, 762, 662, 965, 804]. real-life [762]. Real-time [432, 594, 399, 152, 573, 71, 92, 385, 278, 502, 708, 65, 522, 484, 150, 346, 695, 928, 210, 640, 144, 578, 37, 91, 206, 683, 567, 305, 331, 591, 860, 193, 909, 338, 440, 108, 392, 165, 641, 703, 107, 460, 451, 517, 188, 13, 417, 95, 435, 124, 14, 896, 965, 804]. Real-virtual [783]. Realistic [732, 107, 147, 525, 615, 636, 523, 387, 963, 155, 612, 85, 755, 558, 757, 133, 372, 510]. realities [978]. reality [214, 268, 304, 723, 834, 432, 887, 797, 910, 943, 369, 906, 803, 378, 347, 821, 950, 992, 786, 54, 970, 360, 213, 162, 102, 891, 977, 909, 44, 365, 804, 769]. reality-based [797]. rear [76]. reasoning [527]. recognition [263, 902, 980, 432, 31, 399, 840, 752, 954, 757, 696, 953, 990, 994]. recoloring [886]. Reconstructing [672]. reconstruction [532, 43, 549, 322, 806, 154, 540, 976, 497, 857, 923, 860, 542, 42, 474, 395, 571]. reduce [621]. reduced [619, 420, 816]. reduction [811]. refined [377]. refinement [609]. reflection [132, 542]. reflective [417]. refraction [790, 206]. refractions [708]. refractive [417]. region [364, 94, 519, 624, 142]. region-based [142]. regions [3, 849]. regression [549, 772]. regulate [337]. reinforcement [948, 940]. rejuvenation [36]. related [747, 313]. relational [291]. relationships [744]. relative [869]. relief [955, 921]. Relighting [520]. remeshing [431]. Remote [57, 753]. removal [756]. Render [961]. Render-based [961]. renderer [922]. Rendering [50, 141, 438, 262, 790, 108, 736, 319, 573, 8, 185, 155, 498, 107, 402, 521, 320, 708, 846, 85, 470, 101, 58, 346, 391, 555, 144, 417, 18, 95, 355, 47, 87, 206, 428, 664, 649, 339, 688, 574]. repair [694]. Repairing [350]. representation [450, 660, 398, 574]. representing [752]. reprojection [564]. research [423, 873]. resolution [706, 67, 680, 397, 244, 610]. Resolving [910]. respiration [338]. respond [659]. response [195]. responses [834, 307]. responsible [526]. Responsive [261]. Restoration [554]. retargeting [106, 914, 959, 391, 394, 321]. retargetting [341]. retexturing [93]. retrieval [622, 853, 533, 576, 904, 714, 78, 482]. reusing [209]. revealing [547]. Review [102, 920, 944, 498, 970, 771, 974, 878]. RGB [723, 752, 860, 903, 571]. RGB-D [723, 903, 571]. rhesus [821]. rhinoplasty [348]. rhythmic [153]. rich [897, 158]. richer [547]. Riemannian [879]. rig [690]. rig-space [690]. Rigging [352, 277, 447]. Rigid [514, 777, 675, 731, 933, 286, 377, 679, 246, 282, 331]. rigid-body [675, 933]. Rigid-motion-inspired [514]. Ripple [443]. river [281]. roadmaps [553]. robotic [635]. robots [807, 799]. Robust [262, 844, 815, 779, 562, 172, 20, 485, 912]. rods [370]. Role [940, 31]. Role-based [940]. Rom [621]. rotation [309, 101, 602]. rotation-minimizing [309]. rule [907]. rural [558].

saliency [375]. salient [520, 903, 245]. sampled [98, 592, 20]. sampling [796, 194, 653, 665]. sandstorm [179]. saturation [518]. scalable [443, 222, 726, 144, 157]. scalar [469, 706]. Scale [945, 854, 688, 110, 574, 245]. Scale-aware [854]. scanning [396]. scans [742, 293, 168, 395]. scattering [206]. scenario [528, 588, 993]. scenarios [263]. scene [206, 858, 919]. scenes [960, 155, 408, 558, 55, 95, 699, 429, 867, 574]. scheduling [646]. scheme [406, 222]. science [423]. scissor [819]. scratches [692]. screen [790, 708, 158, 76]. screen-space [708]. Screwing [862]. sculpting [56]. SDF [660]. Search [443]. searching [712]. Segmentation [725, 879, 956, 901, 530, 577, 712, 715, 964, 965, 919, 862]. Segmentation-based [725]. selected [691]. selection [700]. selective [415]. self [361, 748, 757, 643]. self-adaptive [643]. self-disclosure [361, 748]. Semantic [994, 572, 589, 533, 904, 964, 966, 965, 919]. semantics [903]. semi [621, 698, 769, 891]. semi-automatic [698]. semi-dense [891]. semi-Immersive [769]. semi-Lagrangian [621]. sensing [150]. sensitive [797, 376]. sensor [571]. sensorial [32]. sensors [9, 160, 980, 549, 806, 546, 570]. sequence [602, 875]. sequences [310, 445, 593, 152, 350, 375, 135, 752, 99, 459, 530, 602, 148, 131]. sequential [680]. serious [588, 413]. server [663]. Service [984]. services [216]. set [420, 830, 831]. sets [821]. settings [762]. shader [279]. shadow [554, 259]. shadowing [303, 95]. Shape [836, 47, 238, 345, 622, 406, 405, 59, 887, 90, 870, 149, 403, 227, 341, 300, 246, 282, 595, 920, 294, 590]. Shape-constrained [836, 238, 887]. shape-controllable [595]. shape-preserving [59]. shapes [412]. shared [367, 503, 804]. Sharp [945]. Shear [101, 253, 864]. Shear-rotation-warp [101]. shell [183, 278, 285, 518, 649]. short [776]. shot [954, 900]. sickness [911]. sign [397, 288]. signals [4]. Silhouette [76, 277]. Silhouette-based [76]. similarity [229, 714]. Simple [351]. simplification [117, 308]. simplified [935]. simulate [163, 864]. simulated [987]. Simulating [276, 743, 543, 509, 510, 880, 82, 330]. Simulation [734, 739, 388, 983, 984, 221, 462, 683, 140, 269, 644, 36, 176, 657, 371, 267, 881, 648, 938, 594, 691, 358, 944, 887, 820, 2, 839, 926, 619, 476, 935, 885, 385, 251, 353, 384, 402, 612, 668, 746, 289, 796, 119, 465, 484, 512, 346, 669, 565, 651, 690, 348, 386, 473, 234, 364, 647, 505, 252, 463, 210, 852, 133, 511, 144, 583, 771, 693, 735, 772, 592, 758, 483, 615, 791, 570, 17, 930, 401, 372, 435, 823, 518, 91, 760, 69, 436, 24, 497, 694, 832, 795, 646, 649, 14, 866, 896, 305, 297, 331, 707, 523, 383, 909, 327, 198, 253, 387]. simulation [440]. Simulation-as-a-Service [984]. Simulations [945, 67, 751, 191, 254, 306]. simulator [118, 586, 255, 403, 733, 911]. simulators [268]. Simultaneous [628, 844, 891]. Single [475, 952, 573, 614, 695, 519, 920]. single-depth [614]. single-layer [573]. Single-view [952]. singularity [420]. singularity-free [420]. site [55]. situ [327]. Situation [342]. size [31, 158]. sized [57]. Skeletal [848, 82]. skeleton [593, 902, 752, 315, 129, 957]. skeleton-based [902, 752]. skeleton-driven [593]. skeletonization [377]. skeletons [638]. Sketch [887, 842, 922, 356, 654, 576, 991]. Sketch-based [887, 842, 922, 654]. Sketch2VF [842]. sketching [275]. Skill [630]. skin [36, 280, 430, 149, 519, 865, 130, 297, 248]. skinning [725, 581, 296, 134, 200, 377, 865, 129, 196]. Sky [756, 206]. SLAM [966]. sliding [297]. small [897, 509, 830, 831]. smoke [92, 385, 89, 668, 289, 749, 595, 198]. Smooth [339]. smoothed [513, 515, 885, 623, 543, 615, 930, 962, 567]. smoothed-particle [615, 567]. smoothing [756]. smoothing-based [756]. smoothness [469]. Snap [254]. snow [649]. soap [652]. Social [748, 698, 768, 467, 586, 307, 674, 599, 750, 633, 527, 239, 509, 762]. socially [361, 526]. socket [819]. soft [170, 384, 640, 693, 401, 433, 22]. soft-fabrics [384]. soft-touch [170]. software [162]. solid [619, 660, 483, 161, 372, 462, 331, 387]. solids [513, 70, 731, 16]. solution [881, 781]. solutions [611]. solver [512, 962]. song [6]. SOR [962]. sound [141, 933, 100, 808]. source [723]. sourcing [698]. South [944]. Space [444, 11, 572, 790, 915, 325, 745, 708, 531, 430, 534, 690, 555, 970, 852, 112, 321, 456, 653, 245, 357, 920]. Space-time [444, 11]. spaces [123]. SpaceTime [634]. sparse [549, 398]. Spatial [309, 353, 966, 226, 385, 599, 750, 457, 861, 846]. Spatial-temporal [353]. spatio [706]. spatio-temporal [706]. Spatiotemporal [535]. Special [218, 271, 502, 175, 435, 567, 317]. specific [565, 602, 221]. Specifying [4]. spectral [331]. specular [542]. speech [263, 30, 835, 969, 51, 204]. speed [927]. Speeding [476]. SPEM [185]. SPH [295, 788, 612, 796, 286, 864, 305, 643, 798]. sphere [472]. spherical [143, 417]. spider [845]. splatting [470]. spline [918, 433, 917]. split [285]. spray [853]. spring [59]. sprites [155, 618]. square [246, 720]. squares [251, 789]. SRNPD [846]. Stable [186, 355, 615, 636]. standing [815]. state [529, 498, 422, 452]. state-aware [452]. state-of-the-art [498]. states [233]. static [496]. stationary [658]. statistical [745, 347]. SteerBench [316]. steering [657, 359, 199, 342, 316]. step [421, 662, 357]. step-by-step [421]. stereo [899, 941, 217, 173]. stereographic [557]. Stereoscopic [408, 770]. stiffness [779, 300]. stimulus [389]. stochastic [358]. storytelling [770]. strain [46]. strategies [217, 171]. strategy [658, 827]. Streaming [610, 113, 875]. street [61, 558, 664]. street-level [558]. strength [869]. string [845, 578]. strings [446]. strokes [75]. structural [393]. structure [668, 465, 486, 611]. structured [598]. structures [819, 745, 582, 921, 867]. study [834, 951, 925, 258, 760, 14]. Style [228, 430, 124]. styling [124]. stylization [846]. Stylized [204, 279, 167, 355, 664]. Subdivision [431, 611]. subsampling [900]. subspace [486]. subspaces [228]. Subtle [692]. suite [316]. summarization [475]. sums [517]. Supervised [765, 904]. support [160]. supported [127]. Surface [945, 857, 245, 262, 59, 870, 623, 470, 806, 132, 759, 518, 542, 474, 766]. surface-based [766]. surface-deformable [59]. surfaces [98, 660, 220, 322, 638, 540, 68, 172, 610, 611]. Surgeries [683, 694]. surgery [416, 255, 403, 505, 693, 735, 983]. surgical [268, 402, 497]. surrogates [635]. surrounds [905]. survey [423, 213]. suturing [909]. swarm [799, 646, 896]. swarm-adaptive [799]. sweeps [203]. swept [472]. Symmetric [293, 635]. synch [232]. synchronized [37]. synergy [781]. synergy-based [781]. Synthesis [142, 77, 249, 79, 202, 26, 479, 122, 745, 502, 199, 194, 393, 849, 287, 755, 51, 247, 516, 580, 719, 397, 204, 123, 603, 456, 124, 808, 240, 857, 244, 595, 913, 288, 343, 564, 579, 241, 357]. synthesizer [26]. Synthesizing [945, 5, 950, 869]. Synthetic [360, 814, 437]. system [485, 942, 359, 416, 676, 348, 328, 54, 404, 696, 191, 497, 694, 832, 795, 990, 182, 327, 561, 539, 966]. systematic [970]. systems [769, 213, 76, 39, 44].

t [535]. table [378]. tabletop [166, 159]. tactile [949, 905]. Tai [26]. taking [50]. talk [702]. tangible [536, 166, 554]. Target [197, 929]. Target-driven [197, 929]. task [34, 490]. tasks [850]. tearing [446]. technique [992, 799]. techniques [113, 8, 498, 527, 133, 735, 760]. technological [786]. technologies [109, 213]. technology [162, 102]. tele [634]. tele-conference [634]. teleported [634]. telepresence [753, 635]. template [447]. template-based [447]. Temporal [135, 194, 262, 706, 353, 454, 530, 344]. terrain [408]. terrains [477]. test [605, 151]. tetrahedral [294, 665]. tetrahedralization [960]. tetrahedralization-based [960]. text [907, 740]. text-to-gesture [907]. texton [712]. texture [477, 155, 65, 194, 614, 941, 756, 677, 47]. texture-based [677]. textures [281, 278, 285, 221, 649, 244]. TFAN [298]. theater [696]. their [740]. theoretic [622, 27]. theories [633]. theory [388]. thickening [864]. thin [739, 623, 678]. threads [739]. three [556, 540, 162]. three-dimensional [540]. three-way [556]. throwing [781]. tiled [173]. Time [70, 299, 176, 108, 432, 594, 880, 399, 464, 152, 392, 573, 742, 185, 165, 71, 641, 703, 92, 385, 107, 278, 502, 320, 708, 65, 522, 484, 460, 451, 517, 150, 346, 534, 695, 444, 223, 928, 29, 188, 210, 827, 640, 13, 11, 144, 578, 37, 417, 95, 435, 254, 306, 396, 91, 124, 206, 683, 14, 896, 315, 567, 305, 331, 591, 860, 193, 909, 965, 804, 338, 440]. time-based [827]. Time-critical [70, 299]. time-of-flight [396]. time-variant [29]. time-varying [320]. tissue [640, 693, 401, 433, 22]. titration [856]. together [119]. tomography [38]. tool [936, 232, 412, 44]. topic [956, 482]. topological [106, 350]. Topologically [761]. Topology [720, 422, 282]. Topology-aware [720, 282]. toroidal [749]. torsional [370]. touch [187, 170, 307]. touch-enabled [187]. Tour [132, 63]. toy [35]. tracing [960, 701]. tracking [485, 160, 310, 980, 432, 152, 324, 765]. traditional [663, 554]. traffic [648, 935, 558, 585, 647, 587, 783]. training [943, 936, 597, 528, 990, 909]. traits [733]. trajectories [510]. trajectory [405]. Transductive [343]. transfer [915, 430, 695, 352, 519, 457, 478]. Transferring [921]. transform [596]. transformation [703, 747]. transient [157]. transition [207, 462]. transitions [516]. transmission [718]. transnasal [505]. Transparent [901]. transplantation [333]. transport [743, 751]. transportation [948]. travel [119]. tree [650, 100, 761, 461, 699, 190, 611]. tree-like [611]. trees [948, 451, 517]. triangle [605, 17, 151]. triangle-to-triangle [605]. triangular [376, 180, 562, 326, 564, 478]. Tubular [326]. tumour [54]. Turbulence [595, 385]. turbulent [749, 755, 615, 643, 912]. twisting [776, 446].

U [914]. U-Net [914]. ubiquitous [397]. UEGM [389]. UGSC [991]. UGSC-GAN [991]. ultrasound [266, 221]. umbrella [216]. uncertain [388, 389]. Understanding [122, 949, 847, 858]. unification [331]. unified [465, 250, 161, 294]. unit [952]. units [915, 726]. unity [562]. universe [504]. Unsupervised [956]. up-sampled [592]. update [65, 699]. upper [797]. urban [237, 260, 574]. Use [160, 943, 402]. User [472, 991, 214, 686, 174, 159, 158]. User-guided [472, 991]. users [598, 987, 659, 753, 898]. Using [146, 163, 553, 9, 472, 644, 879, 622, 536, 768, 515, 776, 948, 424, 97, 573, 839, 201, 885, 745, 796, 484, 749, 293, 549, 73, 663, 496, 803, 697, 99, 143, 203, 177, 475, 527, 614, 565, 849, 770, 503, 444, 806, 635, 516, 300, 852, 799, 733, 546, 570, 396, 221, 631, 41, 713, 646, 519, 398, 990, 455, 345, 40, 142, 929, 42, 365, 482, 611, 571]. utilizing [820]. utterances [5].

validated [340, 436, 284]. validity [671]. value [453]. variance [900]. variant [29]. variation [235, 579]. variational [914]. variations [106, 477, 437]. varying [320]. vector [923, 383]. Vehicle [648, 994]. velocity [658, 558]. venation [734, 90]. venation-based [90]. ventriculostomy [267]. verbal [361]. verbs [136]. versatile [416, 870]. versus [873]. vessel [641]. vessels [525]. via [945, 520, 596, 955, 578, 157, 962, 937, 282, 591, 903]. viable [742]. Video [450, 453, 240, 40, 452, 160, 310, 152, 820, 93, 320, 181, 524, 695, 786, 921, 519, 131, 429, 900, 365, 979]. video-based [900, 365]. Video-driven [452]. videos [901, 756, 860]. View [230, 700, 335, 158, 952]. View-dependent [230]. viewpoint [622]. views [837]. Virtual [968, 469, 361, 786, 38, 83, 984, 337, 856, 564, 805, 598, 263, 233, 208, 268, 388, 504, 768, 834, 818, 835, 187, 978, 710, 32, 145, 944, 887, 951, 949, 797, 10, 466, 839, 127, 987, 416, 943, 659, 743, 237, 506, 222, 369, 606, 748, 599, 255, 367, 451, 517, 750, 905, 927, 415, 663, 847, 906, 769, 803, 265, 527, 680, 754, 770, 290, 503, 850, 597, 528, 821, 223, 328, 950, 992, 312, 163, 925, 970, 55, 589, 404, 154, 780, 810, 158, 733, 983, 313, 413, 873, 258, 625, 23, 162, 332, 582, 773, 557]. virtual [787, 631, 587, 718, 990, 783, 909, 44, 63, 762, 662, 131, 429, 561, 632, 993, 766, 804, 125, 102]. viscous [45, 18]. Visibility [818, 756, 397]. Visibility-based [818]. visible [51]. vision [163, 360]. vision-based [360]. Visual [746, 596, 137, 30, 936, 840, 415, 503, 928, 324, 854, 832]. visualization [734, 233, 724, 963, 706, 266, 677, 39, 575, 62]. visuo [122]. visuo-motor [122]. vKASS [497]. volcano [764]. Volleyball [968]. Volume [354, 800, 8, 484, 451, 789, 101, 705, 701, 433, 221]. volumes [472]. Volumetric [486, 472, 565, 348, 705, 38, 37, 923]. Vortex [668]. vortices [749]. vorticity [385, 934]. VR [968, 612, 364, 157, 76, 862]. VR-systems [76]. VRCAI’08 [323]. VRCIA [175].

wagging [87]. wake [755]. Walk [702].

walking [905, 927, 112, 41, 662].

walkthrough [146, 61]. walkthroughs [205]. wall [641]. warp [101]. warping [405, 531, 401]. water [642, 746, 623, 382, 132, 483, 921, 518, 857, 859, 193, 243, 383]. waterfall [155]. watering [384]. waters [50]. watery [91]. Wave [945, 938, 871, 859]. wavelet [496, 596, 407]. wavelet-based [407]. waves [107, 382, 808, 383]. way [273, 556]. web [845, 216]. web-inspired [845]. Web-services [216]. Web3D [713]. Weighted [341, 721, 188]. wet [779]. wetting [791]. wheeled [658]. whole [606]. whole-body [606]. Wiley [102]. wind [292, 87, 832]. wire [345]. wireframe [355]. wireless [213]. wisp [210]. withering [473]. Wizard [126]. Wizard-of-Oz [126]. wolf [890]. work [423]. workplaces [215]. world [388, 786]. worlds [119, 663, 718]. wound [393]. wrestling [422]. wrinkles [392, 613].

X [978]. X-person [978]. xvi [102].

years [951]. York [102].

zone [260]. Zoom [992]. Zoom-fwd [992].

References

Magnenat-Thalmann:2004:Ea

[1] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 15(1):i, March 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Egges:2004:GPE

[2] Arjan Egges, Sumedha Kshirsagar, and Nadia Magnenat-Thalmann. Generic personality and emotion simulation for conversational agents. Computer Animation and Virtual Worlds, 15(1):1–13, March 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Fidaleo:2004:ACA

[3] Douglas Fidaleo and Ulrich Neumann. Analysis of co-articulation regions for performance-driven facial animation. Computer Animation and Virtual Worlds, 15(1):15–26, March 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

DeCarlo:2004:SAF

[4] Doug DeCarlo, Matthew Stone, Corey Revilla, and Jennifer J. Venditti. Specifying and animating facial signals for discourse in embodied conversational agents. Computer Animation and Virtual Worlds, 15(1):27–38, March 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kopp:2004:SMU

[5] Stefan Kopp and Ipke Wachsmuth. Synthesizing multimodal utterances for conversational agents. Computer Animation and Virtual Worlds, 15(1):39–52, March 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

King:2004:AS

[6] Scott A. King and Richard E. Parent. Animating song. Computer Animation and Virtual Worlds, 15(1):53–61, March 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2004:Eb

[7] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 15(2):i, May 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ferre:2004:FFM

[8] Maria Ferre, Anna Puig, and Dani Tost. A framework for fusion methods and rendering techniques of multimodal volume data. Computer Animation and Virtual Worlds, 15(2):63–77, May 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Aminian:2004:CHM

[9] Kamiar Aminian and Bijan Najafi. Capturing human motion using body-fixed sensors: outdoor measurement and clinical applications. Computer Animation and Virtual Worlds, 15(2):79–94, May 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Dinerstein:2004:FLB

[10] Jonathan Dinerstein, Parris K. Egbert, Hugo de Garis, and Nelson Dinerstein. Fast and learnable behavioral and cognitive modeling for virtual character animation. Computer Animation and Virtual Worlds, 15(2):95–108, May 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pasko:2004:STB

[11] Galina Pasko, Alexander Pasko, and Tosiyasu Kunii. Space-time blending. Computer Animation and Virtual Worlds, 15(2):109–121, May 2004. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2004:Ec

[12] Nadia Magnenat-Thalmann. Editorial. Computer Animation and Virtual Worlds, 15(3–4):123–124, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2004:LMB

[13] Sang Il Park, Hyun Joon Shin, Tae Hoon Kim, and Sung Yong Shin. On-line motion blending for real-time locomotion generation. Computer Animation and Virtual Worlds, 15(3–4):125–138, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2004:ISR

[14] Enhua Wu, Youquan Liu, and Xuehui Liu. An improved study of realtime fluid simulation on GPU. Computer Animation and Virtual Worlds, 15(3–4):139–146, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hong:2004:CFA

[15] Jeong mo Hong and Chang hun Kim. Controlling fluid animation with geometric potential. Computer Animation and Virtual Worlds, 15(3–4):147–157, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Muller:2004:IFD

[16] Matthias Müller, Simon Schirm, Matthias Teschner, Bruno Heidelberger, and Markus Gross. Interaction of fluids with deformable solids. Computer Animation and Virtual Worlds, 15(3–4):159–171, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shi:2004:IIF

[17] Lin Shi and Yizhou Yu. Inviscid and incompressible fluid simulation on triangle meshes. Computer Animation and Virtual Worlds, 15(3–4):173–181, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Steele:2004:MRV

[18] Kevin Steele, David Cline, Parris K. Egbert, and Jonathan Dinerstein. Modeling and rendering viscous liquids. Computer Animation and Virtual Worlds, 15(3–4):183–192, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Duan:2004:HPB

[19] Ye Duan, Jing Hua, and Hong Qin. HapticFlow: PDE-based mesh editing with haptics. Computer Animation and Virtual Worlds, 15(3–4):193–200, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiao:2004:RMP

[20] Chunxia Xiao, Wenting Zheng, Qunsheng Peng, and A. R. Forrest. Robust morphing of point-sampled geometry. Computer Animation and Virtual Worlds, 15(3–4):201–210, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chang:2004:MFD

[21] Jian Chang and Jian J. Zhang. Meshfree deformations. Computer Animation and Virtual Worlds, 15(3–4):211–218, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2004:HCF

[22] Wen Wu and Pheng Ann Heng. A hybrid condensed finite element model with GPU acceleration for interactive 3D soft tissue cutting. Computer Animation and Virtual Worlds, 15(3–4):219–227, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Schmidl:2004:GDP

[23] Harald Schmidl and Ming C. Lin. Geometry-driven physical interaction between avatars and virtual environments. Computer Animation and Virtual Worlds, 15(3–4):229–236, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2004:PBS

[24] Iris R. Wang, Justin W. L. Wan, and Gladimir V. G. Baranoski. Physicallybased simulation of plant leaf growth. Computer Animation and Virtual Worlds, 15(3–4):237–244, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2004:EBM

[25] Min Je Park and Sung Yong Shin. Example-based motion cloning. Computer Animation and Virtual Worlds, 15(3–4):245–257, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chao:2004:TCS

[26] Shih-Pin Chao, Chih-Yi Chiu, Shi-Nine Yang, and Tsang-Gang Lin. Tai Chi synthesizer: a motion synthesis framework based on key-postures and motion instructions. Computer Animation and Virtual Worlds, 15(3–4):259–268, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

DeSilva:2004:MHA

[27] P. Ravindra De Silva and Nadia Bianchi-Berthouze. Modeling human affective postures: an information theoretic characterization of posture features. Computer Animation and Virtual Worlds, 15(3–4):269–276, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yoon:2004:EN

[28] Jong-Chul Yoon, In-Kwon Lee, and Jung-Ju Choi. Editing noise. Computer Animation and Virtual Worlds, 15(3–4):277–287, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mathur:2004:MMC

[29] Prasun Mathur, Chhavi Upadhyay, Parag Chaudhuri, and Prem Kalra. A measure for mesh compression of timevariant geometry. Computer Animation and Virtual Worlds, 15(3–4):289–296, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bevacqua:2004:EAV

[30] Elisabetta Bevacqua and Catherine Pelachaud. Expressive audio-visual speech. Computer Animation and Virtual Worlds, 15(3–4):297–304, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cunningham:2004:RIS

[31] Douglas W. Cunningham, Manfred Nusseck, Christian Wallraven, and Heinrich H. Bülthoff. The role of image size in the recognition of conversational facial expressions. Computer Animation and Virtual Worlds, 15(3–4):305–310, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Conde:2004:ALE

[32] Toni Conde and Daniel Thalmann. An artificial life environment for autonomous virtual agents with multisensorial and multi-perceptive features. Computer Animation and Virtual Worlds, 15(3–4):311–318, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chittaro:2004:BPA

[33] Luca Chittaro and Milena Serra. Behavioral programming of autonomous characters based on probabilistic automata and personality. Computer Animation and Virtual Worlds, 15(3–4):319–326, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2004:BDT

[34] John R. Lee and Andrew B. Williams. Behavior development through task oriented discourse. Computer Animation and Virtual Worlds, 15(3–4):327–337, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Matsumoto:2004:ACT

[35] Naoko Matsumoto, Akifumi Tokosumi, and Yoko Hirai. Affection for cohabitant toy dolls. Computer Animation and Virtual Worlds, 15(3–4):339–346, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bastanfard:2004:TAS

[36] A. Bastanfard, O. Bastanfard, H. Takahashi, and M. Nakajima. Toward anthropometrics simulation of face rejuvenation and skin cosmetic. Computer Animation and Virtual Worlds, 15(3–4):347–352, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ryan:2004:RTI

[37] John Ryan, Carol O’Sullivan, and Christopher Bell. Real-time interactive volumetric animation of the heart’s electrical cycle from automatically synchronized ECG. Computer Animation and Virtual Worlds, 15(3–4):353–360, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Perchet:2004:VIP

[38] Diane Perchet, Catalin I. Fetita, Laurence Vial, Françoise Prêteux, George Caillibotte, Gabriela Sbiêrlea-Apiou, and Marc Thiriet. Virtual investigation of pulmonary airways in volumetric computed tomography. Computer Animation and Virtual Worlds, 15(3–4):361–376, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yahia-Cherif:2004:MCV

[39] Lydia Yahia-Cherif, Benjamin Gilles, Tom Molet, and Nadia Magnenat-Thalmann. Motion capture and visualization of the hip joint with dynamic MRI and optical systems. Computer Animation and Virtual Worlds, 15(3–4):377–385, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2004:VMU

[40] Yizhou Yu and Qing Wu. Video metamorphosis using dense flow fields. Computer Animation and Virtual Worlds, 15(3–4):387–397, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wagg:2004:AME

[41] David K. Wagg and Mark S. Nixon. Automated markerless extraction of walking people using deformable contour models. Computer Animation and Virtual Worlds, 15(3–4):399–406, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2004:HMR

[42] Jianhui Zhao and Ling Li. Human motion reconstruction from monocular images using genetic algorithms. Computer Animation and Virtual Worlds, 15(3–4):407–414, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jang:2004:MPM

[43] Kyung Ho Jang, Dong Hoon Lee, and Soon Ki Jung. A moving planar mirror based approach for cultural reconstruction. Computer Animation and Virtual Worlds, 15(3–4):415–423, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yuan:2004:VIP

[44] M. L. Yuan, S. K. Ong, and Andrew Y. C. Nee. The virtual interaction panel: an easy control tool in augmented reality systems. Computer Animation and Virtual Worlds, 15(3–4):425–432, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baxter:2004:VPM

[45] William Baxter, Yuanxin Liu, and Ming C. Lin. A viscous paint model for interactive applications. Computer Animation and Virtual Worlds, 15(3–4):433–441, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yan:2004:MBS

[46] Han-Bing Yan, Shi-Min Hu, and Ralph Martin. Morphing based on strain field interpolation. Computer Animation and Virtual Worlds, 15(3–4):443–452, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2004:STP

[47] Haitao Wang, Jian J. Zhang, Stan Z. Li, and Yangsheng Wang. Shape and texture preserved non-photorealistic rendering. Computer Animation and Virtual Worlds, 15(3–4):453–461, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mendez-Feliu:2004:CLA

[48] Àlex Méndez-Feliu and Mateu Sbert. Combining light animation with obscurances for glossy environments. Computer Animation and Virtual Worlds, 15(3–4):463–470, July 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2004:Ed

[49] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 15(5):i, December 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cerezo:2004:RNW

[50] E. Cerezo and F. J. Seron. Rendering natural waters taking fluorescence into account. Computer Animation and Virtual Worlds, 15(5):471–484, December 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ma:2004:AAV

[51] Jiyong Ma, Ronald Cole, Bryan Pellom, Wayne Ward, and Barbara Wise. Accurate automatic visible speech synthesis of arbitrary 3D models based on concatenation of diviseme motion capture data. Computer Animation and Virtual Worlds, 15(5):485–500, December 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bartoli:2004:MP

[52] Adrien Bartoli, Navneet Dalal, and Radu Horaud. Motion panoramas. Computer Animation and Virtual Worlds, 15 (5):501–517, December 2004. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2005:Ea

[53] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 16(1):i–ii, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nicolau:2005:ARS

[54] S. Nicolau, A. Garcia, X. Pennec, L. Soler, and N. Ayache. An augmented reality system to guide radio-frequency tumour ablation. Computer Animation and Virtual Worlds, 16(1):1–10, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Papagiannakis:2005:MVR

[55] George Papagiannakis, Sébastien Schertenleib, Brian O’Kennedy, Marlene Arevalo-Poizat, Nadia Magnenat-Thalmann, Andrew Stoddart, and Daniel Thalmann. Mixing virtual and real scenes in the site of ancient Pompeii. Computer Animation and Virtual Worlds, 16(1):11–24, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Adzhiev:2005:FBA

[56] Valery Adzhiev, Peter Comninos, Maxim Kazakov, and Alexander Pasko. Functionally based augmented sculpting. Computer Animation and Virtual Worlds, 16(1):25–39, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ashdown:2005:RCD

[57] Mark Ashdown and Peter Robinson. Remote collaboration on desk-sized displays. Computer Animation and Virtual Worlds, 16(1):41–51, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2005:ERM

[58] Jinho Lee, Lance C. Burton, Raghu Machiraju, and Donna S. Reese. Efficient rendering of multiblock curvilinear grids with complex boundaries. Computer Animation and Virtual Worlds, 16(1):53–68, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2005:ASD

[59] Yoo-Joo Choi, Min Hong, Min-Hyung Choi, and Myoung-Hee Kim. Adaptive surface-deformable model with shapepreserving spring. Computer Animation and Virtual Worlds, 16(1):69–83, February 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kunii:2005:E

[60] Tosiyasu Kunii and Hock Soon Seah. Editorial. Computer Animation and Virtual Worlds, 16(2):i–ii, May 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2005:MPB

[61] Dong Hoon Lee and Soon Ki Jung. Multiple path-based approach to imagebased street walkthrough. Computer Animation and Virtual Worlds, 16(2): 85–95, May 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zheng:2005:MCC

[62] Jiang Yu Zheng and Min Shi. Mapping cityscapes into cyberspace for visualization. Computer Animation and Virtual Worlds, 16(2):97–107, May 2005. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yuan:2005:DEE

[63] Xiang Yuan and Yam San Chee. Design and evaluation of Elva: an embodied tour guide in an interactive virtual art gallery. Computer Animation and Virtual Worlds, 16(2):109–119, May 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Masa:2005:END

[64] Michal Masa, Jiri Zara, and Pavel Smetana. eGaLab: a new dimension for artists. Computer Animation and Virtual Worlds, 16(2):121–127, May 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Inostroza:2005:RTU

[65] Patricio Inostroza. Real-time update of eye features on the 3D head model texture. Computer Animation and Virtual Worlds, 16(2):129–136, May 2005. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baciu:2005:F

[66] George Baciu, Ming Lin, Rynson W. H. Lau, and Daniel Thalmann. Foreword. Computer Animation and Virtual Worlds, 16(3–4):137–139, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jain:2005:MRC

[67] Nitin Jain, Ilknur Kabul, Naga K. Govindaraju, Dinesh Manocha, and Ming Lin. Multi-resolution collision handling for cloth-like simulations. Computer Animation and Virtual Worlds, 16(3–4):141–151, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wong:2005:GBI

[68] Wingo Sai-Keung Wong and George Baciu. GPU-based intrinsic collision detection for deformable surfaces. Computer Animation and Virtual Worlds, 16(3–4):153–161, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Volino:2005:IMI

[69] Pascal Volino and Nadia Magnenat-Thalmann. Implicit midpoint integration and adaptive damping for efficient cloth simulation. Computer Animation and Virtual Worlds, 16(3–4):163–175, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Dequidt:2005:TCA

[70] J. Dequidt, D. Marchal, and L. Grisoni. Time-critical animation of deformable solids. Computer Animation and Virtual Worlds, 16(3–4):177–187, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2005:RTM

[71] Xiaohu Guo and Hong Qin. Realtime meshless deformation. Computer Animation and Virtual Worlds, 16(3–4):189–200, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wilke:2005:DNH

[72] Lars Wilke, Tom Calvert, Rhonda Ryman, and Ilene Fox. From dance notation to human animation: The Laban-Dancer project. Computer Animation and Virtual Worlds, 16(3–4):201–211, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Komura:2005:ARM

[73] Taku Komura, Edmond S. L. Ho, and Rynson W. H. Lau. Animating reactive motion using momentum-based inverse kinematics. Computer Animation and Virtual Worlds, 16(3–4):213–223, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

So:2005:EBM

[74] Clifford K. F. So and George Baciu. Entropy-based motion extraction for motion capture animation. Computer Animation and Virtual Worlds, 16(3–4):225–235, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oshita:2005:MCS

[75] Masaki Oshita. Motion control with strokes. Computer Animation and Virtual Worlds, 16(3–4):237–244, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thuring:2005:SBM

[76] Sven Thüring, Jörn Herwig, and Alfred Schmitt. Silhouette-based motion capture for interactive VR-systems including a rear projection screen. Computer Animation and Virtual Worlds, 16(3–4):245–257, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Alankus:2005:AMS

[77] Gazihan Alankus, A. Alphan Bayazit, and O. Burchan Bayazit. Automated motion synthesis for dancing characters. Computer Animation and Virtual Worlds, 16(3–4):259–271, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2005:MRB

[78] Tao Yu, Xiaojie Shen, Qilei Li, and Weidong Geng. Motion retrieval based on movement notation language. Computer Animation and Virtual Worlds, 16(3–4):273–282, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Busso:2005:NHM

[79] Carlos Busso, Zhigang Deng, Ulrich Neumann, and Shrikanth Narayanan. Natural head motion synthesis driven by acoustic prosodic features. Computer Animation and Virtual Worlds, 16(3–4):283–290, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2005:FBA

[80] Bongcheol Park, Heejin Chung, Tomoyuki Nishita, and Sung Yong Shin. A feature-based approach to facial expression cloning. Computer Animation and Virtual Worlds, 16(3–4):291–303, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhuang:2005:AGH

[81] Yueting Zhuang, Jun Xiao, Yizi Wu, Tao Yang, and Fei Wu. Automatic generation of human animation based on motion programming. Computer Animation and Virtual Worlds, 16(3–4):305–318, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lemos:2005:MSD

[82] Robson R. Lemos, Jon Rokne, Gladimir V. G. Baranoski, Yasuo Kawakami, and Toshiyuki Kurihara. Modeling and simulating the deformation of human skeletal muscle based on anatomy and physiology. Computer Animation and Virtual Worlds, 16(3–4):319–330, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Strassner:2005:VHP

[83] Johannes Strassner and Marion Langer. Virtual humans with personalized perception and dynamic levels of knowledge. Computer Animation and Virtual Worlds, 16(3–4):331–342, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sakuma:2005:PMA

[84] Takeshi Sakuma, Tomohiko Mukai, and Shigeru Kuriyama. Psychological model for animating crowded pedestrians. Computer Animation and Virtual Worlds, 16(3–4):343–351, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kang:2005:EPM

[85] Byungkwon Kang, Insung Ihm, and Chandrajit Bajaj. Extending the photon mapping method for realistic rendering of hot gaseous fluids. Computer Animation and Virtual Worlds, 16(3–4):353–363, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ihm:2005:CLM

[86] Insung Ihm, Deukhyun Cha, and Byungkwon Kang. Controllable local monotonic cubic interpolation in fluid animations. Computer Animation and Virtual Worlds, 16(3–4):365–375, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2005:DMR

[87] Changbo Wang, Zhangye Wang, Qi Zhou, Chengfang Song, Yu Guan, and Qunsheng Peng. Dynamic modeling and rendering of grass wagging in wind. Computer Animation and Virtual Worlds, 16(3–4):377–389, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jin:2005:BBL

[88] Xiaogang Jin, Shengjun Liu, Charlie C. L. Wang, Jieqing Feng, and Hanqiu Sun. Blob-based liquid morphing. Computer Animation and Virtual Worlds, 16(3–4):391–403, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hong:2005:ASD

[89] Jin-Kyung Hong and Chang-Hun Kim. Animating smoke with dynamic balance. Computer Animation and Virtual Worlds, 16(3–4):405–414, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hong:2005:IVB

[90] Sung Min Hong, Bruce Simpson, and Gladimir V. G. Baranoski. Interactive venation-based leaf shape modeling. Computer Animation and Virtual Worlds, 16(3–4):415–427, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

VanLaerhoven:2005:RTS

[91] Tom Van Laerhoven and Frank Van Reeth. Real-time simulation of watery paint. Computer Animation and Virtual Worlds, 16(3–4):429–439, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

He:2005:RTC

[92] Haitao He and Duanqing Xu. Realtime cartoon animation of smoke. Computer Animation and Virtual Worlds, 16(3–4):441–449, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2005:IVR

[93] Yanwen Guo, Jin Wang, Xiang Zeng, Zhongyi Xie, Hanqiu Sun, and Qunsheng Peng. Image and video retexturing. Computer Animation and Virtual Worlds, 16(3–4):451–461, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qiu:2005:EAC

[94] Jie Qiu, Hock Soon Seah, Feng Tian, Quan Chen, and Zhongke Wu. Enhanced auto coloring with hierarchical region matching. Computer Animation and Virtual Worlds, 16(3–4):463–473, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tamura:2005:DSR

[95] Naoki Tamura, Henry Johan, and Tomoyuki Nishita. Deferred shadowing for real-time rendering of dynamic scenes under environment illumination. Computer Animation and Virtual Worlds, 16(3–4):475–486, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lin:2005:PMM

[96] Chao-Hung Lin, Tong-Yee Lee, Hung-Kuo Chu, and Chih-Yuan Yao. Progressive mesh metamorphosis. Computer Animation and Virtual Worlds, 16(3–4):487–498, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Fan:2005:MMU

[97] Zhengwen Fan, Xiaogang Jin, Jieqing Feng, and Hanqiu Sun. Mesh morphing using polycube-based crossparameterization. Computer Animation and Virtual Worlds, 16(3–4):499–508, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bao:2005:PBM

[98] Yunfan Bao, Xiaohu Guo, and Hong Qin. Physically based morphing of point-sampled surfaces. Computer Animation and Virtual Worlds, 16(3–4): 509–518, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2005:MDU

[99] Tong-Yee Lee, Ping-Hsien Lin, Shaur-Uei Yan, and Chun-Hao Lin. Mesh decomposition using motion information from animation sequences. Computer Animation and Virtual Worlds, 16(3–4):519–529, July 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Matsuyama:2005:GTM

[100] Katsutsugu Matsuyama, Tadahiro Fujimoto, Kazunobu Muraoka, and Norishige Chiba. Generation of tree movement sound effects. Computer Animation and Virtual Worlds, 16(5): 531–545, December 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kye:2005:SRW

[101] Heewon Kye, Byeong-Seok Shin, Yeong Gil Shin, and Helen Hong. Shearrotation-warp volume rendering. Computer Animation and Virtual Worlds, 16 (5):547–557, December 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vexo:2005:BRV

[102] Frederic Vexo. Book review: Virtual reality technology (2nd edn). Grigore C. Burdea and Philippe Coiffet, Wiley, New York, 2003. No. of pages: xvi + 444. ISBN 0-471-36089-9. Computer Animation and Virtual Worlds, 16 (5):559–560, December 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2005:Eb

[103] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 16(5):i, December 2005. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lin:2006:BBM

[104] Qingping Lin, Weihua Wang, Liang Zhang, Jim Mee Ng, and Chor Ping Low. Behaviour-based multiplayer collaborative interaction management. Computer Animation and Virtual Worlds, 17(1):1–19, February 2006. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lerner:2006:ECP

[105] Alon Lerner, Yiorgos Chrysanthou, and Daniel Cohen-Or. Efficient cells-and-portals partitioning. Computer Animation and Virtual Worlds, 17(1):21–40, February 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baciu:2006:MRP

[106] George Baciu and Bartholomew K. C. Iu. Motion retargeting in the presence of topological variations. Computer Animation and Virtual Worlds, 17(1):41–57, February 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2006:RRT

[107] Yaohua Hu, Luiz Velho, Xin Tong, Baining Guo, and Harry Shum. Realistic, real-time rendering of ocean waves. Computer Animation and Virtual Worlds, 17(1):59–67, February 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cheung:2006:ERD

[108] Gary K. L. Cheung, Rynson W. H. Lau, and Frederick W. B. Li. Efficient rendering of deformable objects for real-time applications. Computer Animation and Virtual Worlds, 17(1):69–81, February 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lau:2006:GT

[109] Rynson W. H. Lau, Daniel Cohen-Or, Baining Guo, and John Buchanan. Game technologies. Computer Animation and Virtual Worlds, 17(1):i–ii, February 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2006:FDL

[110] Huaisheng Zhang, Wei Hua, Qing Wang, and Hujun Bao. Fast display of largescale forest with fidelity. Computer Animation and Virtual Worlds, 17(2):83– 97, May 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Benes:2006:HE

[111] Bedřich Beneš, Václav Těšínský, Jan Hornyš, and Sanjiv K. Bhatia. Hydraulic erosion. Computer Animation and Virtual Worlds, 17(2):99–108, May 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pettre:2006:MCB

[112] Julien Pettre and Jean-Paul Laumond. A motion capture-based control-space approach for walking mannequins. Computer Animation and Virtual Worlds, 17(2):109–126, May 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bouras:2006:IMS

[113] Christos J. Bouras, Alexandros Panagopoulos, and Thrasyvoulos Tsiatsos. Implementation of 3D mesh streaming and compression techniques in NVEs. Computer Animation and Virtual Worlds, 17(2):127–140, May 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2006:Ea

[114] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 17(2):i–ii, May 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2006:Eb

[115] Nadia Magnenat-Thalmann, Enhua Wu, Ana Paiva, and Dinesh Pai. Editorial. Computer Animation and Virtual Worlds, 17(3–4):141–144, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kunii:2006:KCA

[116] Tosiyasu L. Kunii and Kenji Ohmori. A kaleidoscope as a cyberworld and its animation: linear architecture and modeling based on an incrementally modular abstraction hierarchy. Computer Animation and Virtual Worlds, 17 (3–4):145–153, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ahn:2006:OMS

[117] Junghyun Ahn, Seungwoo Oh, and Kwangyun Wohn. Optimized motion simplification for crowd animation. Computer Animation and Virtual Worlds, 17(3–4):155–165, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chao:2006:LES

[118] Shih-Pin Chao, Shi-Nine Yang, and Tsang-Gang Lin. An LMA-Effort simulator with dynamics parameters for motion capture animation. Computer Animation and Virtual Worlds, 17(3– 4):167–177, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jorissen:2006:BHP

[119] Pieter Jorissen, Joan De Boeck, and Wim Lamotte. Bringing haptics and physical simulation together: haptic travel through physical worlds. Computer Animation and Virtual Worlds, 17 (3–4):179–187, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2006:DBA

[120] Quan Chen, Feng Tian, HockSoon Seah, Zhongke Wu, Jie Qiu, and Melikhov Konstantin. DBSC-based animation enhanced with feature and motion. Computer Animation and Virtual Worlds, 17 (3–4):189–198, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hartman:2006:AB

[121] Christopher Hartman and Bedřich Beneš. Autonomous boids. Computer Animation and Virtual Worlds, 17(3–4):199–206, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guerra-Filho:2006:UVM

[122] Gutemberg Guerra-Filho and Yiannis Aloimonos. Understanding visuo-motor primitives for motion synthesis and analysis. Computer Animation and Virtual Worlds, 17(3–4):207–217, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shin:2006:MSE

[123] Hyun Joon Shin and Jehee Lee. Motion synthesis and editing in low-dimensional spaces. Computer Animation and Virtual Worlds, 17(3–4):219–227, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2006:KSL

[124] Yi Wang, Zhi-Qiang Liu, and Li-Zhu Zhou. Key-styling: learning motion style for real-time synthesis of 3D animation. Computer Animation and Virtual Worlds, 17(3–4):229–237, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

deMelo:2006:MEV

[125] Celso de Melo and Ana Paiva. Multimodal expression in virtual humans. Computer Animation and Virtual Worlds, 17(3–4):239–248, ???? 2006.

CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Brown:2006:WOP

[126] Edward Brown and Neil Barrett. A Wizard-of-Oz platform for embodied conversational agents. Computer Animation and Virtual Worlds, 17(3– 4):249–257, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Garcia-Rojas:2006:EFE

[127] A. García-Rojas, F. Vexo, D. Thalmann, A. Raouzaiou, K. Karpouzis, S. Kollias, L. Moccozet, and N. Magnenat-Thalmann. Emotional face expression profiles supported by virtual human ontology. Computer Animation and Virtual Worlds, 17(3–4):259–269, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tang:2006:IGF

[128] Bing Tang, Zhigeng Pan, Le Zheng, and Mingmin Zhang. Interactive generation of falling motions. Computer Animation and Virtual Worlds, 17(3–4):271–279, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2006:CSS

[129] Xiaosong Yang, Arun Somasekharan, and Jian J. Zhang. Curve skeleton skinning for human and creature characters. Computer Animation and Virtual Worlds, 17(3–4):281–292, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2006:AMG

[130] Xiaosong Yang and Jian J. Zhang. Automatic muscle generation for character skin deformation. Computer Animation and Virtual Worlds, 17(3–4):293–303, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2006:CPC

[131] Guofeng Zhang, Xueying Qin, Xiaobo An, Wei Chen, and Hujun Bao. As-consistent-as-possible compositing of virtual objects and video sequences. Computer Animation and Virtual Worlds, 17(3–4):305–314, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2006:TPW

[132] Jinho Park, Nambin Heo, Sunghee Choi, and Sung Yong Shin. Tour into the picture with water surface reflection and object movements. Computer Animation and Virtual Worlds, 17 (3–4):315–324, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Paris:2006:EAP

[133] Sébastien Paris, Stéphane Donikian, and Nicolas Bonvalet. Environmental abstraction and path planning techniques for realistic crowd simulation. Computer Animation and Virtual Worlds, 17 (3–4):325–335, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mamou:2006:SAD

[134] Khaled Mamou, Titus Zaharia, and Françoise Prêteux. A skinning approach for dynamic 3D mesh compression. Computer Animation and Virtual Worlds, 17(3–4):337–346, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Heloir:2006:TAC

[135] Alexis Heloir, Nicolas Courty, Sylvie Gibet, and Franck Multon. Temporal alignment of communicative gesture sequences. Computer Animation and Virtual Worlds, 17(3–4):347–357, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hsieh:2006:MDM

[136] Chi-Min Hsieh and Annie Luciani.

Minimal dynamic modeling for dance verbs. Computer Animation and Virtual Worlds, 17(3–4):359–369, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Arya:2006:FAV

[137] Ali Arya, Lisa N. Jefferies, James T.

Enns, and Steve DiPaola. Facial actions as visual cues for personality. Computer Animation and Virtual Worlds, 17 (3–4):371–382, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2006:IMD

[138] Jin Huang, Hongxin Zhang, Xiaohan Shi, Xinguo Liu, and Hujun Bao. Interactive mesh deformation with pseudo material effects. Computer Animation and Virtual Worlds, 17(3–4):383–392, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Garcia:2006:OLF

[139] Marcos García, César Mendoza, Luis Pastor, and Angel Rodríguez. Optimized linear FEM for modeling deformable objects. Computer Animation and Virtual Worlds, 17(3–4):393–402, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2006:SMB

[140] Hongbin Zhu, Xuehui Liu, Youquan Liu, and Enhua Wu. Simulation of miscible binary mixtures based on lattice Boltzmann method. Computer Animation and Virtual Worlds, 17(3–4):403–410, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Avanzini:2006:IPB

[141] Federico Avanzini and Paolo Crosato. Integrating physically based sound models in a multimodal rendering architecture. Computer Animation and Virtual Worlds, 17(3–4):411–419, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2006:SFU

[142] Yu Zhang and Norman I. Badler. Synthesis of 3D faces using region-based morphing under intuitive control. Computer Animation and Virtual Worlds, 17 (3–4):421–432, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2006:GGM

[143] Tong-Yee Lee, Chih-Yuan Yao, Hung-Kuo Chu, Ming-Jen Tai, and Cheng-Chieh Chen. Generating genus-n-to-m mesh morphing using spherical parameterization. Computer Animation and Virtual Worlds, 17(3–4):433–443, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pettre:2006:RTN

[144] Julien Pettré, Pablo de Heras Ciechomski, Jonathan Maïm, Barbara Yersin, Jean-Paul Laumond, and Daniel Thalmann. Real-time navigating crowds: scalable simulation and rendering. Computer Animation and Virtual Worlds, 17 (3–4):445–455, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Conde:2006:IPA

[145] Toni Conde and Daniel Thalmann. An integrated perception for autonomous virtual agents: active and predictive perception. Computer Animation and Virtual Worlds, 17(3–4):457–468, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hung:2006:UPI

[146] Shao-Shin Hung and Damon Shing-Min Liu. Using prefetching to improve walkthrough latency. Computer Animation and Virtual Worlds, 17(3–4):469–478, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2006:RHH

[147] Jieun Lee, Seung-Hyun Yoon, and Myung-Soo Kim. Realistic human hand deformation. Computer Animation and Virtual Worlds, 17(3–4):479–489, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sugisaki:2006:HMC

[148] Eiji Sugisaki, Yosuke Kazama, Shigeo Morishima, Natsuko Tanaka, and Akiko Sato. Hair motion cloning from cartoon animation sequences. Computer Animation and Virtual Worlds, 17(3– 4):491–499, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2006:FSS

[149] Won-Sook Lee and Andrew Soon. Facial shape and 3D skin. Computer Animation and Virtual Worlds, 17(3– 4):501–512, ???? 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Komura:2006:RTL

[150] Taku Komura and Wai-Chun Lam. Real-time locomotion control by sensing gloves. Computer Animation and Virtual Worlds, 17(5):513–525, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tropp:2006:FTT

[151] Oren Tropp, Ayellet Tal, and Ilan Shimshoni. A fast triangle to triangle intersection test for collision detection. Computer Animation and Virtual Worlds, 17(5):527–535, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Douze:2006:RTG

[152] Matthijs Douze and Vincent Charvillat. Real-time generation of augmented video sequences by background tracking. Computer Animation and Virtual Worlds, 17(5):537–550, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2006:PRC

[153] Tsai-Yen Li and Je-Ren Chen. Procedural rhythmic character animation: an interactive Chinese lion dance. Computer Animation and Virtual Worlds, 17 (5):551–564, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pennestri:2006:CAV

[154] E. Pennestrì, E. Pezzuti, P. P. Valentini, and L. Vita. Computer-aided virtual reconstruction of Italian ancient clocks. Computer Animation and Virtual Worlds, 17(5):565–572, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guan:2006:MRR

[155] Yu Guan, Wei Chen, Lincan Zou, Long Zhang, and Qunsheng Peng. Modeling and rendering of realistic waterfall scenes with dynamic texture sprites. Computer Animation and Virtual Worlds, 17(5):573–583, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2006:Ec

[156] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 17(5):iii, December 2006. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Schaeffer:2007:MSV

[157] Benjamin Schaeffer, Peter Brinkmann, George Francis, Camille Goudeseune, Jim Crowell, and Hank Kaczmarski. Myriad: scalable VR via peer-to-peer connectivity, PC clustering, and transient inconsistency. Computer Animation and Virtual Worlds, 18(1):1–17, February 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Polys:2007:EIL

[158] Nicholas F. Polys, Seonho Kim, and Doug A. Bowman. Effects of information layout, screen size, and field of view on user performance in informationrich virtual environments. Computer Animation and Virtual Worlds, 18 (1):19–38, February 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nakashima:2007:ITE

[159] Kousuke Nakashima, Takashi Machida, Kiyoshi Kiyokawa, and Haruo Takemura. A 2D-3D integrated tabletop environment for multi-user collaboration. Computer Animation and Virtual Worlds, 18(1):39–56, February 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Aron:2007:UIS

[160] Michael Aron, Gilles Simon, and Marie-Odile Berger. Use of inertial sensors to support video tracking. Computer Animation and Virtual Worlds, 18(1):57–68, February 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Solenthaler:2007:UPM

[161] Barbara Solenthaler, Jürg Schläfli, and Renato Pajarola. A unified particle model for fluid-solid interactions. Computer Animation and Virtual Worlds, 18(1):69–82, February 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Singh:2007:GEI

[162] Gurminder Singh, Rynson W. H. Lau, and Yiorgos Chrysanthou. Guest editors’ introduction for the first three papers of this issue: advances in virtual reality software and technology. Computer Animation and Virtual Worlds, 18(1):i–ii, February 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Musse:2007:UCV

[163] Soraia R. Musse, Cláudio R. Jung, Julio C. S. Jacques, Jr., and Adriana Braun. Using computer vision to simulate the motion of virtual agents. Computer Animation and Virtual Worlds, 18(2):83– 93, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Leung:2007:DED

[164] Hoi-Chau Leung and Kin-Chuen Hui.

Deformation by examples: a density flow approach. Computer Animation and Virtual Worlds, 18(2):95–105, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Geraerts:2007:CMM

[165] Roland Geraerts and Mark H. Overmars. The corridor map method: a general framework for real-time high-quality path planning. Computer Animation and Virtual Worlds, 18(2):107–119, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2007:BIQ

[166] Jangho Lee, Jun Lee, HyungSeok Kim, and Jee-In Kim. Believable interaction with a quasi-tangible tabletop interface. Computer Animation and Virtual Worlds, 18(2):121–132, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Adabala:2007:CAG

[167] Neeharika Adabala, Manik Varma, and Kentaro Toyama. Computer aided generation of stylized maps. Computer Animation and Virtual Worlds, 18(2):133–140, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Seo:2007:BMB

[168] Hyewon Seo, Frederic Cordier, and Kyunghi Hong. A breast modeler based on analysis of breast scans. Computer Animation and Virtual Worlds, 18(2):141–151, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2007:Ea

[169] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 18(2):i, May 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2007:IST

[170] Hui Chen, Hanqiu Sun, and Xiaogang Jin. Interactive soft-touch dynamic deformations. Computer Animation and Virtual Worlds, 18(3):153–163, July 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiong:2007:PSO

[171] Hua Xiong, Haoyu Peng, Aihong Qin, and Jiaoying Shi. Parallel strategies of occlusion culling on cluster of GPUs. Computer Animation and Virtual Worlds, 18(3):165–177, July 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wong:2007:RCC

[172] Wingo Sai-Keung Wong and George Baciu. Robust continuous collision detection for interactive deformable surfaces. Computer Animation and Virtual Worlds, 18(3):179–192, July 2007.

CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

vanderSchaaf:2007:LLB

[173] Tom van der Schaaf, Desmond Germans, Henri E. Bal, and Michal Koutek. Lessons learned from building and calibrating the ICWall, a stereo tiled display. Computer Animation and Virtual Worlds, 18(3):193–210, July 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Green:2007:ADL

[174] Mark Green and Wai Leng Lee. Automatic design and layout of 3D user interfaces. Computer Animation and Virtual Worlds, 18(3):211–224, July 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sun:2007:GEI

[175] Hanqiu Sun, Enhua Wu, and George Baciu. Guest Editors’ introduction: special issue on ACM VRCIA 2006. Computer Animation and Virtual Worlds, 18 (3):i–ii, July 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bender:2007:IBD

[176] Jan Bender. Impulse-based dynamic simulation in linear time. Computer Animation and Virtual Worlds, 18(4– 5):225–233, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2007:MPE

[177] Tong-Yee Lee, Chao-Hung Lin, Hung-Kuo Chu, Yu-Shuen Wang, Shao-Wei Yen, and Chang-Rung Tsai. Mesh pose-editing using examples. Computer Animation and Virtual Worlds, 18(4– 5):235–245, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lu:2007:AE

[178] Difei Lu, Xiuzi Ye, and Guomin Zhou. Animating by example. Computer Animation and Virtual Worlds, 18(4– 5):247–257, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2007:PBA

[179] Shiguang Liu, Zhangye Wang, Zheng Gong, Lei Huang, and Qunsheng Peng. Physically based animation of sandstorm. Computer Animation and Virtual Worlds, 18(4–5):259–269, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2007:DLM

[180] Jianwei Hu, Ligang Liu, and Guozhao Wang. Dual Laplacian morphing for triangular meshes. Computer Animation and Virtual Worlds, 18(4–5):271– 277, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2007:CV

[181] Eun-Jung Lee, Ji yong Kwon, and In-Kwon Lee. Caricature video. Computer Animation and Virtual Worlds, 18(4– 5):279–288, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zeng:2007:MBD

[182] Ya-Lun Zeng, Charlie Irawan Tan, Wen-Kai Tai, Mau-Tsuen Yang, Cheng-Chin Chiang, and Chin-Chen Chang. A momentum-based deformation system for granular material. Computer Animation and Virtual Worlds, 18(4–5):289– 300, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2007:GBS

[183] Jin Huang, Xinguo Liu, Haiyang Jiang, Qing Wang, and Hujun Bao. Gradientbased shell generation and deformation. Computer Animation and Virtual Worlds, 18(4–5):301–309, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2007:OMN

[184] Min Je Park and Jieun Cho. Organizing motions with a nominal description. Computer Animation and Virtual Worlds, 18(4–5):311–318, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Garcia:2007:SNH

[185] M. García, J. M. Espadero, L. Pastor, and A. Rodríguez. SPEM: a new haptic rendering method based on time coherence. Computer Animation and Virtual Worlds, 18(4–5):319–328, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2007:SAR

[186] Theodore Kim and Ming Lin. Stable advection-reaction-diffusion with arbitrary anisotropy. Computer Animation and Virtual Worlds, 18(4–5):329– 338, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2007:DTE

[187] Hui Chen, Wen Wu, Hanqiu Sun, and Pheng-Ann Heng. Dynamic touchenabled virtual palpation. Computer Animation and Virtual Worlds, 18(4– 5):339–348, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Meredith:2007:ACB

[188] Michael Meredith and Steve Maddock. Approximating character biomechanics with real-time weighted inverse kinematics. Computer Animation and Virtual Worlds, 18(4–5):349–359, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Courty:2007:CMC

[189] N. Courty and T. Corpetti. Crowd motion capture. Computer Animation and Virtual Worlds, 18(4–5):361–370, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2007:PDD

[190] Long Zhang, Yubo Zhang, Zhongding Jiang, Luying Li, Wei Chen, and Qunsheng Peng. Precomputing data-driven tree animation. Computer Animation and Virtual Worlds, 18(4–5):371–382, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ting:2007:QAN

[191] Shang-Ping Ting and Suiping Zhou. Quartz: an autonomous navigation system for MOUT simulations. Computer Animation and Virtual Worlds, 18(4– 5):383–394, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2007:PCM

[192] Yan Li, Jinhui Yu, Kwan liu Ma, and Jiaoying Shi. 3D paper-cut modeling and animation. Computer Animation and Virtual Worlds, 18(4–5):395–403, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2007:RTC

[193] Jinhui Yu, Xinan Jiang, Haiying Chen, and Cheng Yao. Real-time cartoon water animation. Computer Animation and Virtual Worlds, 18(4–5):405–414, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lai:2007:TTS

[194] Chao-Hung Lai and Jiunn-Lin Wu.

Temporal texture synthesis by patchbased sampling and morphing interpolation. Computer Animation and Virtual Worlds, 18(4–5):415–428, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chang:2007:CBB

[195] Jian Chang, Daniel X. Shepherd, and Jian J. Zhang. Cosserat-beam-based dynamic response modelling. Computer Animation and Virtual Worlds, 18(4– 5):429–436, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2007:BND

[196] Jian J. Zhang, Xiaosong Yang, and Yunfeng Zhao. Bar-net driven skinning for character animation. Computer Animation and Virtual Worlds, 18(4–5):437– 446, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shin:2007:TDL

[197] Seung-Ho Shin and Chang-Hun Kim. Target-driven liquid animation with interfacial discontinuities. Computer Animation and Virtual Worlds, 18(4– 5):447–453, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2007:FSS

[198] Yongxia Zhou, Xiehua Sun, Xiangyang Lan, and Shengyou Lin. Fast smoke simulation of moving object. Computer Animation and Virtual Worlds, 18(4– 5):455–461, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kwon:2007:SML

[199] Taesoo Kwon and Sung Yong Shin. A steering model for on-line locomotion synthesis. Computer Animation and Virtual Worlds, 18(4–5):463–472, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pilgrim:2007:PSC

[200] Simon Pilgrim, Anthony Steed, and Alberto Aguado. Progressive skinning for character animation. Computer Animation and Virtual Worlds, 18(4–5):473– 481, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2007:LBF

[201] Yanwen Guo, Fuyan Zhang, Chunxiao Liu, Hanqiu Sun, and Qunsheng Peng. Learning-based 3D face detection using geometric context. Computer Animation and Virtual Worlds, 18(4–5):483– 492, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Carvalho:2007:ILD

[202] Schubert R. Carvalho, Ronan Boulic, and Daniel Thalmann. Interactive low-dimensional human motion synthesis by combining motion models and PIK. Computer Animation and Virtual Worlds, 18(4–5):493–503, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2007:HHA

[203] Jieun Lee and Myung-Soo Kim. Human hand adaptation using sweeps: generating animatable hand models. Computer Animation and Virtual Worlds, 18(4– 5):505–516, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pei:2007:SSF

[204] Yuru Pei and Hongbin Zha. Stylized synthesis of facial speech motions. Computer Animation and Virtual Worlds, 18(4–5):517–526, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sanyal:2007:DQW

[205] Subhajit Sanyal, Subhashis Banerjee, and Prem K. Kalra. Designing quality walkthroughs. Computer Animation and Virtual Worlds, 18(4–5):527– 538, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2007:RTR

[206] Changbo Wang, Zhangye Wang, and Qunsheng Peng. Real-time rendering of sky scene considering scattering and refraction. Computer Animation and Virtual Worlds, 18(4–5):539– 548, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Peng:2007:IFM

[207] Jen-Yu Peng, I-Chen Lin, Jui-Hsiang Chao, Yan-Ju Chen, and Gwo-Hao Juang. Interactive and flexible motion transition. Computer Animation and Virtual Worlds, 18(4–5):549–558, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Badawi:2007:GDM

[208] M. Badawi and S. Donikian. The generic description and management of interaction between autonomous agents and objects in an informed virtual environment. Computer Animation and Virtual Worlds, 18(4–5):559–569, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2007:ACC

[209] Jun Yu, Yueting Zhuang, Jun Xiao, and Cheng Chen. Adaptive control in cartoon data reusing. Computer Animation and Virtual Worlds, 18(4–5):571– 582, September/December 2007. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oshita:2007:RTH

[210] Masaki Oshita. Real-time hair simulation on GPU with a dynamic wisp model. Computer Animation and Virtual Worlds, 18(4–5):583–593, September/December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2007:Eb

[211] Nadia Magnenat-Thalmann, Sung Yong Shin, Fabian Di Fiore, Daniel Thalmann, and Frank Van Reeth. Editorial. Computer Animation and Virtual Worlds, 18(4–5):i–iv, September/ December 2007. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2008:Ea

[212] Nadia Magnenat-Thalmann, Daniel Thalmann, Gurminder Singh, and Georges Papagiannakis. Editorial. Computer Animation and Virtual Worlds, 19(1):1–2, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Papagiannakis:2008:SMW

[213] George Papagiannakis, Gurminder Singh, and Nadia Magnenat-Thalmann. A survey of mobile and wireless technologies for augmented reality systems. Computer Animation and Virtual Worlds, 19(1):3–22, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Barakonyi:2008:ARA

[214] István Barakonyi and Dieter Schmalstieg. Augmented reality agents for user interface adaptation. Computer Animation and Virtual Worlds, 19 (1):23–35, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ueki:2008:FPA

[215] Atsuro Ueki and Masa Inakage. Facilitating public awareness — the design of information distribution for communication in future workplaces. Computer Animation and Virtual Worlds, 19(1):37–47, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Matsumoto:2008:EWE

[216] Takashi Matsumoto, Sho Hashimoto, and Naohito Okude. The embodied Web: embodied Web-services interaction with an umbrella for augmented city experiences. Computer Animation and Virtual Worlds, 19(1):49–66, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mansa:2008:ACS

[217] Ignacio Mansa, Aiert Amundarain, Luis Matey, and Alejandro García-Alonso. Analysis of coherence strategies for stereo occlusion culling. Computer Animation and Virtual Worlds, 19 (1):67–77, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2008:CPS

[218] Anonymous. Call for papers: Special issue on medical applications of computer graphics. Computer Animation and Virtual Worlds, 19(1):78, February 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kallmann:2008:AIK

[219] Marcelo Kallmann. Analytical inverse kinematics with body posture control. Computer Animation and Virtual Worlds, 19(2):79–91, May 2008. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2008:IBM

[220] Man Hee Lee and In Kyu Park. Imagebased modeling of 3D objects with curved surfaces. Computer Animation and Virtual Worlds, 19(2):93–109, May 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vidal:2008:SUG

[221] F. P. Vidal, N. W. John, A. E. Healey, and D. A. Gould. Simulation of ultrasound guided needle puncture using patient specific data with 3D textures and volume haptics. Computer Animation and Virtual Worlds, 19(2):111–127, May 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Han:2008:SIM

[222] Seunghyun Han, Mingyu Lim, Dongman Lee, and Soon J. Hyun. A scalable interest management scheme for distributed virtual environments. Computer Animation and Virtual Worlds, 19(2):129–149, May 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Maciel:2008:EDP

[223] Anderson Maciel and Suvranu De. An efficient dynamic point algorithm for line-based collision detection in real time virtual environments involving haptics. Computer Animation and Virtual Worlds, 19(2):151–163, May 2008. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2008:Eb

[224] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 19(2):i–ii, May 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kondo:2008:DAE

[225] Ryo Kondo and Ken Anjyo. Directable animation of elastic bodies with pointconstraints. Computer Animation and Virtual Worlds, 19(3–4):165–173, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2008:ESK

[226] Byungkuk Choi, Mi You, and Junyong Noh. Extended spatial keyframing for complex character animation. Computer Animation and Virtual Worlds, 19 (3–4):175–188, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lyard:2008:MAB

[227] Etienne Lyard and Nadia Magnenat-Thalmann. Motion adaptation based on character shape. Computer Animation and Virtual Worlds, 19(3–4):189–198, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2008:SSC

[228] Gengdai Liu, Zhigeng Pan, and Zuoyan Lin. Style subspaces for character animation. Computer Animation and Virtual Worlds, 19(3–4):199–209, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tang:2008:EHP

[229] Jeff K. T. Tang, Howard Leung, Taku Komura, and Hubert P. H. Shum. Emulating human perception of motion similarity. Computer Animation and Virtual Worlds, 19(3–4):211–221, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2008:VDD

[230] Sujeong Kim, Stephane Redon, and Young J. Kim. View-dependent dynamics of articulated bodies. Computer Animation and Virtual Worlds, 19 (3–4):223–233, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2008:FAO

[231] Xuecheng Liu, Tianlu Mao, Shihong Xia, Yong Yu, and Zhaoqi Wang. Facial animation by optimized blendshapes from motion capture data. Computer Animation and Virtual Worlds, 19(3– 4):235–245, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kawamoto:2008:ELS

[232] Shin ichi Kawamoto, Tatsuo Yotsukura, Ken Anjyo, and Satoshi Nakamura. Efficient lip-synch tool for 3D cartoon animation. Computer Animation and Virtual Worlds, 19(3–4):247–257, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Arellano:2008:GVE

[233] Diana Arellano, Javier Varona, and Francisco J. Perales. Generation and visualization of emotional states in virtual characters. Computer Animation and Virtual Worlds, 19(3–4):259–270, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Luo:2008:ABH

[234] Linbo Luo, Suiping Zhou, Wentong Cai, Malcolm Yoke Hean Low, Feng Tian, Yongwei Wang, Xian Xiao, and Dan Chen. Agent-based human behavior modeling for crowd simulation. Computer Animation and Virtual Worlds, 19 (3–4):271–281, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Karamouzas:2008:AVP

[235] Ioannis Karamouzas and Mark H. Overmars. Adding variation to path planning. Computer Animation and Virtual Worlds, 19(3–4):283–293, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Silveira:2008:MCG

[236] Renato Silveira, Edson Prestes, and Luciana P. Nedel. Managing coherent groups. Computer Animation and Virtual Worlds, 19(3–4):295–305, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Haciomeroglu:2008:DPL

[237] M. Haciomeroglu, R. G. Laycock, and A. M. Day. Dynamically populating large urban environments with ambient virtual humans. Computer Animation and Virtual Worlds, 19(3–4):307–317, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xu:2008:SCF

[238] Jiayi Xu, Xiaogang Jin, Yizhou Yu, Tian Shen, and Mingdong Zhou. Shapeconstrained flock animation. Computer Animation and Virtual Worlds, 19 (3–4):319–330, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2008:SAP

[239] Andrew Park and Tom Calvert. A social agent pedestrian model. Computer Animation and Virtual Worlds, 19 (3–4):331–340, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiao:2008:VCS

[240] Chunxia Xiao, Shu Liu, Hongbo Fu, Chengchun Lin, Chengfang Song, Zhiyong Huang, Fazhi He, and Qunsheng Peng. Video completion and synthesis. Computer Animation and Virtual Worlds, 19(3–4):341–353, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhuang:2008:PAC

[241] Yueting Zhuang, Jun Yu, Jun Xiao, and Cheng Chen. Perspective-aware cartoon clips synthesis. Computer Animation and Virtual Worlds, 19(3–4):355–364, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

VanLaerhoven:2008:HPA

[242] Tom Van Laerhoven, Fabian Di Fiore, and Frank Van Reeth. Hand-painted animation with intelligent brushes. Computer Animation and Virtual Worlds, 19 (3–4):365–374, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2008:MIB

[243] Jinhui Yu, Jing Liao, and John Patterson. Modeling the interaction between objects and cartoon water. Computer Animation and Virtual Worlds, 19 (3–4):375–385, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xu:2008:MRP

[244] Leilei Xu, Hanqiu Sun, and Lifeng Wang. Multi-resolution parametric synthesis of manipulative dynamic textures. Computer Animation and Virtual Worlds, 19(3–4):387–398, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zou:2008:SMS

[245] Guangyu Zou, Jing Hua, Ming Dong, and Hong Qin. Surface matching with salient keypoints in geodesic scale space. Computer Animation and Virtual Worlds, 19(3–4):399–410, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2008:SDB

[246] Yanzhen Wang, Kai Xu, Yueshan Xiong, and Zhi-Quan Cheng. 2D shape deformation based on rigid square matching. Computer Animation and Virtual Worlds, 19(3–4):411–420, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Moss:2008:CBM

[247] William Moss, Ming C. Lin, and Dinesh Manocha. Constraint-based motion synthesis for deformable models. Computer Animation and Virtual Worlds, 19 (3–4):421–431, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

You:2008:DSD

[248] L. H. You, Xiaosong Yang, and Jian J.

Zhang. Dynamic skin deformation with characteristic curves. Computer Animation and Virtual Worlds, 19(3– 4):433–444, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Amor:2008:GSL

[249] Heni Ben Amor, Guido Heumer, Bernhard Jung, and Arnd Vitzthum. Grasp synthesis from low-dimensional probabilistic grasp models. Computer Animation and Virtual Worlds, 19(3– 4):445–454, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2008:UHI

[250] Jinho Park, Younghwi Kim, Daehyeon Wi, Nahyup Kang, Sung Yong Shin, and Junyong Noh. A unified handling of immiscible and miscible fluids. Computer Animation and Virtual Worlds, 19 (3–4):455–467, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hong:2008:DCM

[251] Jeong-Mo Hong, Jong-Chul Yoon, and Chang-Hun Kim. Divergenceconstrained moving least squares for fluid simulation. Computer Animation and Virtual Worlds, 19(3–4):469–477, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oh:2008:PFM

[252] SeungWoo Oh, Junyong Noh, and Kwangyun Wohn. A physically faithful multigrid method for fast cloth simulation. Computer Animation and Virtual Worlds, 19(3–4):479–492, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2008:SBD

[253] Chuan Zhou, Xiaogang Jin, and Charlie C. L. Wang. Shear buckling and dynamic bending in cloth simulation. Computer Animation and Virtual Worlds, 19(3–4):493–503, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ting:2008:STC

[254] Shang-Ping Ting and Suiping Zhou. Snap: a time critical decision-making framework for MOUT simulations. Computer Animation and Virtual Worlds, 19(3–4):505–514, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2008:VSB

[255] Youngjun Kim, Kunwoo Lee, and Wontae Kim. 3D virtual simulator for breast plastic surgery. Computer Animation and Virtual Worlds, 19(3–4):515–526, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2008:PCA

[256] Kyoungju Park. Physiologically correct animation of the heart. Computer Animation and Virtual Worlds, 19 (3–4):527–535, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2008:E

[257] Gerard J. Kim, Hong Qin, and Nadia Magnenat-Thalmann. Editorial. Computer Animation and Virtual Worlds, 19(3–4):i–iv, ???? 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Rueda:2008:CSA

[258] S. Rueda, P. Morillo, and J. M. Orduña.

A comparative study of awareness methods for peer-to-peer distributed virtual environments. Computer Animation and Virtual Worlds, 19(5):537–552, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2008:GMN

[259] Fan Zhang, Hanqiu Sun, Chong Zhao, and Lifeng Wang. Generalized minimum-norm perspective shadow maps. Computer Animation and Virtual Worlds, 19(5):553–567, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2008:PMU

[260] Liying Wang, Wei Hua, and Hujun Bao.

Procedural modeling of urban zone by optimization. Computer Animation and Virtual Worlds, 19(5):569–578, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gillies:2008:RLB

[261] M. Gillies, X. Pan, M. Slater, and J. Shawe-Taylor. Responsive listening behavior. Computer Animation and Virtual Worlds, 19(5):579–589, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Castellani:2008:RDC

[262] Umberto Castellani, Vincent Gay-Bellile, and Adrien Bartoli. Robust deformation capture from temporal range data for surface rendering. Computer Animation and Virtual Worlds, 19(5): 591–603, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anderson:2008:LLI

[263] James N. Anderson, Nancie Davidson, Hazel Morton, and Mervyn A. Jack. Language learning with interactive virtual agent scenarios and speech recognition: Lessons learned. Computer Animation and Virtual Worlds, 19(5): 605–619, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2008:Ec

[264] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 19(5):iii–iv, December 2008. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2009:CAG

[265] Taek Hee Lee and Yeoung Gil Shin. Coherence aware GPU-based ray casting for virtual colonoscopy. Computer Animation and Virtual Worlds, 20(1): 1–9, January/February 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lim:2009:GBI

[266] Sukhyun Lim, Koojoo Kwon, and Byeong-Seok Shin. GPU-based interactive visualization framework for ultrasound datasets. Computer Animation and Virtual Worlds, 20(1):11– 23, January/February 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cakmak:2009:HVS

[267] H. Çakmak, H. Maaß, C. Trantakis, G. Strauß, E. Nowatius, and U. Kühnapfel. Haptic ventriculostomy simulation in a grid environment. Computer Animation and Virtual Worlds, 20 (1):25–38, January/February 2009. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bayona:2009:NAM

[268] Sofía Bayona, José Manuel Fernández-Arroyo, Pilar Bayona, and Luis Pastor. A new assessment methodology for virtual reality surgical simulators. Computer Animation and Virtual Worlds, 20 (1):39–52, January/February 2009. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Assassi:2009:MAS

[269] Lazhari Assassi, Caecilia Charbonnier, Jérôme Schmid, Pascal Volino, and Nadia Magnenat-Thalmann. From MRI to anatomical simulation of the hip joint. Computer Animation and Virtual Worlds, 20(1):53–66, January/February 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jin:2009:IGM

[270] Ge Jin, Nakhoon Baek, James K. Hahn, Steven Bielamowicz, Rajat Mittal, and Raymond Walsh. Image guided medialization laryngoplasty. Computer Animation and Virtual Worlds, 20(1):67– 77, January/February 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hahn:2009:SIP

[271] James Kwangjune Hahn and Nadia Magnenat Thalmann. Special issue “Physiological Human”. Computer Animation and Virtual Worlds, 20(1): iii, January/February 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nijholt:2009:E

[272] Anton Nijholt, Arjan Egges, and Herwin van Welbergen. Editorial. Computer Animation and Virtual Worlds, 20(2–3):i–iv, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baxter:2009:WMA

[273] William Baxter, Pascal Barla, and Ken Anjyo. N-way morphing for 2D animation. Computer Animation and Virtual Worlds, 20(2–3):79–87, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liang:2009:PDM

[274] Xiubo Liang, Qilei Li, Xiang Zhang, Shun Zhang, and Weidong Geng. Performance-driven motion choreographing with accelerometers. Computer Animation and Virtual Worlds, 20(2–3):89– 99, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yao:2009:CQS

[275] Chih-Yuan Yao, Hung-Kuo Chu, Tao Ju, and Tong-Yee Lee. Compatible quadrangulation by sketching. Computer Animation and Virtual Worlds, 20(2–3):101–109, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Grillon:2009:SGA

[276] Helena Grillon and Daniel Thalmann.

Simulating gaze attention behaviors for crowds. Computer Animation and Virtual Worlds, 20(2–3):111–119, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pan:2009:ARA

[277] JunJun Pan, Xiaosong Yang, Xin Xie, Philip Willis, and Jian J. Zhang. Automatic rigging for animation characters with 3D silhouette. Computer Animation and Virtual Worlds, 20(2–3):121–131, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2009:RTD

[278] Jin Huang, Hanqiu Sun, Kun Zhou, and Hujun Bao. Real-time dynamics for geometric textures in shell. Computer Animation and Virtual Worlds, 20(2–3):133–141, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Todo:2009:SLC

[279] Hideki Todo, Ken Anjyo, and Takeo Igarashi. Stylized lighting for cartoon shader. Computer Animation and Virtual Worlds, 20(2–3):143–152, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kasap:2009:FED

[280] Mustafa Kasap, Parag Chaudhuri, and Nadia Magnenat-Thalmann. Fast EMGdata driven skin deformation. Computer Animation and Virtual Worlds, 20(2–3):153–161, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Burrell:2009:ART

[281] Tim Burrell, Dirk Arnold, and Stephen Brooks. Advected river textures. Computer Animation and Virtual Worlds, 20(2–3):163–173, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2009:SMT

[282] Wenwu Yang and Jieqing Feng. 2D shape manipulation via topology-aware rigid grid. Computer Animation and Virtual Worlds, 20(2–3):175–184, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2009:FC

[283] Hanli Zhao, Ran Fan, Charlie C. L. Wang, Xiaogang Jin, and Yuwei Meng. Fireworks controller. Computer Animation and Virtual Worlds, 20(2–3):185– 194, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

vanVugt:2009:IEE

[284] H. C. van Vugt, J. F. Hoorn, and E. A. Konijn. Interactive engagement with embodied agents: an empirically validated framework. Computer Animation and Virtual Worlds, 20(2–3):195–204, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sheng:2009:FAS

[285] Bin Sheng, Hanqiu Sun, Gang Yang, and Enhua Wu. Furstyling on angle-split shell textures. Computer Animation and Virtual Worlds, 20(2–3):205–213, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oh:2009:IBR

[286] Seungtaik Oh, Younghee Kim, and Byung-Seok Roh. Impulse-based rigid body interaction in SPH. Computer Animation and Virtual Worlds, 20 (2–3):215–224, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liang:2009:CMS

[287] Zhang Liang, Jun Xiao, Yueting Zhuang, and Cheng Chen. Competitive motion synthesis based on hybrid control. Computer Animation and Virtual Worlds, 20(2–3):225–235, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ye:2009:CML

[288] Kejia Ye, Baocai Yin, and Lichun Wang.

CSLML: a markup language for expressive Chinese sign language synthesis. Computer Animation and Virtual Worlds, 20(2–3):237–245, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jang:2009:CKA

[289] Yoojin Jang and Insung Ihm. Chemical kinetics-assisted, path-based smoke simulation. Computer Animation and Virtual Worlds, 20(2–3):247–256, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liew:2009:DCC

[290] Pak-San Liew, Ching-Ling Chin, and Zhiyong Huang. Development of a computational cognitive architecture for intelligent virtual character. Computer Animation and Virtual Worlds, 20(2–3):257–266, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2009:PPD

[291] Cheng Chen, Yueting Zhuang, Jun Xiao, and Zhang Liang. Perceptual 3D pose distance estimation by boosting relational geometric features. Computer Animation and Virtual Worlds, 20(2–3):267–277, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2009:PDM

[292] Shaojun Hu, Tadahiro Fujimoto, and Norishige Chiba. Pseudo-dynamics model of a cantilever beam for animating flexible leaves and branches in wind field. Computer Animation and Virtual Worlds, 20(2–3):279–287, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2009:SDF

[293] Jeong-Sik Kim and Soo-Mi Choi. Symmetric deformation of 3D face scans using facial features and curvatures. Computer Animation and Virtual Worlds, 20(2–3):289–300, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2009:USE

[294] Yong Zhao, Xinguo Liu, Chunxia Xiao, and Qunsheng Peng. A unified shape editing framework based on tetrahedral control mesh. Computer Animation and Virtual Worlds, 20(2–3):301–310, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bao:2009:PCS

[295] Kai Bao, Hui Zhang, Lili Zheng, and Enhua Wu. Pressure corrected SPH for fluid animation. Computer Animation and Virtual Worlds, 20(2–3):311–320, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2009:PCS

[296] Jieun Lee, Myung-Soo Kim, and Seung-Hyun Yoon. Patches: character skinning with local deformation layer. Computer Animation and Virtual Worlds, 20(2–3):321–331, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2009:FSS

[297] Xiaosong Yang, Richard Southern, and Jian Jun Zhang. Fast simulation of skin sliding. Computer Animation and Virtual Worlds, 20(2–3):333–342, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mamou:2009:TLC

[298] Khaled Mamou, Titus Zaharia, and Françoise Prêteux. TFAN: a low complexity 3D mesh compression algorithm. Computer Animation and Virtual Worlds, 20(2–3):343–354, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gissler:2009:TCC

[299] Marc Gissler, Ruediger Schmedding, and Matthias Teschner. Time-critical collision handling for deformable modeling. Computer Animation and Virtual Worlds, 20(2–3):355–364, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ohta:2009:DFU

[300] Makoto Ohta, Yoshihiro Kanamori, and Tomoyuki Nishita. Deformation and fracturing using adaptive shape matching with stiffness adjustment. Computer Animation and Virtual Worlds, 20(2–3):365–373, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2009:PMA

[301] Jae Woo Kim, Hesham Fouad, John L. Sibert, and James K. Hahn. Perceptually motivated automatic dance motion generation for music. Computer Animation and Virtual Worlds, 20(2–3):375–384, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shum:2009:AMG

[302] Hubert P. H. Shum, Taku Komura, and Pranjul Yadav. Angular momentum guided motion concatenation. Computer Animation and Virtual Worlds, 20(2–3):385–394, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sugisaki:2009:ISA

[303] Eiji Sugisaki, Hock Soon Seah, Feng Tian, and Shigeo Morishima. Interactive shadowing for 2D Anime. Computer Animation and Virtual Worlds, 20(2–3):395–404, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Beato:2009:ICK

[304] Nicholas Beato, Yunjun Zhang, Mark Colbert, Kazumasa Yamazawa, and Charles E. Hughes. Interactive chroma keying for mixed reality. Computer Animation and Virtual Worlds, 20 (2–3):405–415, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yan:2009:RTF

[305] He Yan, Zhangye Wang, Jian He, Xi Chen, Changbo Wang, and Qunsheng Peng. Real-time fluid simulation with adaptive SPH. Computer Animation and Virtual Worlds, 20(2–3):417–426, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ting:2009:DDC

[306] Shang-Ping Ting and Suiping Zhou.

Dealing with dynamic changes in time critical decision-making for MOUT simulations. Computer Animation and Virtual Worlds, 20(2–3):427–436, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cramer:2009:GMH

[307] Henriette Cramer, Nicander Kemper, Alia Amin, Bob Wielinga, and Vanessa Evers. ‘give me a hug’: the effects of touch and autonomy on people’s responses to embodied social agents. Computer Animation and Virtual Worlds, 20(2–3):437–445, June 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vasa:2009:CCS

[308] Libor Váša and Václav Skala. Combined compression and simplification of dynamic 3D meshes. Computer Animation and Virtual Worlds, 20(4):447–456, July 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Farouki:2009:SCO

[309] Rida T. Farouki and Carlotta Giannelli. Spatial camera orientation control by rotation-minimizing directed frames. Computer Animation and Virtual Worlds, 20(4):457–472, July 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baiget:2009:GAV

[310] Pau Baiget, Carles Fernández, Xavier Roca, and Jordi Gonzàlez. Generation of augmented video sequences combining behavioral animation and multi-object tracking. Computer Animation and Virtual Worlds, 20(4):473–489, July 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2009:E

[311] Nadia Magnenat-Thalmann. Editorial. Computer Animation and Virtual Worlds, 20(4):iii, July 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Multon:2009:IAV

[312] Franck Multon, Richard Kulpa, Ludovic Hoyet, and Taku Komura. Interactive animation of virtual humans based on motion capture data. Computer Animation and Virtual Worlds, 20(5–6):491– 500, September/December 2009. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Quax:2009:EDE

[313] Peter Quax, Fabian Di Fiore, Wim Lamotte, and Frank Van Reeth. Efficient distribution of emotion-related data through a networked virtual environment architecture. Computer Animation and Virtual Worlds, 20(5–6):501– 510, September/December 2009. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yoshida:2009:MAH

[314] Eiichi Yoshida, Jean-Paul Laumond, Claudia Esteves, Oussama Kanoun, Anthony Mallet, Takeshi Sakaguchi, and Kazuhito Yokoi. Motion autonomy for humanoids: experiments on HRP-2 No. 14. Computer Animation and Virtual Worlds, 20(5–6):511–522, September/December 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiao:2009:RTA

[315] Zhidong Xiao, Hammadi Nait-Charif, and Jian J. Zhang. Real time automatic skeleton and motion estimation for character animation. Computer Animation and Virtual Worlds, 20(5–6):523–531, September/December 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Singh:2009:SBS

[316] Shawn Singh, Mubbasir Kapadia, Petros Faloutsos, and Glenn Reinman. Steer-Bench: a benchmark suite for evaluating steering behaviors. Computer Animation and Virtual Worlds, 20(5–6):533– 548, September/December 2009. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Egges:2009:GEI

[317] Arjan Egges, Arno Kamphuis, and Mark Overmars. Guest editors’ introduction: Special issue on gaming. Computer Animation and Virtual Worlds, 20(5–6):iii– iv, September/December 2009. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2010:Ea

[318] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 21(1): iii, January/February 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Deng:2010:MFF

[319] Qingqiong Deng, Xiaopeng Zhang, Gang Yang, and Marc Jaeger. Multiresolution foliage for forest rendering. Computer Animation and Virtual Worlds, 21(1): 1–23, January/February 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hwang:2010:AHR

[320] Sun-Uk Hwang, Beom-Chan Lee, Jeha Ryu, Kwan H. Lee, and Yong-Gu Lee. Adaptive haptic rendering for timevarying haptic and video frame rates in multi-modal interactions. Computer Animation and Virtual Worlds, 21(1):25– 38, January/February 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Stoiber:2010:FAR

[321] Nicolas Stoiber, Renaud Seguier, and Gaspard Breton. Facial animation retargeting and control based on a human appearance space. Computer Animation and Virtual Worlds, 21(1):39– 54, January/February 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lin:2010:IBD

[322] I-Chen Lin, Wen-Hsing Chang, Yung-Sheng Lo, Jen-Yu Peng, and Chan-Yu Lin. Image-based detail reconstruction of non-Lambertian surfaces. Computer Animation and Virtual Worlds, 21(1): 55–68, January/February 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2010:EIV

[323] Daniel Thalmann and Zhiyong Huang.

Editorial issue VRCAI’08. Computer Animation and Virtual Worlds, 21 (2):i–ii, March/April 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2010:ACA

[324] Hanhoon Park, Jihyun Oh, Byung-Kuk Seo, and Jong-Il Park. Automatic confidence adjustment of visual cues in model-based camera tracking. Computer Animation and Virtual Worlds, 21 (2):69–79, March/April 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Fei:2010:PPA

[325] Guangzheng Fei, Won-Sook Lee, Zijun Xin, Huikai Dong, and Chris Joslin. PASCAL: physics augmented space canvases for animating locomotion. Computer Animation and Virtual Worlds, 21 (2):81–89, March/April 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2010:TTM

[326] Yimin Wang and Jianmin Zheng. Tubular triangular mesh parameterization and applications. Computer Animation and Virtual Worlds, 21(2):91–102, March/April 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2010:DAS

[327] J. Zhang, S. K. Ong, and A. Y. C. Nee. Development of an AR system achieving in situ machining simulation on a 3-axis CNC machine. Computer Animation and Virtual Worlds, 21(2): 103–115, March/April 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Manders:2010:GCS

[328] Corey Manders, Farzam Farbiz, Tang Ka Yin, Yuan Miaolong, and Chua Gim Guan. A gesture control system for intuitive 3D interaction with virtual objects. Computer Animation and Virtual Worlds, 21(2):117–129, March/ April 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Donikian:2010:E

[329] Stéphane Donikian, Elisabeth André, Shi-Min Hu, and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 21(3–4):131–135, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pan:2010:HAS

[330] Jia Pan, Liangjun Zhang, Ming C. Lin, and Dinesh Manocha. A hybrid approach for simulating human motion in constrained environments. Computer Animation and Virtual Worlds, 21(3–4):137–149, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2010:RTH

[331] Yin Yang, Guodong Rong, Luis Torres, and Xiaohu Guo. Real-time hybrid solid simulation: spectral unification of deformable and rigid materials. Computer Animation and Virtual Worlds, 21(3–4):151–159, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Steptoe:2010:EKV

[332] William Steptoe, Oyewole Oyekoya, and Anthony Steed. Eyelid kinematics for virtual characters. Computer Animation and Virtual Worlds, 21(3–4):161–171, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sussmuth:2010:AT

[333] Jochen Süßmuth, Michael Zollhöfer, and Günther Greiner. Animation transplantation. Computer Animation and Virtual Worlds, 21(3–4):173–182, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pronost:2010:CJD

[334] Nicolas Pronost, Anders Sandholm, and Daniel Thalmann. Correlative joint definition for motion analysis and animation. Computer Animation and Virtual Worlds, 21(3–4):183–192, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2010:ACF

[335] Xiang Li, Jun Xu, Yangchun Ren, and Weidong Geng. Animating cartoon faces by multi-view drawings. Computer Animation and Virtual Worlds, 21(3–4):193–201, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sloan:2010:CEF

[336] Robin J. S. Sloan, Brian Robinson, Ken Scott-Brown, Fhionna Moore, and Malcolm Cook. Choreographing emotional facial expressions. Computer Animation and Virtual Worlds, 21(3–4):203–213, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vanhala:2010:VPF

[337] Toni Vanhala, Veikko Surakka, Harri Siirtola, Kari-Jouko Räihä, Benoît Morel, and Laurent Ach. Virtual proximity and facial expressions of computer agents regulate human emotions and attention. Computer Animation and Virtual Worlds, 21(3–4):215–224, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

deMelo:2010:RTE

[338] Celso M. de Melo, Patrick Kenny, and Jonathan Gratch. Real-time expression of affect through respiration. Computer Animation and Virtual Worlds, 21(3–4):225–234, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2010:SFR

[339] Jun Wu, Yuen-Shan Leung, Charlie C. L. Wang, Dangxiao Wang, and Yuru Zhang. Smooth force rendering on coarse polygonal meshes. Computer Animation and Virtual Worlds, 21(3–4):235–244, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Garcia:2010:PVG

[340] Marcos García, Miguel Anguel Otaduy, and Carol O’Sullivan. Perceptually validated global/local deformations. Computer Animation and Virtual Worlds, 21(3–4):245–254, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Na:2010:WLS

[341] Kyung-Gun Na and Moon-Ryul Jung.

Weighted local shape blending for facial motion retargetting. Computer Animation and Virtual Worlds, 21 (3–4):255–265, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Schuerman:2010:SAA

[342] Matthew Schuerman, Shawn Singh, Mubbasir Kapadia, and Petros Faloutsos. Situation agents: agent-based externalized steering logic. Computer Animation and Virtual Worlds, 21 (3–4):267–276, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2010:TGB

[343] Jun Yu, Dongquan Liu, and Hock Soon Seah. Transductive graph based cartoon synthesis. Computer Animation and Virtual Worlds, 21(3–4):277–288, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiao:2010:RNB

[344] Xuezhong Xiao, Hua Huang, and Lizhuang Ma. RBF network-based temporal color morphing. Computer Animation and Virtual Worlds, 21(3–4):289– 296, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

You:2010:SMU

[345] L. H. You, Xiaosong Yang, X. Y. You, Xiaogang Jin, and Jian J. Zhang. Shape manipulation using physically based wire deformations. Computer Animation and Virtual Worlds, 21(3–4):297–309, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2010:RTS

[346] Jun Lee, DongKyum Kim, HyungSeok Kim, Carloa Henzel, Jee-In Kim, and MinGyu Lim. Real-time fur simulation and rendering. Computer Animation and Virtual Worlds, 21(3–4):311–320, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2010:NAO

[347] Yanli Liu, Xueying Qin, Guanyu Xing, and Qunsheng Peng. A new approach to outdoor illumination estimation based on statistical analysis for augmented reality. Computer Animation and Virtual Worlds, 21(3–4):321–330, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2010:IVL

[348] Sheng hui Liao, Ruo feng Tong, Jian-Ping Geng, and Min Tang. Inhomogeneous volumetric Laplacian deformation for rhinoplasty planning and simulation system. Computer Animation and Virtual Worlds, 21(3–4):331–341, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gao:2010:ACA

[349] Yujian Gao, Qinping Zhao, Aimin Hao, T. M. Sezgin, and N. A. Dodgson. Automatic construction of 3D animatable facial avatars. Computer Animation and Virtual Worlds, 21(3–4):343–354, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Feng:2010:RTI

[350] Wei Feng, Hongxin Zhang, Jin Huang, Caoyu Wang, and Hujun Bao. Repairing topological inconsistency of mesh sequences. Computer Animation and Virtual Worlds, 21(3–4):355–364, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2010:SPP

[351] Youngeun Lee and Young J. Kim. Simple and parallel proximity algorithms for general polygonal models. Computer Animation and Virtual Worlds, 21(3–4):365–374, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Seo:2010:RT

[352] Jaewoo Seo, Yeongho Seol, Daehyeon Wi, Younghui Kim, and Junyong Noh. Rigging transfer. Computer Animation and Virtual Worlds, 21(3–4):375–386, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2010:STP

[353] Nan Hu, Suiping Zhou, Zhongke Wu, Mingquan Zhou, and Benjamin Eng Keong Cho. Spatial-temporal patterns and pedestrian simulation. Computer Animation and Virtual Worlds, 21(3–4):387–399, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bao:2010:VFB

[354] Kai Bao, Xiaolong Wu, Hui Zhang, and Enhua Wu. Volume fraction based miscible and immiscible fluid animation. Computer Animation and Virtual Worlds, 21(3–4):401–410, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tang:2010:SSW

[355] Chen Tang, Sheng Li, Guoping Wang, and Yutong Zang. Stable stylized wireframe rendering. Computer Animation and Virtual Worlds, 21(3–4):411–421, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jeon:2010:SIE

[356] Jaewoong Jeon, Hyunho Jang, Soon-Bum Lim, and Yoon-Chul Choy. A sketch interface to empower novices to create 3D animations. Computer Animation and Virtual Worlds, 21 (3–4):423–432, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

vanBasten:2010:SSE

[357] B. J. H. van Basten, P. W. A. M. Peeters, and A. Egges. The step space: example-based footprint-driven motion synthesis. Computer Animation and Virtual Worlds, 21(3–4):433–441, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Courty:2010:CSS

[358] N. Courty and A. Cuzol. Conditional stochastic simulation for character animation. Computer Animation and Virtual Worlds, 21(3–4):443–452, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gerdelan:2010:GFS

[359] Anton Gerdelan and Carol O’Sullivan.

A genetic-fuzzy system for optimising agent steering. Computer Animation and Virtual Worlds, 21(3–4):453–461, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oh:2010:SVB

[360] Sejin Oh, Woonhyuk Baek, and Woontack Woo. Synthetic vision-based perceptual attention for augmented reality agents. Computer Animation and Virtual Worlds, 21(3–4):463–472, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kang:2010:VHE

[361] Sin-Hwa Kang and Jonathan Gratch.

Virtual humans elicit socially anxious interactants’ verbal self-disclosure. Computer Animation and Virtual Worlds, 21(3–4):473–482, May 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2010:Eb

[362] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 21(5):483– 484, September/October 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2010:HCM

[363] Hui Chen, Hanqiu Sun, and Xiaogang Jin. Haptic-constraint modeling based on interactive metaballs. Computer Animation and Virtual Worlds, 21(5):485– 497, September/October 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ma:2010:IVG

[364] Chunyong Ma, Ge Chen, Yong Han, Yongyang Qi, and Yong Chen. An integrated VR–GIS navigation platform for city/region simulation. Computer Animation and Virtual Worlds, 21(5):499– 507, September/October 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2010:HOV

[365] Jiejie Zhu, Zhigeng Pan, Chao Sun, and Wenzhi Chen. Handling occlusions in video-based augmented reality using depth information. Computer Animation and Virtual Worlds, 21(5):509– 521, September/October 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Seo:2010:BGH

[366] Hyewon Seo, Joon-Won Bang, Ji-Man Park, and Soo-Hyun Jeon. 3D billiards game with haptic interface. Computer Animation and Virtual Worlds, 21(5): 523–530, September/October 2010. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2010:AIS

[367] Soo Jeong Kim, Kyung Jun Gil, HyungSeok Kim, Sang Boem Lim, and Jee in Kim. Adaptive interactions in shared virtual environments for heterogeneous devices. Computer Animation and Virtual Worlds, 21(5):531– 543, September/October 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2010:Ec

[368] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 21(6):545, November/December 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hwang:2010:PMP

[369] Jane Hwang and Gerard J. Kim. Provision and maintenance of presence and immersion in hand-held virtual reality through motion based interaction. Computer Animation and Virtual Worlds, 21 (6):547–559, November/December 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Spillmann:2010:IER

[370] Jonas Spillmann and Matthias Harders. Inextensible elastic rods with torsional friction based on Lagrange multipliers. Computer Animation and Virtual Worlds, 21(6):561–572, November/ December 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Buche:2010:FCM

[371] C. Buche, P. Chevaillier, A. Nédélec, M. Parenthoën, and J. Tisseau. Fuzzy cognitive maps for the simulation of individual adaptive behaviors. Computer Animation and Virtual Worlds, 21 (6):573–587, November/December 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sun:2010:PBR

[372] Hongquan Sun and Jiqing Han. Particlebased realistic simulation of fluid– solid interaction. Computer Animation and Virtual Worlds, 21(6):589–595, November/December 2010. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2011:IIa

[373] Anonymous. Issue information. Computer Animation and Virtual Worlds, 22 (1):i–ii, January/February 2011. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2011:Ea

[374] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 22(1): 1, January/February 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Halit:2011:MMS

[375] Cihan Halit and Tolga Capin. Multiscale motion saliency for keyframe extraction from motion capture sequences. Computer Animation and Virtual Worlds, 22 (1):3–14, January/February 2011. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

He:2011:FSD

[376] Jun He, Caiming Zhang, Yu Wei, and Weitao Li. Feature sensitive deformation for triangular mesh models. Computer Animation and Virtual Worlds, 22(1): 15–25, January/February 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vasilakis:2011:GRS

[377] Andreas A. Vasilakis and Ioannis Fudos. GPU rigid skinning based on a refined skeletonization method. Computer Animation and Virtual Worlds, 22(1):27– 46, January/February 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2011:ELT

[378] Shiqi Li and Chi Xu. Efficient lookup table based camera pose estimation for augmented reality. Computer Animation and Virtual Worlds, 22(1):47– 58, January/February 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

vanBasten:2011:CPP

[379] B. J. H. van Basten, A. Egges, and R. Geraerts. Combining path planners and motion graphs. Computer Animation and Virtual Worlds, 22(1):59– 78, January/February 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2011:IIb

[380] Anonymous. Issue information. Computer Animation and Virtual Worlds, 22(2–3):i–ii, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pan:2011:E

[381] Zhigeng Pan, Nadia Magnenat-Thalmann, and Jinxiang Chai. Editorial. Computer Animation and Virtual Worlds, 22 (2–3):79–80, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2011:MOW

[382] Jing Liao, Jinhui Yu, and John Patterson. Modeling ocean waves and interaction between objects and ocean water for cartoon animation. Computer Animation and Virtual Worlds, 22(2– 3):81–89, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2011:FBV

[383] Qizhi Yu, Fabrice Neyret, and Anthony Steed. Feature-based vector simulation of water waves. Computer Animation and Virtual Worlds, 22(2– 3):91–98, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2011:ISF

[384] Chen Huang, Hanqiu Sun, Shiguang Liu, and Ping Li. Interactive soft-fabrics watering simulation on GPU. Computer Animation and Virtual Worlds, 22(2– 3):99–106, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

He:2011:RTS

[385] Shengfeng He, Hon-Cheng Wong, Wai-Man Pang, and Un-Hong Wong. Realtime smoke simulation with improved turbulence by spatial adaptive vorticity confinement. Computer Animation and Virtual Worlds, 22(2–3):107–114, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2011:MSB

[386] Ning Liu, Xiaowei He, Sheng Li, and Guoping Wang. Meshless simulation of brittle fracture. Computer Animation and Virtual Worlds, 22(2–3): 115–124, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2011:RFC

[387] Jian Zhu, Yuanzhang Chang, and Enhua Wu. Realistic, fast, and controllable simulation of solid combustion. Computer Animation and Virtual Worlds, 22(2– 3):125–132, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Buche:2011:STA

[388] Cédric Buche, Anne Jeannin-Girardon, and Pierre De Loor. Simulation theory and anticipation for interactive virtual character in an uncertain world. Computer Animation and Virtual Worlds, 22 (2–3):133–139, April/May 2011. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiang:2011:UUE

[389] Nan Xiang, Haiying Zhao, Xiaojian Zhou, Mingliang Xu, Abdennour El Rhalibi, and Yu Wu. UEGM: uncertain emotion generator under multistimulus. Computer Animation and Virtual Worlds, 22(2–3):141–149, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Singh:2011:FND

[390] Shawn Singh, Mubbasir Kapadia, Glenn Reinman, and Petros Faloutsos. Footstep navigation for dynamic crowds. Computer Animation and Virtual Worlds, 22(2–3):151–158, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2011:FLR

[391] Ko-Yun Liu, Wan-Chun Ma, Chun-Fa Chang, Chuan-Chang Wang, and Paul Debevec. A framework for locally retargeting and rendering facial performance. Computer Animation and Virtual Worlds, 22(2–3):159–167, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Dutreve:2011:EAR

[392] Ludovic Dutreve, Alexandre Meyer, and Sada Bouakaz. Easy acquisition and real-time animation of facial wrinkles. Computer Animation and Virtual Worlds, 22(2–3):169–176, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2011:MLS

[393] Chung-Yeon Lee, Sangyong Lee, and Seongah Chin. Multi-layer structural wound synthesis on 3D face. Computer Animation and Virtual Worlds, 22(2– 3):177–185, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Song:2011:CFR

[394] Jaewon Song, Byungkuk Choi, Yeongho Seol, and Junyong Noh. Characteristic facial retargeting. Computer Animation and Virtual Worlds, 22(2–3): 187–194, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zollhofer:2011:ARP

[395] Michael Zollhöfer, Michael Martinek, Günther Greiner, Marc Stamminger, and Jochen Süßmuth. Automatic reconstruction of personalized avatars from 3D face scans. Computer Animation and Virtual Worlds, 22(2–3):195–202, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tong:2011:BSH

[396] Jing Tong, Mingmin Zhang, Xueqin Xiang, Huaqing Shen, Hao Yan, and Zhengming Chen. 3D body scanning with hairstyle using one time-of-flight camera. Computer Animation and Virtual Worlds, 22(2–3):203–211, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pan:2011:ISL

[397] Jun Pan, Jian min Wang, Shun ting Cao, and Xiao nan Luo. Interactive sign language synthesis based on adaptive display resolution visibility for ubiquitous devices. Computer Animation and Virtual Worlds, 22(2–3):213–220, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiao:2011:PMM

[398] Jun Xiao, Yinfu Feng, and Wenyuan Hu. Predicting missing markers in human motion capture using l_{1}-sparse representation. Computer Animation and Virtual Worlds, 22(2–3):221–228, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Deng:2011:RTM

[399] Liqun Deng, Howard Leung, Naijie Gu, and Yang Yang. Real-time mocap dance recognition for an interactive dancing game. Computer Animation and Virtual Worlds, 22(2–3):229–237, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yeh:2011:ECP

[400] I-Cheng Yeh, Chao-Hung Lin, Hung-Jen Chien, and Tong-Yee Lee. Efficient camera path planning algorithm for human motion overview. Computer Animation and Virtual Worlds, 22(2– 3):239–250, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Si:2011:STW

[401] Weixin Si, Zhiyong Yuan, Xiangyun Liao, Zhaoliang Duan, Yihua Ding, and Jianhui Zhao. 3D soft tissue warping dynamics simulation based on force asynchronous diffusion model. Computer Animation and Virtual Worlds, 22(2– 3):251–259, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2011:NAH

[402] D. Huang, W. Tang, T. R. Wan, N. W. John, D. Gould, Y. Ding, and Y. Chen. A new approach to haptic rendering of guidewires for use in minimally invasive surgical simulation. Computer Animation and Virtual Worlds, 22(2– 3):261–268, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2011:PCS

[403] Sheng-Hui Liao, Yixiong Liang, Ling-Zhi Li, Bei-Ji Zou, Xing-Hao Zhu, and Wei Ai. Practical craniofacial surgery simulator based on GPU accelerated lattice shape matching. Computer Animation and Virtual Worlds, 22(2–3): 269–276, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Peng:2011:VES

[404] Yanjun Peng, Ruisheng Jia, Yuanhong Wang, and Mingmin Zhang. A virtual endoscopy system for virtual medicine. Computer Animation and Virtual Worlds, 22(2–3):277–284, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2011:GFS

[405] Lu Chen, Jin Huang, Hongxin Zhang, and Wei Hua. GPU-friendly shape interpolation based on trajectory warping. Computer Animation and Virtual Worlds, 22(2–3):285–294, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chao:2011:GBS

[406] Min-Wen Chao, Chao-Hung Lin, Chih-Chieh Chang, and Tong-Yee Lee. A graph-based shape matching scheme for 3D articulated objects. Computer Animation and Virtual Worlds, 22(2–3): 295–305, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2011:EWB

[407] Chong Zhao, Hanqiu Sun, and Kaihuai Qin. Efficient wavelet-based geometry compression. Computer Animation and Virtual Worlds, 22(2–3):307–315, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hwang:2011:SIG

[408] Huicheol Hwang, Kyehyun Kim, Roger Blanco i Ribera, and Junyong Noh. Stereoscopic image generation of background terrain scenes. Computer Animation and Virtual Worlds, 22(2–3): 317–323, April/May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

VanLaerhoven:2011:PGA

[409] Tom Van Laerhoven, Fabian Di Fiore, William Van Haevre, and Frank Van Reeth. Paint-on-glass animation: the fellowship of digital paint and artisanal control. Computer Animation and Virtual Worlds, 22(2–3):325–332, April/ May 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2011:IIc

[410] Anonymous. Issue information. Computer Animation and Virtual Worlds, 22(4):1–2, July/August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2011:EI

[411] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 22.4. Computer Animation and Virtual Worlds, 22(4):333, July/August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Meng:2011:IDC

[412] Min Meng, Lubin Fan, and Ligang Liu. iCutter: a direct cut-out tool for 3D shapes. Computer Animation and Virtual Worlds, 22(4):335–342, July/ August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Rodriguez:2011:CPM

[413] Inmaculada Rodriguez, Anna Puig, and Marc Esteva. Cross-platform management of intelligent objects behaviors in serious virtual environments. Computer Animation and Virtual Worlds, 22(4): 343–350, July/August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2011:HMG

[414] Gengdai Liu, Mingliang Xu, Zhigeng Pan, and Abdennour El Rhalibi. Human motion generation with multifactor models. Computer Animation and Virtual Worlds, 22(4):351–359, July/ August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kokkinara:2011:MSV

[415] Elena Kokkinara, Oyewole Oyekoya, and Anthony Steed. Modelling selective visual attention for autonomous virtual characters. Computer Animation and Virtual Worlds, 22(4):361–369, July/ August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gourret:2011:EIA

[416] Jean-Paul Gourret, Amir Hariri, and Philippe Liverneaux. Explicit and implicit animation with fuzzy constraints of a versatile multi-body system for virtual hand surgery. Computer Animation and Virtual Worlds, 22(4): 371–392, July/August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Slomp:2011:PRT

[417] Marcos Slomp, Matthew W. Johnson, Toru Tamaki, and Kazufumi Kaneda. Photorealistic real-time rendering of spherical raindrops with hierarchical reflective and refractive maps. Computer Animation and Virtual Worlds, 22(4): 393–404, July/August 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2011:IId

[418] Anonymous. Issue information. Computer Animation and Virtual Worlds, 22 (5):1–2, September/October 2011. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2011:E

[419] Anonymous. Editorial. Computer Animation and Virtual Worlds, 22(5): 405, September/October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Raunhardt:2011:ISF

[420] Daniel Raunhardt and Ronan Boulic. Immersive singularity-free full-body interactions with reduced marker set. Computer Animation and Virtual Worlds, 22(5):407–419, September/ October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Olivier:2011:SSM

[421] Anne-Hélène Olivier, Richard Kulpa, Julien Pettré, and Armel Crétual. A step-by-step modeling, analysis and annotation of locomotion. Computer Animation and Virtual Worlds, 22(5):421– 433, September/October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ho:2011:FSM

[422] Edmond S. L. Ho and Taku Komura. A finite state machine based on topology coordinates for wrestling games. Computer Animation and Virtual Worlds, 22 (5):435–443, September/October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Donikian:2011:SRW

[423] Stéphane Donikian and Paolo Petta. A survey of research work in computer science and cognitive science dedicated to the modeling of reactive human behaviors. Computer Animation and Virtual Worlds, 22(5):445–455, September/ October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chou:2011:CCA

[424] Yun-Feng Chou and Zen-Chung Shih.

Comic character animation using Bayesian estimation. Computer Animation and Virtual Worlds, 22(5):457– 470, September/October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xu:2011:PDA

[425] Jun Xu, Xiang Li, Yangchun Ren, and Weidong Geng. Performancedriven animation of hand-drawn cartoon faces. Computer Animation and Virtual Worlds, 22(5):471–483, September/ October 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2011:IIe

[426] Anonymous. Issue information. Computer Animation and Virtual Worlds, 22 (6):1–2, November/December 2011. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2011:Ec

[427] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 22(6):485, November/December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2011:ALB

[428] Changbo Wang, Chenhui Li, Jinqiu Dai, and Yang Li. Adaptive latticebased light rendering of participating media. Computer Animation and Virtual Worlds, 22(6):487–498, November/ December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2011:OIV

[429] Yijiang Zhang, Julien Pettré, Jan Ondřej, Xueying Qin, Qunsheng Peng, and Stéphane Donikian. Online inserting virtual characters into dynamic video scenes. Computer Animation and Virtual Worlds, 22(6):499–510, November/December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2011:PSP

[430] Jong-Hyuk Kim, Jung-Ju Choi, and Christoph M. Hoffmann. Pose space parameterization and style transfer of skin deformation. Computer Animation and Virtual Worlds, 22(6):511–518, November/December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2011:SCR

[431] Jianping Hu, Xiuping Liu, and Qi Xie. Subdivision connectivity remeshing and its applications. Computer Animation and Virtual Worlds, 22(6):519–528, November/December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cho:2011:RTR

[432] Kyusung Cho, Jinki Jung, Sang-Wook Lee, Sang Ok Lim, and Hyun Seung Yang. Real-time recognition and tracking for augmented reality books. Computer Animation and Virtual Worlds, 22 (6):529–541, November/December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ulusoy:2011:ADV

[433] Ilkay Ulusoy, Erdem Akagunduz, and Murat Yirci. Anatomical and dynamic volume spline model applied to facial soft tissue. Computer Animation and Virtual Worlds, 22(6):543–554, November/December 2011. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:IIa

[434] Anonymous. Issue information. Computer Animation and Virtual Worlds, 23 (1):i–ii, January/February 2012. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2012:ECS

[435] Daniel Thalmann, Saad Ali, and Petros Faloutsos. Editorial for the CAVW special issue on real-time crowd simulation. Computer Animation and Virtual Worlds, 23(1):1, January/February 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wagoum:2012:EVS

[436] Armel Ulrich Kemloh Wagoum, Mohcine Chraibi, Jonas Mehlich, Armin Seyfried, and Andreas Schadschneider. Efficient and validated simulation of crowds for an evacuation assistant. Computer Animation and Virtual Worlds, 23(1):3–15, January/February 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sun:2012:ASD

[437] Libo Sun, Alexander Shoulson, Pengfei Huang, Nicole Nelson, Wenhu Qin, Ani Nenkova, and Norman I. Badler.

Animating synthetic dyadic conversations with variations based on context and agent attributes. Computer Animation and Virtual Worlds, 23(1):17– 32, January/February 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Beacco:2012:ERA

[438] A. Beacco, C. Andujar, N. Pelechano, and B. Spanlang. Efficient rendering of animated characters through optimized per-joint impostors. Computer Animation and Virtual Worlds, 23(1): 33–47, January/February 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Musse:2012:TQA

[439] Soraia R. Musse, Vinicius J. Cassol, and Cláudio R. Jung. Towards a quantitative approach for comparing crowds. Computer Animation and Virtual Worlds, 23 (1):49–57, January/February 2012. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

vanToll:2012:RTD

[440] Wouter G. van Toll, Atlas F. Cook Iv, and Roland Geraerts. Real-time densitybased crowd simulation. Computer Animation and Virtual Worlds, 23(1):59– 69, January/February 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:EB

[441] Anonymous. Editorial board. Computer Animation and Virtual Worlds, 23(2):i–ii, March/April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:EI

[442] Anonymous. Editorial issue 23.2. Computer Animation and Virtual Worlds, 23(2):71, March/April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Brand:2012:MCS

[443] Sandy Brand and Rafael Bidarra. Multicore scalable and efficient pathfinding with Parallel Ripple Search. Computer Animation and Virtual Worlds, 23 (2):73–85, March/April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lopez:2012:STP

[444] Thomas Lopez, Fabrice Lamarche, and Tsai-Yen Li. Space-time planning in changing environments: using dynamic objects for accessibility. Computer Animation and Virtual Worlds, 23(2): 87–99, March/April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Casas:2012:PAP

[445] Dan Casas, Margara Tejera, Jean-Yves Guillemaut, and Adrian Hilton. Parametric animation of performancecaptured mesh sequences. Computer Animation and Virtual Worlds, 23(2): 101–111, March/April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Rungjiratananon:2012:AST

[446] Witawat Rungjiratananon, Yoshihiro Kanamori, Napaporn Metaaphanon, Yosuke Bando, Bing-Yu Chen, and Tomoyuki Nishita. Animating strings with twisting, tearing and flicking effects. Computer Animation and Virtual Worlds, 23(2):113–124, March/ April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pantuwong:2012:NTB

[447] Natapon Pantuwong and Masanori Sugimoto. A novel template-based automatic rigging algorithm for articulatedcharacter animation. Computer Animation and Virtual Worlds, 23(2): 125–141, March/April 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:IIb

[448] Anonymous. Issue information. Computer Animation and Virtual Worlds, 23 (3–4):i–ii, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:E

[449] Anonymous. Editorial. Computer Animation and Virtual Worlds, 23(3–4): 143–144, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jiang:2012:VER

[450] Kai Jiang, Xiaowu Chen, Yu Zhang, and Qinping Zhao. Video event representation and inference on And-Or graph. Computer Animation and Virtual Worlds, 23(3–4):145–154, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2012:EMN

[451] Jinmo Kim and Hyungje Cho. Efficient modeling of numerous trees by introducing growth volume for real-time virtual ecosystems. Computer Animation and Virtual Worlds, 23(3–4):155–165, May/ August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zeng:2012:VDS

[452] Ming Zeng, Lin Liang, Xinguo Liu, and Hujun Bao. Video-driven stateaware facial animation. Computer Animation and Virtual Worlds, 23(3–4): 167–178, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shen:2012:VCO

[453] Yang Shen, Xiao Lin, Yan Gao, Bin Sheng, and Qisong Liu. Video composition by optimized 3D mean-value coordinates. Computer Animation and Virtual Worlds, 23(3–4):179–190, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kang:2012:GCM

[454] Dongwann Kang, Yongjin Ohn, Myounghun Han, and Kyunghyun Yoon. Generation of coherent mosaic animations: enhancement and evaluation of temporal coherence. Computer Animation and Virtual Worlds, 23(3–4): 191–202, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yoshiyasu:2012:EBI

[455] Yusuke Yoshiyasu and Nobutoshi Yamazaki. Example-based inverse kinematics using cage. Computer Animation and Virtual Worlds, 23(3–4): 203–213, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tan:2012:CPR

[456] Charlie Irawan Tan and Wen Kai Tai. Characteristics preserving racer animation: a data-driven race path synthesis in formation space. Computer Animation and Virtual Worlds, 23(3–4): 215–223, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yoshiyasu:2012:DAS

[457] Yusuke Yoshiyasu and Nobutoshi Yamazaki. Detail-aware spatial deformation transfer. Computer Animation and Virtual Worlds, 23(3–4):225–233, May/ August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ma:2012:BMI

[458] Wan-Chun Ma, Yi-Hua Wang, Graham Fyffe, Bing-Yu Chen, and Paul Debevec. A blendshape model that incorporates physical interaction. Computer Animation and Virtual Worlds, 23(3–4): 235–243, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2012:IBA

[459] Sang Won Lee and Kang Hoon Lee. Interactive buildup of animation sequences with captured motion data. Computer Animation and Virtual Worlds, 23(3–4): 245–251, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kang:2012:PCR

[460] Young-Min Kang and Chang-Sik Cho. Photorealistic cloth in real-time applications. Computer Animation and Virtual Worlds, 23(3–4):253–265, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2012:ICB

[461] Meng Yang, Longsheng Jiang, Xiaosheng Li, Youquan Liu, Xuehui Liu, and Enhua Wu. Interactive coupling between a tree and raindrops. Computer Animation and Virtual Worlds, 23(3–4): 267–277, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2012:SMF

[462] Changbo Wang, Qiang Zhang, Huajun Xiao, and Qiuyan Shen. Simulation of multiple fluids with solid–liquid phase transition. Computer Animation and Virtual Worlds, 23(3–4):279–289, May/ August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oh:2012:PSH

[463] Seungtaik Oh, Seunghyup Shin, and Hyeryeong Jun. Practical simulation of hierarchical brittle fracture. Computer Animation and Virtual Worlds, 23(3–4): 291–300, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Djado:2012:PBD

[464] Khalid Djado, Richard Egli, and Fabrice Granger. Particle-based drop animation on meshes in real time. Computer Animation and Virtual Worlds, 23 (3–4):301–309, May/August 2012. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jund:2012:USC

[465] Thomas Jund, Pierre Kraemer, and David Cazier. A unified structure for crowd simulation. Computer Animation and Virtual Worlds, 23(3–4): 311–320, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ennis:2012:PPF

[466] Cathy Ennis and Carol O’Sullivan. Perceptually plausible formations for virtual conversers. Computer Animation and Virtual Worlds, 23(3–4):321– 329, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2012:AEE

[467] Ahyoung Choi, Celso De Melo, Woontack Woo, and Jonathan Gratch. Affective engagement to emotional facial expressions of embodied social agents in a decision-making game. Computer Animation and Virtual Worlds, 23(3–4): 331–342, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2012:CDL

[468] Weizi Li, Zichao Di, and Jan M. Allbeck. Crowd distribution and location preference. Computer Animation and Virtual Worlds, 23(3–4):343–351, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2012:VCC

[469] Xiaowu Chen, Hengyuan Zhang, and Qinping Zhao. Virtual calligraphic carving through smoothness scalar field and brush-pressure distribution. Computer Animation and Virtual Worlds, 23(3–4): 353–361, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2012:ASS

[470] Hyeon-Joong Kim, A. Cengiz Öztireli, Markus Gross, and Soo-Mi Choi. Adaptive surface splatting for facial rendering. Computer Animation and Virtual Worlds, 23(3–4):363–373, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Du:2012:FCC

[471] Peng Du, Min Tang, and Ruofeng Tong.

Fast continuous collision culling with deforming noncollinear filters. Computer Animation and Virtual Worlds, 23(3–4): 375–383, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bae:2012:UGV

[472] Myungsoo Bae, Jinwook Kim, and Young J. Kim. User-guided volumetric approximation using swept sphere volumes for physically based animation. Computer Animation and Virtual Worlds, 23(3–4):385–394, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2012:PBO

[473] Youquan Liu, Yanyun Chen, Wen Wu, Nelson Max, and Enhua Wu. Physically based object withering simulation. Computer Animation and Virtual Worlds, 23 (3–4):395–406, May/August 2012. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2012:DFP

[474] Xu Zhao, Zhong Zhou, Ye Duan, and Wei Wu. Detail-feature-preserving surface reconstruction. Computer Animation and Virtual Worlds, 23(3–4):

407–416, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2012:SIS

[475] Hwan-Jik Lee, Hyun Joon Shin, and Jung-Ju Choi. Single image summarization of 3D animation using depth images. Computer Animation and Virtual Worlds, 23(3–4):417–424, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gutierrez:2012:SSD

[476] Luis F. Gutiérrez, Iker Aguinaga, Matthias Harders, and Félix Ramos. Speeding up the simulation of deformable objects through mesh improvement. Computer Animation and Virtual Worlds, 23(3–4):425–433, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ferraris:2012:FBP

[477] John Ferraris, Feng Tian, and Christos Gatzidis. Feature-based probabilistic texture blending with feature variations for terrains. Computer Animation and Virtual Worlds, 23(3–4):435–445, May/ August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2012:DDD

[478] Yong Zhao, Bin Pan, Chunxia Xiao, and Qunsheng Peng. Dual-domain deformation transfer for triangular meshes. Computer Animation and Virtual Worlds, 23(3–4):447–456, May/ August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Fratarcangeli:2012:PBF

[479] Marco Fratarcangeli. Position-based facial animation synthesis. Computer Animation and Virtual Worlds, 23(3–4): 457–466, May/August 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:IIc

[480] Anonymous. Issue information. Computer Animation and Virtual Worlds, 23 (5):i–ii, September/October 2012. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2012:EIa

[481] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 23.5. Computer Animation and Virtual Worlds, 23 (5):467, September/October 2012. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2012:HMR

[482] Mingyang Zhu, Huaijiang Sun, Rongyi Lan, and Bin Li. Human motion retrieval using topic model. Computer Animation and Virtual Worlds, 23(5):469– 476, September/October 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shao:2012:PBS

[483] Xuqiang Shao, Zhong Zhou, and Wei Wu. Particle-based simulation of bubbles in water–solid interaction. Computer Animation and Virtual Worlds, 23 (5):477–487, September/October 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jung:2012:RTC

[484] Hoeryong Jung and Doo Yong Lee. Realtime cutting simulation of meshless deformable object using dynamic bounding volume hierarchy. Computer Animation and Virtual Worlds, 23(5):489– 501, September/October 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Alvarez:2012:NMD

[485] Hugo Álvarez, Ibai Leizea, and Diego Borro. A new marker design for a robust marker tracking system against occlusions. Computer Animation and Virtual Worlds, 23(5):503–518, September/ October 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2012:VSM

[486] Yong Zhao and Jianhui Liu. Volumetric subspace mesh deformation with structure preservation. Computer Animation and Virtual Worlds, 23(5):519– 532, September/October 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2012:III

[487] Anonymous. Issue information: Issue information. Computer Animation and Virtual Worlds, 23(6):i–ii, November 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2012:EIb

[488] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 23.6. Computer Animation and Virtual Worlds, 23(6):533, November 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

vanToll:2012:NMD

[489] Wouter G. van Toll, Atlas F. Cook IV, and Roland Geraerts. A navigation mesh for dynamic environments. Computer Animation and Virtual Worlds, 23 (6):535–546, November 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zong:2012:PIT

[490] Dan Zong, Chunpeng Li, Shihong Xia, and Zhaoqi Wang. Planning interactive task for intelligent characters. Computer Animation and Virtual Worlds, 23 (6):547–557, November 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jin:2012:OKE

[491] Chao Jin, Thomas Fevens, and Sudhir Mudur. Optimized keyframe extraction for 3D character animations. Computer Animation and Virtual Worlds, 23 (6):559–568, November 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Allen:2012:PII

[492] Brian F. Allen, Nadia Magnenat-Thalmann, and Daniel Thalmann. Politeness improves interactivity in dense crowds. Computer Animation and Virtual Worlds, 23(6):569–578, November 2012. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2013:IIIa

[493] Anonymous. Issue information: Issue information. Computer Animation and Virtual Worlds, 24(1):i–ii, January 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2013:EIa

[494] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 24.1. Computer Animation and Virtual Worlds, 24 (1):1, January 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Karim:2013:PLM

[495] Ahmad Abdul Karim, Thibaut Gaudin, Alexandre Meyer, Axel Buendia, and Saida Bouakaz. Procedural locomotion of multilegged characters in dynamic environments. Computer Animation and Virtual Worlds, 24(1):3–15, January 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kumar:2013:NPP

[496] Amit Kumar and Aparajita Ojha. Natural path planning using wavelet noise in static environment. Computer Animation and Virtual Worlds, 24(1):17–24, January 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2013:VSP

[497] Yanzhen Wang, Yueshan Xiong, Kai Xu, and Dong Liu. vKASS: a surgical procedure simulation system for arthroscopic anterior cruciate ligament reconstruction. Computer Animation and Virtual Worlds, 24(1):25–41, January 2013. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hegde:2013:PRT

[498] Siddharth Hegde, Christos Gatzidis, and Feng Tian. Painterly rendering techniques: a state-of-the-art review of current approaches. Computer Animation and Virtual Worlds, 24(1):43–64, January 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

apCenydd:2013:EAA

[499] Llyr ap Cenydd and Bill Teahan. An embodied approach to arthropod animation. Computer Animation and Virtual Worlds, 24(1):65–83, January 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2013:IIIb

[500] Anonymous. Issue information: Issue information. Computer Animation and Virtual Worlds, 24(2):i–ii, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2013:EIb

[501] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 24.2. Computer Animation and Virtual Worlds, 24 (2):85, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2013:SIP

[502] Ting-Chieh Huang, Yi-Jheng Huang, and Wen-Chieh Lin. Special issue paper: Real-time horse gait synthesis. Computer Animation and Virtual Worlds, 24(2):87–95, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2013:NMC

[503] Guanyang Liu and Keke Lu. Networked multiplayer cooperative interaction using decoupled motion control method in a shared virtual environment with haptic, visual and movement feedback. Computer Animation and Virtual Worlds, 24(2):97–109, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Buche:2013:ABV

[504] C. Buche and P. De Loor. Anticipatory behavior in virtual universe, application to a virtual juggler. Computer Animation and Virtual Worlds, 24(2):111–125, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Neubauer:2013:HCH

[505] André Neubauer, Rupert Brooks, Iman Brouwer, Patricia Debergue, and Denis Laroche. Haptic collision handling for simulation of transnasal surgery. Computer Animation and Virtual Worlds, 24(2):127–141, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Haciomeroglu:2013:HAD

[506] Murat Haciomeroglu, Cumhur Yigit Ozcan, Oner Barut, Levent Seckin, and Hayri Sever. Hardware-accelerated dynamic clustering of virtual crowd members. Computer Animation and Virtual Worlds, 24(2):143–153, March 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2013:IIIc

[507] Anonymous. Issue information: Issue Information. Computer Animation and Virtual Worlds, 24(3–4):i–ii, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2013:E

[508] Daniel Thalmann, Tolga Capin, and Selim Balcisoy. Editorial. Computer Animation and Virtual Worlds, 24(3–4):153–154, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2013:SAS

[509] Seung In Park, Francis Quek, and Yong Cao. Simulating and animating social dynamics: embedding small pedestrian groups in crowds. Computer Animation and Virtual Worlds, 24(3–4):155–164, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sun:2013:SRC

[510] Libo Sun, Xiaona Li, and Wenhu Qin.

Simulating realistic crowd based on agent trajectories. Computer Animation and Virtual Worlds, 24(3–4):165–172, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2013:CAB

[511] Jin Hyoung Park, Francisco Arturo Rojas, and Hyun Seung Yang. A collision avoidance behavior model for crowd simulation based on psychological findings. Computer Animation and Virtual Worlds, 24(3–4):173–183, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jung:2013:HCG

[512] Hwi-Ryong Jung, Sun-Tae Kim, Junyong Noh, and Jeong-Mo Hong. A heterogeneous CPU–GPU parallel approach to a multigrid Poisson solver for incompressible fluid simulation. Computer Animation and Virtual Worlds, 24(3–4):185–193, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Akinci:2013:CES

[513] Nadir Akinci, Jens Cornelis, Gizem Akinci, and Matthias Teschner. Coupling elastic solids with smoothed particle hydrodynamics fluids. Computer Animation and Virtual Worlds, 24(3–4):195–203, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2013:RMI

[514] Guijuan Zhang, Dianjie Lu, Dengming Zhu, Lei Lv, Hong Liu, and Xiangxu Meng. Rigid-motion-inspired liquid character animation. Computer Animation and Virtual Worlds, 24(3–4):205–213, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2013:FRA

[515] Feibin Chen, Changbo Wang, Buying Xie, and Hong Qin. Flexible and rapid animation of brittle fracture using the smoothed particle hydrodynamics formulation. Computer Animation and Virtual Worlds, 24(3–4):215–224, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nikfetrat:2013:FPA

[516] Nima Nikfetrat and Won-Sook Lee. Fire pattern analysis and synthesis using EigenFires and motion transitions. Computer Animation and Virtual Worlds, 24(3–4):225–235, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2013:PMT

[517] Jinmo Kim, Daeyeoul Kim, and Hyungje Cho. Procedural modeling of trees based on convolution sums of divisor functions for real-time virtual ecosystems. Computer Animation and Virtual Worlds, 24(3–4):237–246, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Um:2013:PDS

[518] Kiwon Um, Tae-Yong Kim, Youngdon Kwon, and JungHyun Han. Porous deformable shell simulation with surface water flow and saturation. Computer Animation and Virtual Worlds, 24(3–4):247–254, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2013:FPI

[519] Hongyu Wu, Xiaowu Chen, Mengxia Yang, and Zhihong Fang. Facial performance illumination transfer from a single video using interpolation in nonskin region. Computer Animation and Virtual Worlds, 24(3–4):255–263, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2013:RAI

[520] Chunxiao Liu, Hong Li, Qunsheng Peng, Xun Wang, and Enhua Wu. Relighting abstracted image via salient edge-guided luminance field optimization. Computer Animation and Virtual Worlds, 24(3–4):265–274, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2013:PBC

[521] Cheng-Guo Huang, Tsung-Shian Huang, Wen-Chieh Lin, and Jung-Hong Chuang. Physically based cosmetic rendering. Computer Animation and Virtual Worlds, 24(3–4):275–283, May 2013.

CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jaklin:2013:RTP

[522] Norman Jaklin, Atlas Cook IV, and Roland Geraerts. Real-time path planning in heterogeneous environments. Computer Animation and Virtual Worlds, 24(3–4):285–295, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

You:2013:RPS

[523] Mi You, Taekwon Jang, Seunghoon Cha, Jihwan Kim, and Junyong Noh. Realistic paint simulation based on fluidity, diffusion, and absorption. Computer Animation and Virtual Worlds, 24(3–4):297–306, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liang:2013:PDA

[524] Dongxue Liang and Kyoungju Park. Pencil drawing animation from a video. Computer Animation and Virtual Worlds, 24(3–4):307–316, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Park:2013:RDH

[525] Jaesung Park, Minsub Shim, Seon-Young Park, Yunku Kang, and Myung-Soo Kim. Realistic deformation of 3D human blood vessels. Computer Animation and Virtual Worlds, 24 (3–4):317–325, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Moussa:2013:TSR

[526] Maher Ben Moussa and Nadia Magnenat-Thalmann. Toward socially responsible agents: integrating attachment and learning in emotional decisionmaking. Computer Animation and Virtual Worlds, 24(3–4):327–334, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2013:TPV

[527] JeeHang Lee, Tingting Li, and Julian Padget. Towards polite virtual agents using social reasoning techniques. Computer Animation and Virtual Worlds, 24(3–4):335–343, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Luo:2013:ISG

[528] Linbo Luo, Haiyan Yin, Wentong Cai, Michael Lees, and Suiping Zhou. Interactive scenario generation for missionbased virtual training. Computer Animation and Virtual Worlds, 24(3–4): 345–354, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cimen:2013:CHM

[529] Gokcen Cimen, Hacer Ilhan, Tolga Capin, and Hasmet Gurcay. Classification of human motion based on affective state descriptors. Computer Animation and Virtual Worlds, 24(3–4):355–363, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Luo:2013:CMS

[530] Guoliang Luo, Frederic Cordier, and Hyewon Seo. Compression of 3D mesh sequences by temporal segmentation. Computer Animation and Virtual Worlds, 24(3–4):365–375, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jeong:2013:DSW

[531] Moon-Hwan Jeong and Hyeong-Seok Ko. Draft-space warping: grading of clothes based on parametrized draft. Computer Animation and Virtual Worlds, 24(3–4):377–386, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2013:LAD

[532] Yadang Chen, Chuanyan Hao, Zhongmou Cai, Wen Wu, and Enhua Wu. Live accurate and dense reconstruction from a handheld camera. Computer Animation and Virtual Worlds, 24(3–4):387–397, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qi:2013:SFH

[533] Tian Qi, Yinfu Feng, Jun Xiao, Yueting Zhuang, Xiaosong Yang, and Jianjun Zhang. A semantic feature for human motion retrieval. Computer Animation and Virtual Worlds, 24(3–4):399–407, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2013:IEM

[534] Siwang Li, Jin Huang, Mathieu Desbrun, and Xiaogang Jin. Interactive elastic motion editing through space– time position constraints. Computer Animation and Virtual Worlds, 24 (3–4):409–417, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

LeNaour:2013:SCM

[535] T. Le Naour, N. Courty, and S. Gibet. Spatiotemporal coupling with the 3D+t motion Laplacian. Computer Animation and Virtual Worlds, 24(3–4):419–428, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bozgeyikli:2013:ITO

[536] Gamze Bozgeyikli, Evren Bozgeyikli, and Veysi İşler. Introducing tangible objects into motion controlled gameplay using Microsoft Kinect^{T M}. Computer Animation and Virtual Worlds, 24(3–4):429–441, May 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2013:IIId

[537] Anonymous. Issue information: Issue information. Computer Animation and Virtual Worlds, 24(5):i–ii, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2013:EIc

[538] Nadia Magnenat-Thalmann. Editorial issue 24.5. Computer Animation and Virtual Worlds, 24(5):443, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2013:SAA

[539] Yuanfeng Zhu, Ajay Sundar Ramakrishnan, Bernd Hamann, and Michael Neff. A system for automatic animation of piano performances. Computer Animation and Virtual Worlds, 24 (5):445–457, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Perriollat:2013:CMB

[540] Mathieu Perriollat and Adrien Bartoli. A computational model of bounded developable surfaces with application to image-based three-dimensional reconstruction. Computer Animation and Virtual Worlds, 24(5):459–476, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tence:2013:COL

[541] F. Tence, L. Gaubert, J. Soler, P. De Loor, and C. Buche. Chameleon: online learning for believable behaviors based on humans imitation in computer games. Computer Animation and Virtual Worlds, 24(5):477–495, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2013:FSR

[542] Mingqi Yu and Hongyan Quan. Fluid surface reconstruction based on specular reflection model. Computer Animation and Virtual Worlds, 24(5): 497–510, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ozgen:2013:SCF

[543] Oktar Ozgen, Selcuk Sumengen, Marcelo Kallmann, Carlos Fm Coimbra, and Selim Balcisoy. Simulating colliding flows in smoothed particle hydrodynamics with fractional derivatives. Computer Animation and Virtual Worlds, 24 (5):511–523, September 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2013:IIIe

[544] Anonymous. Issue information: Issue information. Computer Animation and Virtual Worlds, 24(6):i–ii, November 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kallmann:2013:EI

[545] Marcelo Kallmann, Kostas Bekris, Nadia Magnenat-Thalmann, and Daniel Thalmann. Editorial issue 24.6. Computer Animation and Virtual Worlds, 24(6):525, November 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sandilands:2013:ICU

[546] Peter Sandilands, Myung Geol Choi, and Taku Komura. Interaction capture using magnetic sensors. Computer Animation and Virtual Worlds, 24 (6):527–538, November 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ahn:2013:AFE

[547] Junghyun Ahn, Stephane Gobron, Daniel Thalmann, and Ronan Boulic. Asymmetric facial expressions: revealing richer emotions for embodied conversational agents. Computer Animation and Virtual Worlds, 24(6):539–551, November 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Backman:2013:DCP

[548] Robert Backman and Marcelo Kallmann. Designing controllers for physicsbased characters with motion networks. Computer Animation and Virtual Worlds, 24(6):553–563, November 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2013:HMR

[549] Jongmin Kim, Yeongho Seol, and Jehee Lee. Human motion reconstruction from sparse 3D motion sensors using kernel CCA-based regression. Computer Animation and Virtual Worlds, 24 (6):565–576, November 2013. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2014:IIa

[550] Anonymous. Issue information. Computer Animation and Virtual Worlds, 25(1):i–ii, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2014:EIa

[551] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 25.1. Computer Animation and Virtual Worlds, 25 (1):1, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Feng:2014:FAC

[552] Andrew Feng, Yazhou Huang, Yuyu Xu, and Ari Shapiro. Fast, automatic character animation pipelines. Computer Animation and Virtual Worlds, 25(1):3–16, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Rantanen:2014:UPR

[553] Mika T. Rantanen and Martti Juhola. Using probabilistic roadmaps in changing environments. Computer Animation and Virtual Worlds, 25(1):17–31, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shi:2014:RTC

[554] Yan Shi, Fangtian Ying, Xuan Chen, Zhigeng Pan, and Jinhui Yu. Restoration of traditional Chinese shadow play — Piying art from tangible interaction. Computer Animation and Virtual Worlds, 25(1):33–43, January 2014. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Luo:2014:DSR

[555] Jianxin Luo, Guyu Hu, and Guiqiang Ni. Dual-space ray casting for height field rendering. Computer Animation and Virtual Worlds, 25(1):45–56, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

He:2014:FEH

[556] Zhiying He, Xiaohui Liang, Jian Wang, Qinping Zhao, and Chengyu Guo. Flexible editing of human motion by threeway decomposition. Computer Animation and Virtual Worlds, 25(1):57–68, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tripicchio:2014:MPP

[557] Paolo Tripicchio, Claudio Loconsole, Andrea Piarulli, Emanuele Ruffaldi, Franco Tecchia, and Massimo Bergamasco. On multiuser perspectives in passive stereographic virtual environments. Computer Animation and Virtual Worlds, 25(1):69–81, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lu:2014:AFA

[558] Xuequan Lu, Wenzhi Chen, Mingliang Xu, Zonghui Wang, Zhigang Deng, and Yangdong Ye. AA-FVDM: an accidentavoidance full velocity difference model for animating realistic street-level traffic in rural scenes. Computer Animation and Virtual Worlds, 25(1):83–97, January 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2014:IIb

[559] Anonymous. Issue information. Computer Animation and Virtual Worlds, 25 (2):i–ii, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2014:EIb

[560] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 25.2. Computer Animation and Virtual Worlds, 25(2):99–100, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2014:CPE

[561] Mingmin Zhang, Zhigeng Pan, Xixi Huang, Nan Xiang, Shuwen Wang, and Pengyu Zhu. Call for papers: Easy-Home: an online virtual home decoration system. Computer Animation and Virtual Worlds, 25(2):101–113, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2014:RML

[562] Weitao Li, Yuanfeng Zhou, Caiming Zhang, and Xuemei Li. Robust multilevel partition of unity implicits from triangular meshes. Computer Animation and Virtual Worlds, 25(2):115–127, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kirmizibayrak:2014:IFC

[563] Can Kirmizibayrak, Mike Wakid, Yeny Yim, Dimitre Hristov, and James K. Hahn. Interactive focus + context medical data exploration and editing. Computer Animation and Virtual Worlds, 25(2):129–141, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2014:VCP

[564] Chunxiao Zhang, Eric Dubois, and Yan Zhao. Virtual cubic panorama synthesis based on triangular reprojection. Computer Animation and Virtual Worlds, 25(2):143–154, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2014:IDC

[565] Shuai Li, Qinping Zhao, Shengfa Wang, Aimin Hao, and Hong Qin. Interactive deformation and cutting simulation directly using patient-specific volumetric images. Computer Animation and Virtual Worlds, 25(2):155–169, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lv:2014:OBG

[566] Lei Lv, Tianlu Mao, Xuecheng Liu, and Zhaoqi Wang. Optimization-based group performance deducing. Computer Animation and Virtual Worlds, 25(2):171–184, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xu:2014:RTG

[567] Tianchen Xu, Wen Wu, and Enhua Wu. Real-time generation of smoothedparticle hydrodynamics-based special effects in character animation. Computer Animation and Virtual Worlds, 25(2):185–198, March 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2014:IIc

[568] Anonymous. Issue information. Computer Animation and Virtual Worlds, 25(3–4):i–ii, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2014:EE

[569] Anonymous. Editorial: Editorial. Computer Animation and Virtual Worlds, 25(3–4):199–200, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shapiro:2014:RAC

[570] Ari Shapiro, Andrew Feng, Ruizhe Wang, Hao Li, Mark Bolas, Gerard Medioni, and Evan Suma. Rapid avatar capture and simulation using commodity depth sensors. Computer Animation and Virtual Worlds, 25(3–4):201–211, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zollhofer:2014:IMB

[571] Michael Zollhöfer, Justus Thies, Matteo Colaianni, Marc Stamminger, and Günther Greiner. Interactive modelbased reconstruction of the human head using an RGB-D sensor. Computer Animation and Virtual Worlds, 25(3–4):213–222, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Castillo:2014:SSF

[572] Susana Castillo, Christian Wallraven, and Douglas W. Cunningham. The semantic space for facial communication. Computer Animation and Virtual Worlds, 25(3–4):225–233, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Fang:2014:RTD

[573] Xiaoxin Fang, Bin Sheng, Wen Wu, Zengzhi Fan, and Lizhuang Ma. Realtime depth-of-field rendering using single-layer composition. Computer Animation and Virtual Worlds, 25(3–4): 235–243, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2014:HLD

[574] Shengchuan Zhou, Innfarn Yoo, Bedrich Benes, and Ge Chen. A hybrid level-of-detail representation for largescale urban scenes rendering. Computer Animation and Virtual Worlds, 25(3–4):245–255, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yasmin:2014:HEN

[575] Shamima Yasmin, Nan Du, James Chen, and Yusheng Feng. A haptic-enabled novel approach to cardiovascular visualization. Computer Animation and Virtual Worlds, 25(3–4):257–271, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tang:2014:HMR

[576] Zhangpeng Tang, Jun Xiao, Yinfu Feng, Xiaosong Yang, and Jian Zhang. Human motion retrieval based on freehand sketch. Computer Animation and Virtual Worlds, 25(3–4):273–281, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lv:2014:GAA

[577] Na Lv, Yan Huang, Zhiquan Feng, and Jingliang Peng. A genetic algorithm approach to human motion capture data segmentation. Computer Animation and Virtual Worlds, 25(3–4):283–292, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qi:2014:RTM

[578] Tian Qi, Jun Xiao, Yueting Zhuang, Hanzhi Zhang, Xiaosong Yang, Jianjun Zhang, and Yinfu Feng. Realtime motion data annotation via action string. Computer Animation and Virtual Worlds, 25(3–4):293–302, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2014:HMV

[579] Liuyang Zhou, Lifeng Shang, Hubert P. H. Shum, and Howard Leung. Human motion variation synthesis with multivariate Gaussian processes. Computer Animation and Virtual Worlds, 25(3–4):303–311, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oshita:2014:IMS

[580] Masaki Oshita. Interactive motion synthesis with optimal blending. Computer Animation and Virtual Worlds, 25(3–4):313–321, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2014:BFD

[581] YoungBeom Kim and JungHyun Han. Bulging-free dual quaternion skinning. Computer Animation and Virtual Worlds, 25(3–4):323–331, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Stuvel:2014:HSC

[582] Sybren A. Stüvel, Nadia Magnenat-Thalmann, Daniel Thalmann, Arjan Egges, and A. Frank van der Stappen. Hierarchical structures for collision checking between virtual characters. Computer Animation and Virtual Worlds, 25(3–4):333–342, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pitiot:2014:DPA

[583] Thomas Pitiot, David Cazier, Thomas Jund, Arash Habibi, and Pierre Kraemer. Deformable polygonal agents in crowd simulation. Computer Animation and Virtual Worlds, 25(3–4):343–352, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2014:FMA

[584] Xinjie Wang, Linling Zhou, Zhigang Deng, and Xiaogang Jin. Flock morphing animation. Computer Animation and Virtual Worlds, 25(3–4):353–362, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lu:2014:PMA

[585] Xuequan Lu, Zonghui Wang, Mingliang Xu, Wenzhi Chen, and Zhigang Deng. A personality model for animating heterogeneous traffic behaviors. Computer Animation and Virtual Worlds, 25(3–4):363–373, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chu:2014:MSB

[586] Mei Ling Chu, Paolo Parigi, Kincho Law, and Jean-Claude Latombe. Modeling social behaviors in an evacuation simulator. Computer Animation and Virtual Worlds, 25(3–4):375–384, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2014:AOE

[587] Hua Wang, Tianlu Mao, Xingchen Kang, and Zhaoqi Wang. An all-in-one efficient lane-changing model for virtual traffic. Computer Animation and Virtual Worlds, 25(3–4):385–393, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Luo:2014:TDD

[588] Linbo Luo, Haiyan Yin, Wentong Cai, Michael Lees, Nasri Bin Othman, and Suiping Zhou. Towards a data-driven approach to scenario generation for serious games. Computer Animation and Virtual Worlds, 25(3–4):395–404, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pelkey:2014:PSV

[589] Cameron D. Pelkey and Jan M. Allbeck. Populating semantic virtual environments. Computer Animation and Virtual Worlds, 25(3–4):405–412, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2014:HMD

[590] Yong Zhao, Junyu Dong, Bin Pan, and Chunxia Xiao. Hierarchical mesh deformation with shape preservation. Computer Animation and Virtual Worlds, 25(3–4):413–422, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2014:RTP

[591] Chen Yang, Shuai Li, Lili Wang, Aimin Hao, and Hong Qin. Real-time physical deformation and cutting of heterogeneous objects via hybrid coupling of meshless approach and finite element method. Computer Animation and Virtual Worlds, 25(3–4):423–435, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Saito:2014:MMD

[592] Shunsuke Saito, Nobuyuki Umetani, and Shigeo Morishima. Macroscopic and microscopic deformation coupling in up-sampled cloth simulation. Computer Animation and Virtual Worlds, 25(3–4):437–446, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2014:ASD

[593] Xue Chen and Jieqing Feng. Adaptive skeleton-driven cages for mesh sequences. Computer Animation and Virtual Worlds, 25(3–4):447–455, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2014:RTS

[594] Min Gyu Choi. Real-time simulation of ductile fracture with oriented particles. Computer Animation and Virtual Worlds, 25(3–4):457–465, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2014:TSS

[595] Ben Yang and Xiaogang Jin. Turbulence synthesis for shape-controllable smoke animation. Computer Animation and Virtual Worlds, 25(3–4):467–474, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2014:VFA

[596] Shiguang Liu, Yixin Xu, Junyong Noh, and Yiying Tong. Visual fluid animation via lifting wavelet transform. Computer Animation and Virtual Worlds, 25(3–4):475–485, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lopez:2014:CVT

[597] Thomas Lopez, Pierre Chevaillier, Valérie Gouranton, Paul Evrard, Florian Nouviale, Mukesh Barange, Rozenn Bouville, and Bruno Arnaldi. Collaborative virtual training with physical and communicative autonomous agents. Computer Animation and Virtual Worlds, 25(3–4):487–495, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Almajano:2014:AAA

[598] Pablo Almajano, Maite Lopez-Sanchez, Inmaculada Rodriguez, and Tomas Trescak. Assistant agents to advice users in hybrid structured 3D virtual environments. Computer Animation and Virtual Worlds, 25(3–4):497–506, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Karimaghalou:2014:MSS

[599] Nahid Karimaghalou, Ulysses Bernardet, and Steve DiPaola. A model for social spatial behavior in virtual characters. Computer Animation and Virtual Worlds, 25(3–4):507–519, May 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2014:IId

[600] Anonymous. Issue information. Computer Animation and Virtual Worlds, 25(5–6):i–ii, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2014:EI

[601] Anonymous. Editorial issue 25.5-6. Computer Animation and Virtual Worlds, 25(5–6):519, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Perumal:2014:EAC

[602] Logah Perumal. Euler angles: conversion of arbitrary rotation sequences to specific rotation sequence. Computer Animation and Virtual Worlds, 25(5– 6):521–529, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shum:2014:NPB

[603] Hubert P. H. Shum, Ludovic Hoyet, Edmond S. L. Ho, Taku Komura, and Franck Multon. Natural preparation behavior synthesis. Computer Animation and Virtual Worlds, 25(5–6): 531–542, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2014:PRI

[604] Qiao Huang and Jie Yang. Person re-identification with local descriptors across multicameras. Computer Animation and Virtual Worlds, 25(5–6): 543–551, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ling-yu:2014:FTT

[605] Wei Ling-yu. A faster triangle-to-triangle intersection test algorithm. Computer Animation and Virtual Worlds, 25(5–6):553–559, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jang:2014:EGI

[606] Bong gyu Jang and Gerard Jounghyun Kim. Evaluation of grounded isometric interface for whole-body navigation in virtual environments. Computer Animation and Virtual Worlds, 25(5–6): 561–575, September 2014. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:IIa

[607] Anonymous. Issue information. Computer Animation and Virtual Worlds, 26(1):i–ii, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2015:EI

[608] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 26.1. Computer Animation and Virtual Worlds, 26 (1):1, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2015:FAD

[609] Changbo Wang, Qiang Zhang, Fanlong Kong, and Yusheng Gao. Fast animation of debris flow with mixed adaptive grid refinement. Computer Animation and Virtual Worlds, 26(1):3–14, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2015:SDS

[610] Lin Zhang, Fei Dou, Zhong Zhou, and Wei Wu. Streaming 3D deforming surfaces with dynamic resolution control. Computer Animation and Virtual Worlds, 26(1):15–28, January 2015. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2015:AST

[611] Xiaoqiang Zhu, Xiaogang Jin, and Lihua You. Analytical solutions for treelike structure modelling using subdivision surfaces. Computer Animation and Virtual Worlds, 26(1):29–42, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2015:POS

[612] Chen Huang, Jian Zhu, Hanqiu Sun, and Enhua Wu. Parallel-optimizing SPH fluid simulation for realistic VR environments. Computer Animation and Virtual Worlds, 26(1):43–54, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Warburton:2015:PBF

[613] Mark Warburton and Steve Maddock.

Physically-based forehead animation including wrinkles. Computer Animation and Virtual Worlds, 26(1):55–68, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2015:CBP

[614] Sang-Gil Lee and Byeong-Seok Shin.

Contour-based polygonal ambient occlusion using a single-depth texture. Computer Animation and Virtual Worlds, 26(1):69–77, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shao:2015:RSS

[615] Xuqiang Shao, Zhong Zhou, Jinsong Zhang, and Wei Wu. Realistic and stable simulation of turbulent details behind objects in smoothed-particle hydrodynamics fluids. Computer Animation and Virtual Worlds, 26(1):79–94, January 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:IIb

[616] Anonymous. Issue information. Computer Animation and Virtual Worlds, 26 (2):i–ii, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2015:EI

[617] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 26.2. Computer Animation and Virtual Worlds, 26(2):95–96, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jones:2015:DSA

[618] Ben Jones, Jovan Popovic, James Mc-Cann, Wilmot Li, and Adam Bargteil. Dynamic sprites: artistic authoring of interactive animations. Computer Animation and Virtual Worlds, 26(2):97– 108, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gerszewski:2015:BES

[619] Dan Gerszewski, Ladislav Kavan, Peter-Pike Sloan, and Adam W. Bargteil. Basis enrichment and solid–fluid coupling for model-reduced fluid simulation. Computer Animation and Virtual Worlds, 26(2):109–117, March 2015.

CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ninomiya:2015:PAC

[620] Kai Ninomiya, Mubbasir Kapadia, Alexander Shoulson, Francisco Garcia, and Norman Badler. Planning approaches to constraint-aware navigation in dynamic environments. Computer Animation and Virtual Worlds, 26(2):119–139, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2015:LAC

[621] Zhanpeng Huang, Liang Han, and Guanghong Gong. A local adaptive Catmull–Rom to reduce numerical dissipation of semi-Lagrangian advection. Computer Animation and Virtual Worlds, 26(2):141–146, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bonaventura:2015:SRU

[622] Xavier Bonaventura, Jianwei Guo, Weiliang Meng, Miquel Feixas, Xiaopeng Zhang, and Mateu Sbert. 3D shape retrieval using viewpoint informationtheoretic measures. Computer Animation and Virtual Worlds, 26(2):147–156, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jang:2015:GAA

[623] Taekwon Jang and Junyong Noh. A geometric approach to animating thin surface features in smoothed particle hydrodynamics water. Computer Animation and Virtual Worlds, 26(2):161–172, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xian:2015:EEC

[624] Chuhua Xian, Guiqing Li, and Yunhui Xiong. Efficient and effective cage generation by region decomposition. Computer Animation and Virtual Worlds, 26(2):173–184, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sasaki:2015:FFP

[625] Naoki Sasaki, Hsiang-Ting Chen, Daisuke Sakamoto, and Takeo Igarashi. Facetons: face primitives for building 3D architectural models in virtual environments. Computer Animation and Virtual Worlds, 26(2):185–194, March 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:IIc

[626] Anonymous. Issue information. Computer Animation and Virtual Worlds, 26(3–4):i–ii, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:E

[627] Anonymous. Editorial. Computer Animation and Virtual Worlds, 26 (3–4):195–196, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2015:SEE

[628] Shan Yang and Ming Lin. Simultaneous estimation of elasticity for multiple deformable bodies. Computer Animation and Virtual Worlds, 26(3–4):197–206, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liang:2015:AOD

[629] Hui Liang, Jian Chang, Xiaosong Yang, Lihua You, Shaojun Bian, and Jian Jun Zhang. Advanced ordinary differential equation based head modelling for Chinese marionette art preservation. Computer Animation and Virtual Worlds, 26(3–4):207–216, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

enAimen:2015:SLB

[630] Gökçen Çimen, Zümra Kavafoğlu, Ersan Kavafoğlu, Tolga Çapın, and Hašmet Gürcay. Skill learning based catching motion control. Computer Animation and Virtual Worlds, 26(3–4):217–225, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vogt:2015:BGI

[631] David Vogt, Ben Lorenz, Steve Grehl, and Bernhard Jung. Behavior generation for interactive virtual humans using context-dependent interaction meshes and automated constraint extraction. Computer Animation and Virtual Worlds, 26(3–4):227–235, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2015:PPC

[632] Juzheng Zhang, Jianmin Zheng, and Nadia Magnenat-Thalmann. PCMD: personality-characterized mood dynamics model toward personalized virtual characters. Computer Animation and Virtual Worlds, 26(3–4):237–245, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kochanowicz:2015:DDT

[633] Jaroslaw Kochanowicz, Ah-Hwee Tan, and Daniel Thalmann. Dramaturgical and dissonance theories in explicit social context modeling for complex agents. Computer Animation and Virtual Worlds, 26(3–4):247–257, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jo:2015:SAC

[634] Dongsik Jo, Ki-Hong Kim, and Gerard Jounghyun Kim. SpaceTime: adaptive control of the teleported avatar for improved AR tele-conference experience. Computer Animation and Virtual Worlds, 26(3–4):259–269, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nagendran:2015:STU

[635] Arjun Nagendran, Anthony Steed, Brian Kelly, and Ye Pan. Symmetric telepresence using robotic humanoid surrogates. Computer Animation and Virtual Worlds, 26(3–4):271–280, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yao:2015:RSA

[636] Junfeng Yao, Lei Lan, Wenlin Lin, Yingying She, and Binbin Zhang. Realistic and stable animation of cloth. Computer Animation and Virtual Worlds, 26(3–4):281–289, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jeong:2015:GCP

[637] Moon-Hwan Jeong, Dong-Hoon Han, and Hyeong-Seok Ko. Garment capture from a photograph. Computer Animation and Virtual Worlds, 26 (3–4):291–300, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pang:2015:ESE

[638] Zhiqiang Pang, Yong Zhao, and Chunxia Xiao. Effective skeletons extraction for animated surfaces based on geometry propagation. Computer Animation and Virtual Worlds, 26(3–4):301–309, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2015:IIH

[639] Xingzi Zhang and Alexei Sourin. Imageinspired haptic interaction. Computer Animation and Virtual Worlds, 26(3–4):311–319, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pan:2015:RTH

[640] Junjun Pan, Junxuan Bai, Xin Zhao, Aimin Hao, and Hong Qin. Realtime haptic manipulation and cutting of hybrid soft tissue models by extended position-based dynamics. Computer Animation and Virtual Worlds, 26(3–4):321–335, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2015:GAR

[641] Jiaxiang Guo, Zhiyong Yuan, Xiangyun Liao, Yaoyi Bai, and Qianfeng Lai. GPU-assisted real-time coupling of blood flow and vessel wall. Computer Animation and Virtual Worlds, 26(3–4):337–345, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baek:2015:MWA

[642] Seungho Baek, Kiwon Um, and JungHyun Han. Muddy water animation with different details. Computer Animation and Virtual Worlds, 26(3–4):347–355, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2015:SFC

[643] Xiaoyong Zhang and Shiguang Liu. SPH fluid control with self-adaptive turbulent details. Computer Animation and Virtual Worlds, 26(3–4):357–366, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Barbosa:2015:ACS

[644] Charles Welton Ferreira Barbosa, Yoshinori Dobashi, and Tsuyoshi Yamamoto. Adaptive cloud simulation using position based fluids. Computer Animation and Virtual Worlds, 26(3–4):367–375, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Berseth:2015:EOC

[645] Glen Berseth, Muhammad Usman, Brandon Haworth, Mubbasir Kapadia, and Petros Faloutsos. Environment optimization for crowd evacuation. Computer Animation and Virtual Worlds, 26(3–4):377–386, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wong:2015:GPS

[646] Sai-Keung Wong, Pao-Kun Tang, Fu-Shun Li, Zong-Min Wang, and Shih-Ting Yu. Guidance path scheduling using particle swarm optimization in crowd simulation. Computer Animation and Virtual Worlds, 26(3–4):387–395, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mao:2015:ELM

[647] Tianlu Mao, Hua Wang, Zhigang Deng, and Zhaoqi Wang. An efficient lane model for complex traffic simulation. Computer Animation and Virtual Worlds, 26(3–4):397–403, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chao:2015:VIM

[648] Qianwen Chao, Zhigang Deng, and Xiaogang Jin. Vehicle–pedestrian interaction for mixed traffic simulation. Computer Animation and Virtual Worlds, 26(3–4):405–412, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wong:2015:HBS

[649] Sai-Keung Wong and I-Ting Fu. Hybridbased snow simulation and snow rendering with shell textures. Computer Animation and Virtual Worlds, 26(3–4):413–421, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2015:ITM

[650] Yong-Joon Kim, Jong-Hwa Woo, Myung-Soo Kim, and Gershon Elber. Interactive tree modeling and deformation with collision detection and avoidance. Computer Animation and Virtual Worlds, 26(3–4):423–432, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2015:BDF

[651] Jianfang Li, Min Liu, Weiwei Xu, Haiyi Liang, and Ligang Liu. Boundarydominant flower blooming simulation. Computer Animation and Virtual Worlds, 26(3–4):433–443, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2015:GSB

[652] Namjung Kim, SaeWoon Oh, and Kyoungju Park. Giant soap bubble creation model. Computer Animation and Virtual Worlds, 26(3–4):445–455, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2015:CDS

[653] Kyungyong Yang, Kibeom Youn, Kyungho Lee, and Jehee Lee. Controllable data sampling in the space of human poses. Computer Animation and Virtual Worlds, 26(3–4):457–467, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kazmi:2015:ESB

[654] Ismail Khalid Kazmi, Lihua You, Xiaosong Yang, Xiaogang Jin, and Jian J. Zhang. Efficient sketch-based creation of detailed character models through data-driven mesh deformations. Computer Animation and Virtual Worlds, 26(3–4):469–481, May 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:IId

[655] Anonymous. Issue information. Computer Animation and Virtual Worlds, 26(5):i–ii, September 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:IIe

[656] Anonymous. Issue information. Computer Animation and Virtual Worlds, 26(5):i–ii, September 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Boatright:2015:GMP

[657] Cory D. Boatright, Mubbasir Kapadia, Jennie M. Shapira, and Norman I. Badler. Generating a multiplicity of policies for agent steering in crowd simulation. Computer Animation and Virtual Worlds, 26(5):483–494, September 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kumar:2015:AVB

[658] Amit Kumar and Aparajita Ojha. Anticipated velocity based guidance strategy for wheeled mobile evader amidst stationary and moving obstacles in bounded environment. Computer Animation and Virtual Worlds, 26(5): 495–507, September 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2015:MIU

[659] Rongkai Guo, Gayani Samaraweera, and John Quarles. Mobility impaired users respond differently than healthy users in virtual environments. Computer Animation and Virtual Worlds, 26(5): 509–526, September 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2015:HSD

[660] Jong-Hyun Kim, Chang-Hun Kim, and Jung Lee. A hybrid SDF for the detailed representation of liquid–solid mixed surfaces. Computer Animation and Virtual Worlds, 26(5):527–536, September 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2015:IIf

[661] Anonymous. Issue information. Computer Animation and Virtual Worlds, 26(6):i–ii, November 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zank:2015:PAS

[662] Markus Zank, Thomas Nescher, and Andreas Kunz. Predicting audio step feedback for real walking in virtual environments. Computer Animation and Virtual Worlds, 26(6):537–547, November 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kotsilieris:2015:IDV

[663] Theodore Kotsilieris, George T. Karetsos, Ioannis Anagnostopoulos, and Nikoletta A. Dimopoulou. Interconnecting distributed virtual worlds using Metabots: performance evaluation against the traditional client-server model. Computer Animation and Virtual Worlds, 26(6):549–561, November 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Way:2015:SAI

[664] Der-Lor Way and Cheng-Han Hsieh. 3D street art illusions: embedding chalk stylized rendering of 3D objects into a pavement photo. Computer Animation and Virtual Worlds, 26(6): 563–575, November 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2015:ETM

[665] Yuanfeng Zhou, Caiming Zhang, and Pengbo Bo. Efficient tetrahedral mesh generation based on sampling optimization. Computer Animation and Virtual Worlds, 26(6):577–587, November 2015. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2016:IIa

[666] Anonymous. Issue information. Computer Animation and Virtual Worlds, 27(1):1–2, January 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2016:CIB

[667] Liang Zhang and Guido Brunnett. Combining inverse blending and Jacobianbased inverse kinematics to improve accuracy in human motion generation. Computer Animation and Virtual Worlds, 27(1):3–13, January 2016. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2016:VPS

[668] Zhanpeng Huang, Guanghong Gong, and Liang Han. Vortex particle smoke simulation with an octree data structure. Computer Animation and Virtual Worlds, 27(1):14–23, January 2016. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lemercier:2016:TMB

[669] Samuel Lemercier and Jean-Michel Auberlet. Towards more behaviours in crowd simulation. Computer Animation and Virtual Worlds, 27(1):24–34, January 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Aristidou:2016:EFM

[670] Andreas Aristidou, Yiorgos Chrysanthou, and Joan Lasenby. Extending FABRIK with model constraints. Computer Animation and Virtual Worlds, 27(1):35–57, January 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Etemad:2016:PVA

[671] S. Ali Etemad, Ali Arya, Avi Parush, and Steve DiPaola. Perceptual validity in animation of human motion. Computer Animation and Virtual Worlds, 27(1):58–71, January 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2016:RHM

[672] Guang Chen, Jituo Li, Bei Wang, Jiping Zeng, Guodong Lu, and Dongliang Zhang. Reconstructing 3D human models with a Kinect. Computer Animation and Virtual Worlds, 27(1):72–85, January 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2016:IIb

[673] Anonymous. Issue information. Computer Animation and Virtual Worlds, 27(2):87–88, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gutierrez-Garcia:2016:CSA

[674] J. Octavio Gutierrez-Garcia and Luis-Felipe Rodríguez. Corruptible social agents. Computer Animation and Virtual Worlds, 27(2):89–102, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Deul:2016:PBR

[675] Crispin Deul, Patrick Charrier, and Jan Bender. Position-based rigid-body dynamics. Computer Animation and Virtual Worlds, 27(2):103–112, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2016:IDG

[676] Yong Sun Kim, Yongwan Kim, and Ki-Hong Kim. Interactive digital graffiti canvas system. Computer Animation and Virtual Worlds, 27(2):113–121, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lu:2016:ELD

[677] Daying Lu, Dengming Zhu, Zhaoqi Wang, and Jinzhu Gao. Efficient level of detail for texture-based flow visualization. Computer Animation and Virtual Worlds, 27(2):123–140, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tang:2016:ITE

[678] Wen Tang, Tao Ruan Wan, and Donjing Huang. Interactive thin elastic materials. Computer Animation and Virtual Worlds, 27(2):141–150, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vezzaro:2016:ICD

[679] Luca Vezzaro, Davide Zerbato, and Paolo Fiorini. Interactive constrained dynamics for rigid and deformable objects. Computer Animation and Virtual Worlds, 27(2):151–162, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2016:ARA

[680] Jun Lee, Mingyu Lim, SungJun Park, HyungSeok Kim, Heedong Ko, and Jee-In Kim. Approximate resolution of asynchronous conflicts among sequential collaborations in dynamic virtual environments. Computer Animation and Virtual Worlds, 27(2):163–180, March 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2016:IIc

[681] Anonymous. Issue information. Computer Animation and Virtual Worlds, 27(3–4):181–182, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2016:E

[682] Anonymous. Editorial. Computer Animation and Virtual Worlds, 27 (3–4):183–184, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2016:RTS

[683] Yanzhen Wang, Ferdinand Serracino-Inglott, Xiaodong Yi, Xue-Feng Yuan, and Xue-Jun Yang. Real-time simulation of catheterization in endovascular surgeries. Computer Animation and Virtual Worlds, 27(3–4):185–194, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Berseth:2016:ACB

[684] Glen Berseth, Mubbasir Kapadia, and Petros Faloutsos. ACCLMesh: curvature-based navigation mesh generation. Computer Animation and Virtual Worlds, 27(3–4):195–204, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2016:PBE

[685] Kang Hoon Lee and Myung Geol Choi.

A path browser for exploratory motion assembly. Computer Animation and Virtual Worlds, 27(3–4):205–212, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2016:PBU

[686] Myung Geol Choi and Kang Hoon Lee.

Points-based user interface for character posing. Computer Animation and Virtual Worlds, 27(3–4):213–220, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2016:DDD

[687] Chenlei Wu and Takashi Kanai. Datadriven detailed hair animation for game characters. Computer Animation and Virtual Worlds, 27(3–4):221–230, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xue:2016:EGC

[688] Junjie Xue, Gang Zhao, and Wenlei Xiao. An efficient GPU out-of-core framework for interactive rendering of large-scale CAD models. Computer Animation and Virtual Worlds, 27(3–4):231–240, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2016:DEP

[689] Ping Li and Hanqiu Sun. Densityenhanced perceptual mosaic on GPU. Computer Animation and Virtual Worlds, 27(3–4):241–249, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2016:ERS

[690] Guiqing Li, Yaobin Ouyang, Guodong Wei, Zhibang Zhang, and Aihua Mao. Enhanced rig-space simulation. Computer Animation and Virtual Worlds, 27(3–4):250–261, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chung:2016:PBD

[691] Se-Joon Chung and Nancy Pollard. Predictable behavior during contact simulation: a comparison of selected physics engines. Computer Animation and Virtual Worlds, 27(3–4):262–270, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2016:SFI

[692] Jong-Hyun Kim, Jaeho Im, Chang-Hun Kim, and Jung Lee. Subtle features of ice with cloudy effects and scratches from collision damage. Computer Animation and Virtual Worlds, 27(3–4):271–279, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qian:2016:EST

[693] Kun Qian, Tao Jiang, Meili Wang, Xiaosong Yang, and Jianjun Zhang. Energized soft tissue dissection in surgery simulation. Computer Animation and Virtual Worlds, 27(3–4):280–289, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2016:ICB

[694] Yanzhen Wang, Ferdinand Serracino-Inglott, Xiaodong Yi, Xue-Jun Yang, and Xue-Feng Yuan. An interactive computer-based simulation system for endovascular aneurysm repair surgeries. Computer Animation and Virtual Worlds, 27(3–4):290–300, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2016:RTF

[695] Shuang Liu, Xiaosong Yang, Zhao Wang, Zhidong Xiao, and Jianjun Zhang. Real-time facial expression transfer with single video camera. Computer Animation and Virtual Worlds, 27(3–4):301–310, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Senecal:2016:CBE

[696] Simon Senecal, Louis Cuel, Andreas Aristidou, and Nadia Magnenat-Thalmann. Continuous body emotion recognition system during theater performances. Computer Animation and Virtual Worlds, 27(3–4):311–320, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Laraba:2016:DPE

[697] Sohaib Laraba and Joëlle Tilmanne.

Dance performance evaluation using hidden Markov models. Computer Animation and Virtual Worlds, 27(3–4): 321–329, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kochanowicz:2016:SCC

[698] Jaroslaw Kochanowicz, Ah-Hwee Tan, and Daniel Thalmann. Social context cognition crowd-sourcing and semiautomatic parametrization. Computer Animation and Virtual Worlds, 27 (3–4):330–339, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2016:DFT

[699] Xin Yang, Qi Liu, Pengfei Zhang, Lutong Xin, Dongsheng Zhou, Yuxin Wang, Qiang Zhang, and Xiaopeng Wei. DKD: a fast k–d tree update design for dynamic scenes. Computer Animation and Virtual Worlds, 27(3–4):340–350, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Gao:2016:EVS

[700] Tianhao Gao, Wencheng Wang, and Honglei Han. Efficient view selection by measuring proxy information. Computer Animation and Virtual Worlds, 27(3–4):351–357, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Olivares:2016:ECB

[701] Ulises Olivares, Héctor G. Rodríguez, Arturo García, and Félix F. Ramos. Efficient construction of bounding volume hierarchies into a complete octree for ray tracing. Computer Animation and Virtual Worlds, 27(3–4):358–368, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2016:WTC

[702] Yingying Wang, Kerstin Ruhland, Michael Neff, and Carol O’Sullivan. Walk the talk: coordinating gesture with locomotion for conversational characters. Computer Animation and Virtual Worlds, 27(3–4):369–377, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Han:2016:ORT

[703] Daseong Han, Seokpyo Hong, Junyong Noh, Xiaogang Jin, and Joseph S. Shin. Online real-time locomotive motion transformation based on biomechanical observations. Computer Animation and Virtual Worlds, 27(3–4): 378–384, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2016:PIC

[704] Ning Liu, Dengming Zhu, Zhaoqi Wang, Hong Qin, Jianfeng Zhan, and Jinzhu Gao. Pipelining image compositing in heterogeneous networking environments. Computer Animation and Virtual Worlds, 27(3–4):385–393, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2016:PLV

[705] Ning Liu, Dengming Zhu, Zhaoqi Wang, Yi Wei, and Min Shi. Progressive light volume for interactive volumetric illumination. Computer Animation and Virtual Worlds, 27(3–4):394–404, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cui:2016:PBS

[706] Liqun Cui, Chunyong Ma, and Ge Chen.

Physical-based spatio-temporal resolution enhancement of scalar data for fluid visualization. Computer Animation and Virtual Worlds, 27(3–4):405–414, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2016:DDP

[707] Cheng Yang, Xubo Yang, and Xiangyun Xiao. Data-driven projection method in fluid simulation. Computer Animation and Virtual Worlds, 27(3–4):415–424, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Imai:2016:RTS

[708] Takuya Imai, Yoshihiro Kanamori, and Jun Mitani. Real-time screen-space liquid rendering with complex refractions. Computer Animation and Virtual Worlds, 27(3–4):425–434, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2016:ICB

[709] Myung Geol Choi and Kang Hoon Lee. Interactive control of big-object manipulation animation. Computer Animation and Virtual Worlds, 27(3–4):435–442, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2016:LVJ

[710] Jong-In Choi, Sun-Jeong Kim, Chang-Hun Kim, and Jung Lee. Let’s be a virtual juggler. Computer Animation and Virtual Worlds, 27(3–4):443–450, May 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2016:IId

[711] Anonymous. Issue information. Computer Animation and Virtual Worlds, 27 (5):451–453, September 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ma:2016:GCB

[712] Wei Ma, Yu Zhang, Luwei Yang, and Lijuan Duan. Graph-cut based interactive image segmentation with randomized texton searching. Computer Animation and Virtual Worlds, 27(5):

454–465, September 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wen:2016:FAW

[713] Laixiang Wen, Ning Xie, and Jinyuan Jia. Fast accessing Web3D contents using lightweight progressive meshes. Computer Animation and Virtual Worlds, 27(5):466–483, September 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Valcik:2016:ASM

[714] Jakub Valcik, Jan Sedmidubsky, and Pavel Zezula. Assessing similarity models for human-motion retrieval applications. Computer Animation and Virtual Worlds, 27(5):484–500, September 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Weiwei:2016:NMA

[715] Xing Weiwei, Wang Weiqiang, Bao Peng, Sun Liya, and Tong Leiming. A novel method for automated human behavior segmentation. Computer Animation and Virtual Worlds, 27(5): 501–514, September 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2016:IIe

[716] Anonymous. Issue information. Computer Animation and Virtual Worlds, 27 (6):515–517, November 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnenat-Thalmann:2016:EI

[717] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial issue 27.6. Computer Animation and Virtual Worlds, 27(6):518, November 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2016:IDA

[718] Mingfei Wang, Jinyuan Jia, Ning Xie, and Chenxi Zhang. Interestdriven avatar neighbor-organizing for P2P transmission in distributed virtual worlds. Computer Animation and Virtual Worlds, 27(6):519–531, November 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oshita:2016:CMS

[719] Masaki Oshita, Hayato Oshima, Yuta Senju, and Syun Morishige. Character motion synthesis by principal component analysis and motion control interface by hands. Computer Animation and Virtual Worlds, 27(6):532–545, November 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2016:TAM

[720] Xun Wang, Wenwu Yang, Wangbin Kou, Bailin Yang, and Guozheng Wang. Topology-aware moving least square deformation for 2D characters. Computer Animation and Virtual Worlds, 27 (6):546–555, November 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hajizadeh:2016:PCA

[721] Mohammadali Hajizadeh and Hossein Ebrahimnezhad. Predictive compression of animated 3D models by optimized weighted blending of key-frames. Computer Animation and Virtual Worlds, 27 (6):556–576, November 2016. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2017:IIa

[722] Anonymous. Issue information. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Boom:2017:ILS

[723] Bastiaan J. Boom, Sergio Orts-Escolano, Xin X. Ning, Steven McDonagh, Peter Sandilands, and Robert B. Fisher. Interactive light source position estimation for augmented reality with an RGB-D camera. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2017:PBH

[724] Congkun Chen, Yun Sheng, Fang Li, Guixu Zhang, and Hassan Ugail. A PDE-based head visualization method with CT data. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Flores:2017:SBS

[725] Jorge Eduardo Ramírez Flores and Antonio Susín Sánchez. Segmentationbased skinning. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lacle:2017:SGM

[726] Francis Laclé and Nicolas Pronost. A scalable geometrical model for musculotendon units. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pino:2017:MFP

[727] Angel Zaldivar Pino, Manuel Gonzalez Bedia, and Francisco José Serón Arbeloa. Modeling flocks with perceptual agents from a dynamicist perspective. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2017:EIa

[728] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 28.1. Computer Animation and Virtual Worlds, 28(1):??, January 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2017:IIb

[729] Anonymous. Issue information. Computer Animation and Virtual Worlds, 28(2):e1735:1–e1735:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2017:EIb

[730] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 28.2. Computer Animation and Virtual Worlds, 28(2):e1753:1–e1753:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2017:IPM

[731] Jong-Hyun Kim, Chang-Hun Kim, and Jung Lee. Incorporating particle motion into an ADF for fast coupling of fluids with rigid and deformable solids. Computer Animation and Virtual Worlds, 28(2):e1689:1–e1689:??, March– April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baravalle:2017:RMP

[732] Rodrigo Baravalle, Leonardo Scandolo, Claudio Delrieux, Cristian García Bauza, and Elmar Eisemann. Realistic modeling of porous materials. Computer Animation and Virtual Worlds, 28(2): e1719:1–e1719:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Possani-Espinosa:2017:DPT

[733] Andre Possani-Espinosa, J. Octavio Gutierrez-Garcia, and Isaac Vargas Gordillo. Determining personality traits of racing game players using the open racing car simulator: toward believable virtual drivers. Computer Animation and Virtual Worlds, 28(2):e1722:1– e1722:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Alsweis:2017:SVA

[734] Monssef Alsweis, Oliver Deussenn, and Jia Liu. Simulation and visualization of adapting venation patterns. Computer Animation and Virtual Worlds, 28(2): e1723:1–e1723:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qian:2017:ETL

[735] Kun Qian, Junxuan Bai, Xiaosong Yang, Junjun Pan, and Jianjun Zhang. Essential techniques for laparoscopic surgery simulation. Computer Animation and Virtual Worlds, 28(2):e1724:1– e1724:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cui:2017:IBE

[736] Dele Cui, Yun Sheng, and Guixu Zhang. Image-based embroidery modeling and rendering. Computer Animation and Virtual Worlds, 28(2):e1725:1– e1725:??, March–April 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2017:E

[737] Anonymous. Editorial. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2017:IIc

[738] Anonymous. Issue information. Computer Animation and Virtual Worlds, 28 (3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Baek:2017:STL

[739] Seungho Baek and JungHyun Han. Simulation of thin liquid jets with threads. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Balint:2017:AAL

[740] J. Timothy Balint and Jan Allbeck. ALET: Agents Learning their Environment through Text. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cai:2017:DCF

[741] Zhongxuan Cai, Xuefeng Chang, Yanzhen Wang, Xiaodong Yi, and Xue-Jun Yang. Distributed control for flocking and group maneuvering of nonholonomic agents. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Feng:2017:JTV

[742] Andrew Feng, Evan Suma Rosenberg, and Ari Shapiro. Just-in-time, viable, 3-D avatars from scans. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2017:SCT

[743] Shihui Guo, Meili Wang, Gabriel Notman, Jian Chang, Jianjun Zhang, and Minghong Liao. Simulating collective transport of virtual ants. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Haworth:2017:DFR

[744] Brandon Haworth, Muhammad Usman, Glen Berseth, Mubbasir Kapadia, and Petros Faloutsos. On densityflow relationships during crowd evacuation. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Herrmann:2017:ASH

[745] Erik Herrmann, Martin Manns, Han Du, Somayeh Hosseini, and Klaus Fischer. Accelerating statistical human motion synthesis using space partitioning data structures. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Im:2017:VSR

[746] Jaeho Im, Jong-Hyun Kim, Wook Kim, Nuri Park, Taehyeong Kim, Young Bin Kim, Jung Lee, and Chang-Hun Kim. Visual simulation of rapidly freezing water based on crystallization. Computer Animation and Virtual Worlds, 28(3–4): ??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jung:2017:ARG

[747] Sunjin Jung, Seokpyo Hong, Kyungmin Cho, Haegwang Eom, Byungkuk Choi, and Junyong Noh. Age-related gait motion transformation based on biomechanical observations. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kang:2017:SIH

[748] Sin-Hwa Kang, David M. Krum, Peter Khooshabeh, Thai Phan, Chien-Yen Chang, Ori Amir, and Rebecca Lin. Social influence of humor in virtual human counselor’s self-disclosure. Computer Animation and Virtual Worlds, 28(3–4): ??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Khorloo:2017:CAS

[749] Oyundolgor Khorloo and Enkhbayar Altantsetseg. Constructive approach for smoke plume animation using turbulent toroidal vortices. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2017:EVH

[750] Kangsoo Kim, Divine Maloney, Gerd Bruder, Jeremy N. Bailenson, and Gregory F. Welch. The effects of virtual human’s spatial and behavioral coherence with physical objects on social presence in AR. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kravchenko:2017:HFI

[751] Boris Kravchenko, Gladimir V. G. Baranoski, Tenn Francis Chen, Erik Miranda, and Spencer R. Van Leeuwen. Highfidelity iridal light transport simulations at interactive rates. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Laraba:2017:SBA

[752] Sohaib Laraba, Mohammed Brahimi, Joëlle Tilmanne, and Thierry Dutoit. 3D skeleton-based action recognition by representing motion capture sequences as 2D-RGB images. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2017:MIB

[753] Youjin Lee, Sukwon Lee, and Sung-Hee Lee. Multifinger interaction between remote users in avatar-mediated telepresence. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2017:MMB

[754] Jiwon Lee, Kisung Jeong, and Jinmo Kim. MAVE: Maze-based immersive virtual environment for new presence and experience. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017.

CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2017:FBR

[755] Xiangyun Liao, Weixin Si, Biao Xu, Zhiyong Yuan, Qiong Wang, and Pheng-Ann Heng. Filament-based realistic turbulent wake synthesis. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2017:SDT

[756] Chunxiao Liu, Yiyun Shen, Yaqi Shao, Jinwei Zhao, and Xun Wang. Sky detection- and texture smoothing-based high-visibility haze removal from images and videos. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Narang:2017:MRS

[757] Sahil Narang, Andrew Best, Andrew Feng, Sin hwa Kang, Dinesh Manocha, and Ari Shapiro. Motion recognition of self and others on realistic 3D avatars. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sato:2017:FCF

[758] Syuhei Sato, Keisuke Mizutani, Yoshinori Dobashi, Tomoyuki Nishita, and Tsuyoshi Yamamoto. Feedback control of fire simulation based on computational fluid dynamics. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tisserand:2017:AGP

[759] Yvain Tisserand, Louis Cuel, and Nadia Magnenat-Thalmann. Automatic 3D garment positioning based on surface metric. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vermeulen:2017:CSK

[760] Jordi L. Vermeulen, Arne Hillebrand, and Roland Geraerts. A comparative study of k-nearest neighbour techniques in crowd simulation. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2017:TCL

[761] Yutong Wang, Luyuan Wang, Zhigang Deng, and Xiaogang Jin. Topologically consistent leafy tree morphing. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yumak:2017:ASG

[762] Zerrin Yumak, Bram van den Brink, and Arjan Egges. Autonomous social gaze model for an interactive virtual character in real-life settings. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yun:2017:PII

[763] Jeongsu Yun, Myungbae Son, Byungyoon Choi, Theodore Kim, and Sung-Eui Yoon. Physically inspired, interactive lightning generation. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2017:HMM

[764] Shenfan Zhang, Fanlong Kong, Chen Li, Changbo Wang, and Hong Qin. Hybrid modeling of multiphysical processes for particle-based volcano animation. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2017:SCD

[765] Yongqiang Zhang, Shuang Liu, Xiaosong Yang, Jianjun Zhang, and Daming Shi. Supervised coordinate descent method with a 3D bilinear model for face alignment and tracking. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2017:BCS

[766] Xiaoqiang Zhu, Lei Song, Lihua You, Mengyao Zhu, Xiangyang Wang, and Xiaogang Jin. Brush2Model: Convolution surface-based brushes for 3D modelling in head-mounted display-based virtual environments. Computer Animation and Virtual Worlds, 28(3–4):??, May 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2017:IId

[767] Anonymous. Issue information. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bulbul:2017:PVC

[768] Abdullah Bulbul and Rozenn Dahyot.

Populating virtual cities using social media. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kyriakou:2017:IVC

[769] Marios Kyriakou, Xueni Pan, and Yiorgos Chrysanthou. Interaction with virtual crowd in Immersive and semi-Immersive Virtual Reality systems. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liang:2017:EMI

[770] Hui Liang, Jian Chang, Shujie Deng, Can Chen, Ruofeng Tong, and Jian Jun Zhang. Exploitation of multiplayer interaction and development of virtual puppetry storytelling using gesture control and stereoscopic devices. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Polceanu:2017:CMS

[771] Mihai Polceanu and Cédric Buche. Computational mental simulation: a review. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qiu:2017:NFD

[772] Yuxing Qiu, Lipeng Yang, Shuai Li, Qing Xia, Hong Qin, and Aimin Hao. Novel fluid detail enhancement based on multi-layer depth regression analysis and FLIP fluid simulation. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Stuvel:2017:PCB

[773] Sybren A. Stüvel, A. Frank van der Stappen, and Arjan Egges. Perception of collisions between virtual characters. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2017:EIc

[774] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 28.5. Computer Animation and Virtual Worlds, 28(5):??, September 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2017:IIe

[775] Anonymous. Issue information. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2017:IAB

[776] Kan Chen and Henry Johan. Interactive authoring of bending and twisting motions of short plants using hand gestures. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2017:LCR

[777] Jiaxu Chen, Long Zhang, Xiaoxu Li, Bo Zhang, and Zhongfu Ye. Locally controlled as-rigid-as-possible deformation for 2D characters. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Haworth:2017:CCO

[778] Brandon Haworth, Muhammad Usman, Glen Berseth, Mahyar Khayatkhoei, Mubbasir Kapadia, and Petros Faloutsos. CODE: Crowd-optimized design of environments. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2017:RHC

[779] Jong-Hyun Kim, Wook Kim, Young Bin Kim, Jaeho Im, Jung Lee, and Sun-Jeong Kim. Robust handling of clumping and stiffness in wet hair animation. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Pino:2017:PML

[780] Angel Zaldivar Pino, Manuel Gonzalez Bedia, and Francisco José Seron Arbeloa. A parsimonious model for locomotor in virtual agents based on dynamical coupling with the environment. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ruiz:2017:SBC

[781] Ana Lucia Cruz Ruiz, Charles Pontonnier, Jonathan Levy, and Georges Dumont. A synergy-based control solution for overactuated characters: Application to throwing. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2017:EId

[782] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 28.6. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2017:RVF

[783] Xin Yang, Wanchao Su, Jian Deng, Xiaogang Jin, Guozhen Tan, and Zhigeng Pan. Real-virtual fusion model for traffic animation. Computer Animation and Virtual Worlds, 28(6):??, November 2017. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2018:IIa

[784] Anonymous. Issue information. Computer Animation and Virtual Worlds, 29 (1):e1789:1–e1789:??, January–February 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2018:EIa

[785] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 29.1. Computer Animation and Virtual Worlds, 29 (1):e1803:1–e1803:??, January–February 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nevelsteen:2018:VWD

[786] Kim J. L. Nevelsteen. Virtual world, defined from a technological perspective and applied to video games, mixed reality, and the Metaverse. Computer Animation and Virtual Worlds, 29 (1):e1752:1–e1752:??, January–February 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tsai:2018:DVC

[787] Tsung-Yu Tsai, Sai-Keung Wong, Yi-Hung Chou, and Guan-Wen Lin. Directing virtual crowds based on dynamic adjustment of navigation fields. Computer Animation and Virtual Worlds, 29 (1):e1765:1–e1765:??, January–February 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Feng:2018:DPS

[788] Gang Feng and Shiguang Liu. Detailpreserving SPH fluid control with deformation constraints. Computer Animation and Virtual Worlds, 29(1):e1781:1– e1781:??, January–February 2018. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kwon:2018:HVD

[789] Soonhyeon Kwon, Younguk Kim, Kihyuk Kim, and Sungkil Lee. Heterogeneous volume deformation and animation authoring with density-aware moving least squares. Computer Animation and Virtual Worlds, 29(1):e1784:1– e1784:??, January–February 2018. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chan:2018:PMC

[790] Ka-Hou Chan, Wei Ke, and Sio-Kei Im.

Particle-mesh coupling in the interaction of fluid and deformable bodies with screen space refraction rendering. Computer Animation and Virtual Worlds, 29 (1):e1787:1–e1787:??, January–February 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shao:2018:PBS

[791] Xuqiang Shao, Wei Wu, and Baoyi Wang. Position-based simulation of cloth wetting phenomena. Computer Animation and Virtual Worlds, 29 (1):e1788:1–e1788:??, January–February 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2018:CI

[792] Xiaoyong Zhang and Shiguang Liu. Cover image. Computer Animation and Virtual Worlds, 29(2):i, March–April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2018:IIb

[793] Anonymous. Issue information. Computer Animation and Virtual Worlds, 29(2):e1790:1–e1790:??, March–April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2018:EIb

[794] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 29.2. Computer Animation and Virtual Worlds, 29(2):e1809:1–e1809:??, March–April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Waszak:2018:LMS

[795] Bartlomiej Waszak. Limbless movement simulation with a particle-based system. Computer Animation and Virtual Worlds, 29(2):e1795:1–e1795:??, March– April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jiang:2018:EBD

[796] Min Jiang, Richard Southern, and Jiang J. Zhang. Energy-based dissolution simulation using SPH sampling. Computer Animation and Virtual Worlds, 29(2):e1798:1–e1798:??, March– April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Dhiman:2018:IAP

[797] Ashish Dhiman, Dhaval Solanki, Ashu Bhasin, Abhijit Das, and Uttama Lahiri. An intelligent, adaptive, performancesensitive, and virtual reality-based gaming platform for the upper limb. Computer Animation and Virtual Worlds, 29(2):e1800:1–e1800:??, March–April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2018:PSF

[798] Xiaoyong Zhang and Shiguang Liu. Parallel SPH fluid control with dynamic details. Computer Animation and Virtual Worlds, 29(2):e1801:1–e1801:??, March– April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Parhi:2018:NMH

[799] Dayal R. Parhi, Chinmaya Sahu, and Priyadarshi Biplab Kumar. Navigation of multiple humanoid robots using hybrid adaptive swarm-adaptive ant colony optimisation technique. Computer Animation and Virtual Worlds, 29(2):e1802:1–e1802:??, March–April 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lu:2018:CIV

[800] Yao Lu, Shang Zhao, Naji Younes, and James K. Hahn. Cover image, volume 29, issue 5. Computer Animation and Virtual Worlds, 29(5):e1863:1– e1863:??, September–October 2018. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2018:IIc

[801] Anonymous. Issue information. Computer Animation and Virtual Worlds, 29(5):e1793:1–e1793:??, September– October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2018:EIc

[802] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 29.5. Computer Animation and Virtual Worlds, 29(5):e1862:1–e1862:??, September– October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kyrlitsias:2018:ACE

[803] Christos Kyrlitsias and Despina Michael-Grigoriou. Asch conformity experiment using immersive virtual reality. Computer Animation and Virtual Worlds, 29(5):e1804:1–e1804:??, September–October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2018:SAV

[804] Yu Zhu, Shiying Li, Xi Luo, Kang Zhu, Qiang Fu, Xilin Chen, Huixing Gong, and Jingyi Yu. A shared augmented virtual environment for real-time mixed reality applications. Computer Animation and Virtual Worlds, 29(5):e1805:1– e1805:??, September–October 2018. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Agil:2018:GBA

[805] Umut Agil and Ugur Güdükbay. A group-based approach for gaze behavior of virtual crowds incorporating personalities. Computer Animation and Virtual Worlds, 29(5):e1806:1–e1806:??, September–October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lu:2018:ANH

[806] Yao Lu, Shang Zhao, Naji Younes, and James K. Hahn. Accurate nonrigid 3D human body surface reconstruction using commodity depth sensors. Computer Animation and Virtual Worlds, 29(5): e1807:1–e1807:??, September–October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2018:DCC

[807] Wanrong Huang, Yanzhen Wang, Xiaodong Yi, and Xue-Jun Yang. Distributed coordination with connectivity maintenance for nonholonomic robots. Computer Animation and Virtual Worlds, 29(5):e1856:1–e1856:??, September–October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2018:EBS

[808] Kai Wang and Shiguang Liu. Examplebased synthesis for sound of ocean waves caused by bubble dynamics. Computer Animation and Virtual Worlds, 29(3–4):e1857:1–e1857:??, September– October 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2018:IId

[809] Anonymous. Issue information. Computer Animation and Virtual Worlds, 29 (6):e1794:1–e1794:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Plaku:2018:MGM

[810] Erion Plaku, Sara Rashidian, and Stefan Edelkamp. Multi-group motion planning in virtual environments. Computer Animation and Virtual Worlds, 29(6): e1688:1–e1688:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Rabbani:2018:ABC

[811] Amir H. Rabbani, Michiel van de Panne, and Paul G. Kry. Anticipatory balance control and dimension reduction. Computer Animation and Virtual Worlds, 29(6):e1726:1–e1726:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oshita:2018:GHM

[812] Masaki Oshita and Yuta Senju. Generating hand motion from body motion based on hand pose estimation. Computer Animation and Virtual Worlds, 29(6):e1730:1–e1730:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Normoyle:2018:EAC

[813] Aline Normoyle and Sophie Jörg. The effect of animation controller and avatar on player perceptions. Computer Animation and Virtual Worlds, 29(6), November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Krejtz:2018:PES

[814] Krzysztof Krejtz, Andrew Duchowski, Heng Zhou, Sophie Jörg, and Anna Niedzielska. Perceptual evaluation of synthetic gaze jitter. Computer Animation and Virtual Worlds, 29(6): e1745:1–e1745:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kavafoglu:2018:RSC

[815] Zumra Kavafoglu, Ersan Kavafoglu, and Arjan Egges. Robust standing control with posture optimization. Computer Animation and Virtual Worlds, 29(6): e1746:1–e1746:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Schroder:2018:DER

[816] Matthias Schröder, Thomas Waltemate, Jonathan Maycock, Tobias Röhlig, Helge Ritter, and Mario Botsch. Design and evaluation of reduced marker layouts for hand motion capture. Computer Animation and Virtual Worlds, 29(6): e1751:1–e1751:??, November–December 2018. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2019:IIa

[817] Anonymous. Issue information. Computer Animation and Virtual Worlds, 30 (1):e1849:1–e1849:??, January–February 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cevikbas:2019:PVB

[818] Safak Burak Çevikbas and Veysi Isler.

Phaneros: Visibility-based framework for massive peer-to-peer virtual environments. Computer Animation and Virtual Worlds, 30(1):e1808:1–e1808:??, January–February 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2019:DDS

[819] Xuejin Chen, Haoming Jiang, Tingting Xuan, Lihan Huang, and Ligang Liu. Designing deployable 3D scissor structures with ball-and-socket joints. Computer Animation and Virtual Worlds, 30 (1):e1848:1–e1848:??, January–February 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Dupre:2019:HGB

[820] Rob Dupre and Vasileios Argyriou. A human and group behavior simulation evaluation framework utilizing composition and video analysis. Computer Animation and Virtual Worlds, 30 (1):e1844:1–e1844:??, January–February 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Luo:2019:ABE

[821] Yanlin Luo, Bin Gao, Yiyi Deng, Xiaoming Zhu, Tianzi Jiang, Xudong Zhao, and Zhengyi Yang. Automated brain extraction and immersive exploration of its layers in virtual reality for the rhesus macaque MRI data sets. Computer Animation and Virtual Worlds, 30 (1):e1841:1–e1841:??, January–February 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2019:EIa

[822] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 30.1. Computer Animation and Virtual Worlds, 30 (1):e1872:1–e1872:??, January–February 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tian:2019:MGF

[823] Ye Tian, Yong Hu, and Xukun Shen. A multi-GPU finite element computation and hybrid collision handling process framework for brain deformation simulation. Computer Animation and Virtual Worlds, 30(1):e1846:1– e1846:??, January–February 2019. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2019:IIb

[824] Anonymous. Issue information. Computer Animation and Virtual Worlds, 30(2):e1850:1–e1850:??, March–April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kumar:2019:INH

[825] Priyadarshi Biplab Kumar, Saktiswarup Mohapatra, and Dayal R. Parhi. An intelligent navigation of humanoid NAO in the light of classical approach and computational intelligence. Computer Animation and Virtual Worlds, 30(2): e1858:1–e1858:??, March–April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2019:DME

[826] Mingyuan Li, Xiaoheng Jiang, Ningbo Gu, Weiwei Xu, Junxiao Xue, Bing Zhou, and Mingliang Xu. Disassembling a 3D mechanism for efficient packing. Computer Animation and Virtual Worlds, 30(2):e1861:1–e1861:??, March– April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ozcan:2019:TBG

[827] Cumhur Yigit Ozcan, Ebru Akcapinar Sezer, and Murat Haciomeroglu. A timebased global path planning strategy for crowd navigation. Computer Animation and Virtual Worlds, 30(2):e1864:1– e1864:??, March–April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Podila:2019:APP

[828] Sahithi Podila and Ying Zhu. Animating predator and prey fish interactions. Computer Animation and Virtual Worlds, 30(2):e1866:1–e1866:??, March– April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2019:EIb

[829] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 30.2. Computer Animation and Virtual Worlds, 30(2):e1876:1–e1876:??, March–April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vanakittistien:2019:GRHa

[830] Nuttapon Vanakittistien, Attawith Sudsang, and Nuttapong Chentanez. Gameready 3D hair model from a small set of images. Computer Animation and Virtual Worlds, 30(2):e1855:1– e1855:??, March–April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Vanakittistien:2019:GRHb

[831] Nuttapon Vanakittistien, Attawith Sudsang, and Nuttapong Chentanez. Gameready 3D hair model from a small set of images. Computer Animation and Virtual Worlds, 30(2):e1855:1– e1855:??, March–April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2019:VSO

[832] Lin Wang, Hyuncheol Kim, Imgyu Kim, and Soonhung Han. A visual simulation of ocean floating wind power system. Computer Animation and Virtual Worlds, 30(2):e1859:1–e1859:??, March– April 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2019:IIc

[833] Anonymous. Issue information. Computer Animation and Virtual Worlds, 30(3–4):e1860:1–e1860:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cebeci:2019:CSA

[834] Berk Cebeci, Ufuk Celikcan, and Tolga K. Capin. A comprehensive study of the affective and physiological responses induced by dynamic virtual reality environments. Computer Animation and Virtual Worlds, 30(3–4):e1893:1– e1893:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Charalambous:2019:ADE

[835] Constantinos Charalambous, Zerrin Yumak, and A. Frank van der Stappen. Audio-driven emotional speech animation for interactive virtual characters. Computer Animation and Virtual Worlds, 30(3–4):e1892:1–e1892:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2019:SCF

[836] Qiang Chen, Guoliang Luo, Yang Tong, Xiaogang Jin, and Zhigang Deng. Shape-constrained flying insects animation. Computer Animation and Virtual Worlds, 30(3–4):e1902:1–e1902:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Choi:2019:OCV

[837] Myung Geol Choi, Ji Hye Lee, Wansu Ha, and Kang Hoon Lee. Optimal close-up views for precise 3D manipulation. Computer Animation and Virtual Worlds, 30(3–4):e1884:1–e1884:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ecormier-Nocca:2019:IBA

[838] Pierre Ecormier-Nocca, Julien Pettré, Pooran Memari, and Marie-Paule Cani. Image-based authoring of herd animations. Computer Animation and Virtual Worlds, 30(3–4):e1903:1–e1903:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Favaretto:2019:ICA

[839] Rodolfo Migon Favaretto, Roberto Rosa dos Santos, Soraia Raupp Musse, Felipe Vilanova, and Angelo Brandelli Costa. Investigating cultural aspects in the fundamental diagram using convolutional neural networks and virtual agent simulation. Computer Animation and Virtual Worlds, 30(3–4):e1899:1– e1899:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hou:2019:MFD

[840] Yong Hou, Hangzai Luo, Wanqing Zhao, Xiang Zhang, Jun Wang, and Jinye Peng. Multilayer feature descriptors fusion CNN models for fine-grained visual recognition. Computer Animation and Virtual Worlds, 30(3–4):e1897:1– e1897:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2019:CI

[841] Zhongyuan Hu, Haoran Xie, Tsukasa Fukusato, Takahiro Sato, and Takeo Igarashi. Cover image. Computer Animation and Virtual Worlds, 30(3– 4):e1912:1–e1912:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2019:SSB

[842] Zhongyuan Hu, Haoran Xie, Tsukasa Fukusato, Takahiro Sato, and Takeo Igarashi. Sketch2VF: Sketch-based flow design with conditional generative adversarial network. Computer Animation and Virtual Worlds, 30(3–4):e1889:1– e1889:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2019:ICI

[843] Jiawei Huang and Shiguang Liu. Inside cover image. Computer Animation and Virtual Worlds, 30(3–4):e1913:1– e1913:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2019:RSL

[844] Jiawei Huang and Shiguang Liu. Robust simultaneous localization and mapping in low-light environment. Computer Animation and Virtual Worlds, 30(3– 4):e1895:1–e1895:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Je:2019:CAS

[845] Seungwoo Je, Yekaterina Abileva, Andrea Bianchi, and Jean-Charles Bazin. A computational approach for spider webinspired fabrication of string art. Computer Animation and Virtual Worlds, 30(3–4):e1904:1–e1904:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jin:2019:SSR

[846] Yuxi Jin, Ping Li, Bin Sheng, Yongwei Nie, Jinman Kim, and Enhua Wu. SRNPD: Spatial rendering network for pencil drawing stylization. Computer Animation and Virtual Worlds, 30(3– 4):e1890:1–e1890:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Krogmeier:2019:HVC

[847] Claudia Krogmeier, Christos Mousas, and David Whittinghill. Human-virtual character interaction: Toward understanding the influence of haptic feedback. Computer Animation and Virtual Worlds, 30(3–4):e1883:1–e1883:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

LeNaour:2019:SMA

[848] T. Le Naour, N. Courty, and S. Gibet. Skeletal mesh animation driven by few positional constraints. Computer Animation and Virtual Worlds, 30(3– 4):e1900:1–e1900:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2019:MCD

[849] Ping Li, Yuxi Jin, Bin Sheng, Di Lin, Yongwei Nie, and Enhua Wu. Multiview-coherent disocclusion synthesis using connected regions optimization. Computer Animation and Virtual Worlds, 30(3–4):e1894:1–e1894:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2019:IPR

[850] Kuan-Yu Liu, Matias Volonte, Yu-Chun Hsu, Sabarish V. Babu, and Sai-Keung Wong. Interaction with proactive and reactive agents in box manipulation tasks in virtual environments. Computer Animation and Virtual Worlds, 30(3– 4):e1881:1–e1881:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Nunnari:2019:YAL

[851] Fabrizio Nunnari and Alexis Heloir. Yet another low-level agent handler. Computer Animation and Virtual Worlds, 30(3–4):e1891:1–e1891:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Oshita:2019:ANU

[852] Masaki Oshita. Agent navigation using deep learning with agent space heat map for crowd simulation. Computer Animation and Virtual Worlds, 30(3– 4):e1878:1–e1878:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Peng:2019:DDR

[853] Chen Peng, Zipeng Zhao, Chen Li, Changbo Wang, Hong Qin, and Hongyan Quan. Data-driven retrieval of spray details with random forestbased distance. Computer Animation and Virtual Worlds, 30(3–4):e1901:1– e1901:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sun:2019:SAC

[854] Manhui Sun, Shaowu Yang, and Henzhu Liu. Scale-aware camera localization in 3D LiDAR maps with a monocular visual odometry. Computer Animation and Virtual Worlds, 30(3–4):e1879:1– e1879:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2019:E

[855] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. Computer Animation and Virtual Worlds, 30(3– 4):e1911:1–e1911:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2019:VTL

[856] Bo-Jyun Wu, Sai-Keung Wong, and Ta-Wei Li. Virtual titration laboratory experiment with differentiated instruction. Computer Animation and Virtual Worlds, 30(3–4):e1882:1–e1882:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xie:2019:MLM

[857] Xueguang Xie, Xiao Zhai, Fei Hou, Aimin Hao, and Hong Qin. Multitask learning on monocular water images: Surface reconstruction and image synthesis. Computer Animation and Virtual Worlds, 30(3–4):e1896:1– e1896:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2019:CND

[858] Xin Yang, Haoran Wang, Shaozhe Chen, Xinglin Piao, Dongsheng Zhou, Qiang Zhang, Baocai Yin, and Xiaopeng Wei. Cascaded network with deep intensity manipulation for scene understanding. Computer Animation and Virtual Worlds, 30(3–4):e1888:1– e1888:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2019:OWW

[859] Zhongyao Yang and Shiguang Liu. An optimization on water wave diffraction approximation based on wave packets. Computer Animation and Virtual Worlds, 30(3–4):e1886:1–e1886:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yiannakides:2019:RTH

[860] Anastasios Yiannakides, Andreas Aristidou, and Yiorgos Chrysanthou. Realtime 3D human pose and motion reconstruction from monocular RGB videos. Computer Animation and Virtual Worlds, 30(3–4):e1887:1–e1887:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2019:CAM

[861] Xun Zhang, Davide Schaumann, Brandon Haworth, Petros Faloutsos, and Mubbasir Kapadia. Coupling agent motivations and spatial behaviors for authoring multiagent narratives. Computer Animation and Virtual Worlds, 30(3–4):e1898:1–e1898:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2019:SAO

[862] Xiaoqiang Zhu, Lei Song, Nan Wang, Ruiheng Zhang, Shenshuai Chen, Xiangyang Wang, Mengyao Zhu, Lihua You, Zhigang Deng, and Xiaogang Jin. Screwing assembly oriented interactive model segmentation in HMD VR environment. Computer Animation and Virtual Worlds, 30(3–4):e1880:1– e1880:??, May–August 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2019:IIe

[863] Anonymous. Issue information. Computer Animation and Virtual Worlds, 30(5):e1853:1–e1853:??, September– October 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ozgen:2019:SMS

[864] Oktar Ozgen, Marcelo Kallmann, and Eric Brown. An SPH model to simulate the dynamic behavior of shear thickening fluids. Computer Animation and Virtual Worlds, 30(5):e1870:1– e1870:??, September–October 2019. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2019:ACS

[865] Huayue Wu. An adaptive character skinning algorithm based on the property of real skin. Computer Animation and Virtual Worlds, 30(5):e1868:1– e1868:??, September–October 2019. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiang:2019:BIA

[866] Wei Xiang, Jiaping Ren, Kuan Wang, Zhigang Deng, and Xiaogang Jin. Biologically inspired ant colony simulation. Computer Animation and Virtual Worlds, 30(5):e1867:1–e1867:??, September–October 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2019:BHS

[867] Xi Zhao, Zhenqiang Su, and Xinyu Yang. Building hierarchical structures for 3D scenes based on normalized cut. Computer Animation and Virtual Worlds, 30(5):e1869:1–e1869:??, September–October 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2019:IIf

[868] Anonymous. Issue information. Computer Animation and Virtual Worlds, 30 (6):e1854:1–e1854:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chan:2019:GFE

[869] Jacky C. P. Chan, Hubert P. H. Shum, He Wang, Li Yi, Wei Wei, and Edmond S. L. Ho. A generic framework for editing and synthesizing multimodal data with relative emotion strength. Computer Animation and Virtual Worlds, 30(6):e1871:1–e1871:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Huang:2019:PBF

[870] Yuhang Huang, Yonghang Yu, and Takashi Kanai. Predicting brittle fracture surface shape from a versatile database. Computer Animation and Virtual Worlds, 30(6):e1865:1–e1865:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Jiang:2019:LGW

[871] Quan Jiang and Xiliang Chen. Landslidegenerated wave hazard prediction based on multiphase flow model of Dual-SPHysics. Computer Animation and Virtual Worlds, 30(6):e1874:1–e1874:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2019:HPB

[872] Na Liu, Xingce Wang, Shaolong Liu, Zhongke Wu, Jiale He, Peng Cheng, Chunyan Miao, and Nadia Magnenat Thalmann. Hierarchical planningbased crowd formation. Computer Animation and Virtual Worlds, 30(6): e1875:1–e1875:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ronchi:2019:MPH

[873] Enrico Ronchi, David Mayorga, Ruggiero Lovreglio, Jonathan Wahlqvist, and Daniel Nilsson. Mobile-powered head-mounted displays versus cave automatic virtual environment experiments for evacuation research. Computer Animation and Virtual Worlds, 30(6): e1873:1–e1873:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2019:EIc

[874] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial Issue 30.6. Computer Animation and Virtual Worlds, 30 (6):e1918:1–e1918:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2019:CDM

[875] Bailin Yang, Zhaoyi Jiang, Jiantao Shangguan, Frederick W. B. Li, Chao Song, Yibo Guo, and Mingliang Xu. Compressed dynamic mesh sequence for progressive streaming. Computer Animation and Virtual Worlds, 30(6): e1847:1–e1847:??, November–December 2019. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2020:IIa

[876] Anonymous. Issue information. Computer Animation and Virtual Worlds, 31 (1):e1905:1–e1905:??, January–February 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2020:EIa

[877] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 31.1. Computer Animation and Virtual Worlds, 31 (1):e1921:1–e1921:??, January–February 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiao:2020:LRD

[878] Zhang Xiao and Yang Deling. Literature review: the distributed postproduction of cultural knowledge for artworks in online museums. Computer Animation and Virtual Worlds, 31(1):e1877:1– e1877:??, January–February 2020. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bin:2020:MCD

[879] Wang Bin, Liu Weibin, and Xing Weiwei. Motion capture data segmentation using Riemannian manifold learning. Computer Animation and Virtual Worlds, 31(1):e1885:1–e1885:??, January–February 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Demsar:2020:HMS

[880] Jure Demsar, Will Blewitt, and Iztok Lebar Bajec. A hybrid model for simulating grazing herds in real time. Computer Animation and Virtual Worlds, 31(1):e1914:1–e1914:??, January–February 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cao:2020:ISD

[881] Wei Cao, Zhixin Yang, Xiaohua Ren, Luan Lyu, Bob Zhang, Yanci Zhang, and Enhua Wu. An improved solution for deformation simulation of nonorthotropic geometric models. Computer Animation and Virtual Worlds, 31(1):e1915:1– e1915:??, January–February 2020. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cui:2020:CI

[882] Xiaoyu Cui, Ruifan Cai, Xiangjun Tang, Zhigang Deng, and Xiaogang Jin. Cover image. Computer Animation and Virtual Worlds, 31(2):e1926:1– e1926:??, March–April 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2020:IIb

[883] Anonymous. Issue information. Computer Animation and Virtual Worlds, 31(2):e1906:1–e1906:??, March–April 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2020:EIb

[884] Daniel Thalmann and Nadia Magnenat Thalmann. Editorial issue 31.2. Computer Animation and Virtual Worlds, 31(2):e1925:1–e1925:??, March– April 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hassaballah:2020:IFF

[885] M. Hassaballah, Abdelraheem M. Aly, and A. Abdelnaim. Interactive fluid flow simulation in computer graphics using incompressible smoothed particle hydrodynamics. Computer Animation and Virtual Worlds, 31(2):e1916:1– e1916:??, March–April 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lin:2020:EBI

[886] Xianxuan Lin, Xun Wang, Frederick W. B. Li, Jinyu Li, Bailin Yang, Kaili Zhang, and Tianxiang Wei. Examplebased image recoloring in an indoor environment. Computer Animation and Virtual Worlds, 31(2):e1917:1–e1917:??, March–April 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cui:2020:SBS

[887] Xiaoyu Cui, Ruifan Cai, Xiangjun Tang, Zhigang Deng, and Xiaogang Jin. Sketch-based shape-constrained fireworks simulation in head-mounted virtual reality. Computer Animation and Virtual Worlds, 31(2):e1920:1– e1920:??, March–April 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2020:IIc

[888] Anonymous. Issue information. Computer Animation and Virtual Worlds, 31 (3):e1907:1–e1907:??, May–June 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2020:EIc

[889] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 31.3. Computer Animation and Virtual Worlds, 31 (3):e1930:1–e1930:??, May–June 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Muni:2020:IGW

[890] Manoj K. Muni, Dayal R. Parhi, and Priyadarshi Biplab Kumar. Implementation of grey wolf optimization controller for multiple humanoid navigation. Computer Animation and Virtual Worlds, 31 (3):e1919:1–e1919:??, May–June 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2020:NAR

[891] Lianyao Wu, Wanggen Wan, Xiaoqing Yu, Chunkai Ye, and A A M Muzahid. A novel augmented reality framework based on monocular semi-dense simultaneous localization and mapping. Computer Animation and Virtual Worlds, 31 (3):e1922:1–e1922:??, May–June 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2020:ICP

[892] Jongmin Kim, Yeongho Seol, Hoemin Kim, and Taesoo Kwon. Interactive character posing with efficient collision handling. Computer Animation and Virtual Worlds, 31(3):e1923:1– e1923:??, May–June 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiang:2020:CI

[893] Wei Xiang, Xinran Yao, He Wang, and Xiaogang Jin. Cover image. Computer Animation and Virtual Worlds, 31(4-5), July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2020:CI

[894] Yu Zhang, Zhongke Wu, and Xingce Wang. Cover image. Computer Animation and Virtual Worlds, 31(4-5), July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2020:E

[895] Jian Jun Zhang, Nadia Magnenat Thalmann, Daniel Thalmann, Xiaosong Yang, Weiwei Xu, Jian Chang, and Feng Tian. Editorial. Computer Animation and Virtual Worlds, 31(4-5): e1968:1–e1968:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiang:2020:FDD

[896] Wei Xiang, Xinran Yao, He Wang, and Xiaogang Jin. FASTSWARM: a datadriven framework for real-time flying insect swarm simulation. Computer Animation and Virtual Worlds, 31(4-5): e1957:1–e1957:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2020:GSG

[897] Chien-Yuan Chen, Sai-Keung Wong, and Wen-Yun Liu. Generation of small groups with rich behaviors from natural language interface. Computer Animation and Virtual Worlds, 31(4-5): e1960:1–e1960:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2020:NLI

[898] Wen-Yun Liu, Sai-Keung Wong, and Chien-Yuan Chen. A natural language interface with casual users for crowd animation. Computer Animation and Virtual Worlds, 31(4-5):e1965:1–e1965:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2020:FPC

[899] Jie Liao, Yanping Fu, Qingan Yan, and Chunxia Xiao. Folding patch correspondence for multiview stereo. Computer Animation and Virtual Worlds, 31(4-5): e1938:1–e1938:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2020:OSV

[900] Jing Zhao, Wenjing Yang, Mingliang Yang, Wanrong Huang, Qiong Yang, and Hongguang Zhang. One-shot videobased person re-identification with variance subsampling algorithm. Computer Animation and Virtual Worlds, 31(4-5): e1964:1–e1964:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2020:TOS

[901] Jie Liao, Yanping Fu, Qingan Yan, and Chunxia Xiao. Transparent object segmentation from casually captured videos. Computer Animation and Virtual Worlds, 31(4-5):e1950:1–e1950:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2020:HFS

[902] Zhangmeng Chen, Junjun Pan, Xiaosong Yang, and Hong Qin. Hybrid features for skeleton-based action recognition based on network fusion. Computer Animation and Virtual Worlds, 31(4-5): e1952:1–e1952:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhao:2020:RDS

[903] Shimin Zhao, Miaomiao Chen, Pengjie Wang, Ying Cao, Pingping Zhang, and Xin Yang. RGB-D salient object detection via deep fusion of semantics and details. Computer Animation and Virtual Worlds, 31(4-5):e1954:1–e1954:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tian:2020:DSM

[904] Yu Tian, Wenjing Yang, Qingsong Liu, and Qiong Yang. Deep supervised multimodal semantic autoencoder for cross-modal retrieval. Computer Animation and Virtual Worlds, 31(4-5): e1962:1–e1962:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Koilias:2020:FMC

[905] Alexandros Koilias, Christos Mousas, and Christos-Nikolaos Anagnostopoulos. I feel a moving crowd surrounds me: Exploring tactile feedback during immersive walking in a virtual crowd. Computer Animation and Virtual Worlds, 31(4-5):e1963:1–e1963:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Krogmeier:2020:EFE

[906] Claudia Krogmeier and Christos Mousas. Eye fixations and electrodermal activity during low-budget virtual reality embodiment. Computer Animation and Virtual Worlds, 31(4-5):e1941:1– e1941:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ali:2020:ATG

[907] Ghazanfar Ali, Myungho Lee, and Jae-In Hwang. Automatic text-to-gesture rule generation for embodied conversational agents. Computer Animation and Virtual Worlds, 31(4-5): e1944:1–e1944:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ertugrul:2020:EMO

[908] Egemen Ertugrul, Han Zhang, Fang Zhu, Ping Lu, Ping Li, Bin Sheng, and Enhua Wu. Embedding 3D models in offline physical environments. Computer Animation and Virtual Worlds, 31(4-5): e1959:1–e1959:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yu:2020:RTS

[909] Peng Yu, Junjun Pan, Hong Qin, Aimin Hao, and Haipeng Wang. Real-time suturing simulation for virtual reality medical training. Computer Animation and Virtual Worlds, 31(4-5): e1940:1–e1940:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Feng:2020:RHO

[910] Qi Feng, Hubert P. H. Shum, and Shigeo Morishima. Resolving hand-object occlusion for mixed reality with joint deep learning and model optimization. Computer Animation and Virtual Worlds, 31 (4-5):e1956:1–e1956:??, July 2020. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tauscher:2020:ENP

[911] Jan-Philipp Tauscher, Alexandra Witt, Sebastian Bosse, Fabian W. Schottky, Steve Grogorick, Susana Castillo, and Marcus Magnor. Exploring neural and peripheral physiological correlates of simulator sickness. Computer Animation and Virtual Worlds, 31(4-5): e1953:1–e1953:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2020:ATF

[912] Jian Zhu, Silong Li, Ruichu Cai, Zhifeng Hao, Guoheng Huang, Bin Sheng, and Enhua Wu. Animating turbulent fluid with a robust and efficient high-order advection method. Computer Animation and Virtual Worlds, 31(4-5): e1951:1–e1951:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2020:BMC

[913] Lijie Yang, Tianchen Xu, Jixiang Du, Hongbo Zhang, and Enhua Wu. Brushwork master: Chinese ink painting synthesis for animating brushwork process. Computer Animation and Virtual Worlds, 31(4-5):e1949:1–e1949:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2020:VUN

[914] Seong Uk Kim, Hanyoung Jang, and Jongmin Kim. A variational U-Net for motion retargeting. Computer Animation and Virtual Worlds, 31(4-5): e1947:1–e1947:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Fan:2020:FEA

[915] Yachun Fan, Feng Tian, Xiaohui Tan, and Housen Cheng. Facial expression animation through action units transfer in latent space. Computer Animation and Virtual Worlds, 31(4-5): e1946:1–e1946:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Li:2020:DCG

[916] Yanran Li, Lingteng Qiu, Li Wang, Fangde Liu, Zhao Wang, Sebastian Iulian Poiana, Xiaosong Yang, and Jianjun Zhang. Densely connected GCN model for motion prediction. Computer Animation and Virtual Worlds, 31(4-5): e1958:1–e1958:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2020:DDB

[917] Yu Zhang, Zhongke Wu, and Xingce Wang. Dynamic disk B-spline curves. Computer Animation and Virtual Worlds, 31(4-5):e1955:1–e1955:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2020:EBS

[918] Xinyue Liu, Xingce Wang, and Zhongke Wu. Extending B-spline by piecewise polynomial. Computer Animation and Virtual Worlds, 31(4-5):e1942:1– e1942:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2020:PCS

[919] Zhaoxuan Zhang, Kun Li, Xuefeng Yin, Xinglin Piao, Yuxin Wang, Xin Yang, and Baocai Yin. Point cloud semantic scene segmentation based on coordinate convolution. Computer Animation and Virtual Worlds, 31(4-5): e1948:1–e1948:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2020:BRF

[920] Dan Zhang, Chenlei Lv, Na Liu, Zhongke Wu, and Xingce Wang. Book review: 3D face modeling from single image based on discrete shape space. Computer Animation and Virtual Worlds, 31(4-5):e1943:1–e1943:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tseng:2020:TVD

[921] Yi-Chun Tseng and Sai-Keung Wong.

Transferring video to dynamic relief structures with water interaction. Computer Animation and Virtual Worlds, 31 (4-5):e1961:1–e1961:??, July 2020. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xiang:2020:SBM

[922] Nan Xiang, Ruibin Wang, Tao Jiang, Li Wang, Yanran Li, Xiaosong Yang, and Jianjun Zhang. Sketch-based modeling with a differentiable renderer. Computer Animation and Virtual Worlds, 31 (4-5):e1939:1–e1939:??, July 2020. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ye:2020:HCF

[923] Zehao Ye, Guiqing Li, Biyuan Yao, and Chuhua Xian. HAO-CNN: Filamentaware hair reconstruction based on volumetric vector fields. Computer Animation and Virtual Worlds, 31(4-5): e1945:1–e1945:??, July 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2020:EId

[924] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial Issue 31.6. Computer Animation and Virtual Worlds, 31 (6):e1982:1–e1982:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Netek:2020:CTV

[925] Rostislav Netek, Tomas Burian, and Martin Macecek. From 360^{∘} camera toward to virtual map app: Designing low-cost pilot study. Computer Animation and Virtual Worlds, 31(6): e1924:1–e1924:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Frerichs:2020:CGS

[926] Dhana Frerichs, Andrew Vidler, and Christos Gatzidis. Computer graphics simulation of natural mummification by desiccation. Computer Animation and Virtual Worlds, 31(6):e1927:1–e1927:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Koilias:2020:IWV

[927] Alexandros Koilias, Michael G. Nelson, Christos-Nikolaos Anagnostopoulos, and Christos Mousas. Immersive walking in a virtual crowd: the effects of the density, speed, and direction of a virtual crowd on human movement behavior. Computer Animation and Virtual Worlds, 31(6):e1928:1–e1928:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Magnoux:2020:RTV

[928] Vincent Magnoux and Benoît Ozell.

Real-time visual and physical cutting of a meshless model deformed on a background grid. Computer Animation and Virtual Worlds, 31(6):e1929:1–e1929:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2020:TDC

[929] Zili Zhang, Yunfei Li, Bailin Yang, Frederick W. B. Li, and Xiaohui Liang. Target-driven cloud evolution using position-based fluids. Computer Animation and Virtual Worlds, 31(6): e1937:1–e1937:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Shi:2020:AHS

[930] Jiajun Shi, Chen Li, Changbo Wang, Hong Qin, and Gaoqi He. An advanced hybrid smoothed particle hydrodynamics-fluid implicit particle method on adaptive grid for condensation simulation. Computer Animation and Virtual Worlds, 31(6):e1967:1– e1967:??, November–December 2020. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2021:IIa

[931] Anonymous. Issue information. Computer Animation and Virtual Worlds, 32 (1):e1931:1–e1931:??, January–February 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2021:EIa

[932] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 32.1. Computer Animation and Virtual Worlds, 32 (1):e1991:1–e1991:??, January–February 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2021:AEE

[933] Shiguang Liu, Si Gao, and Siqi Xu. Animating explosion with exploding sound and rigid-body sound. Computer Animation and Virtual Worlds, 32 (1):e1970:1–e1970:??, January–February 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhu:2021:CVL

[934] Jian Zhu, Silong Li, Ruichu Cai, Zhifeng Hao, Guoheng Huang, Bin Sheng, and Enhua Wu. Compensating the vorticity loss during advection with an adaptive vorticity confinement force. Computer Animation and Virtual Worlds, 32 (1):e1973:1–e1973:??, January–February 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Han:2021:SFM

[935] Yi Han, Qianwen Chao, and Xiaogang Jin. A simplified force model for mixed traffic simulation. Computer Animation and Virtual Worlds, 32(1):e1974:1– e1974:??, January–February 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Guo:2021:PVA

[936] Hong Guo, Shanchen Zou, Chuying Lai, and Hongxin Zhang. PhyCoVIS: a visual analytic tool of physical coordination for cheer and dance training. Computer Animation and Virtual Worlds, 32 (1):e1975:1–e1975:??, January–February 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xu:2021:KEH

[937] Chenxu Xu, Wenjie Yu, Yanran Li, Xuequan Lu, Meili Wang, and Xiaosong Yang. KeyFrame extraction for human motion capture data via multiple binomial fitting. Computer Animation and Virtual Worlds, 32(1):e1976:1– e1976:??, January–February 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2021:LWP

[938] Qiang Chen, Guoliang Luo, Yang Tong, Xiaogang Jin, and Zhigang Deng. A linear wave propagation-based simulation model for dense and polarized crowds. Computer Animation and Virtual Worlds, 32(1):e1977:1–e1977:??, January–February 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2021:EIb

[939] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 32.2. Computer Animation and Virtual Worlds, 32(2):e1994:1–e1994:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yang:2021:RBA

[940] Dong Yang, Wenjing Yang, Minglong Li, and Qiong Yang. Role-based attention in deep reinforcement learning for games. Computer Animation and Virtual Worlds, 32(2):e1978:1– e1978:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liao:2021:DMS

[941] Jie Liao, Mengqiang Wei, Yanping Fu, Qingan Yan, and Chunxia Xiao. Dense multiview stereo based on image texture enhancement. Computer Animation and Virtual Worlds, 32(2):e1979:1– e1979:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bruno:2021:THM

[942] Alessandro Bruno, Morgan Moore, Jinglu Zhang, Stéphane Lancette, Ville P. Ward, and Jian Chang. Toward a head movement-based system for multilayer digital content exploration. Computer Animation and Virtual Worlds, 32(2):e1980:1–e1980:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Grabowski:2021:UVR

[943] Andrzej Grabowski and Kamil Jach. The use of virtual reality in the training of professionals: with the example of firefighters. Computer Animation and Virtual Worlds, 32(2):e1981:1– e1981:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Csi:2021:ERV

[944] Benjamin S. Botha M. Sc Csi, Lizette de Wet Ph.D, and Yvonne Botma Ph.D. Experts’ review of a virtual environment for virtual clinical simulation in South Africa. Computer Animation and Virtual Worlds, 32(2):e1983:1– e1983:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2021:SLS

[945] Jong-Hyun Kim and Jung Lee. Synthesizing large-scale fluid simulations with surface and wave foams via sharp wave pattern and cloudy foam. Computer Animation and Virtual Worlds, 32(2): e1984:1–e1984:??, March–April 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2021:IIb

[946] Anonymous. Issue information. Computer Animation and Virtual Worlds, 32 (3-4):e2027:1–e2027:??, June 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lee:2021:EI

[947] Won Sook Lee, Chris Joslin, Daniel Thalmann, Eric Paquette, and Sung-Hee Lee. Editorial issue 32.3. Computer Animation and Virtual Worlds, 32(3-4): e2030:1–e2030:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2021:GMA

[948] Shao-Chieh Chen, Guan-Ting Liu, and Sai-Keung Wong. Generation of multiagent animation for object transportation using deep reinforcement learning and blend-trees. Computer Animation and Virtual Worlds, 32(3-4): e2017:1–e2017:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cui:2021:TUE

[949] Dixuan Cui, Dominic Kao, and Christos Mousas. Toward understanding embodied human-virtual character interaction through virtual and tactile hugging. Computer Animation and Virtual Worlds, 32(3-4):e2009:1–e2009:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mazumdar:2021:SAV

[950] Angshuman Mazumdar and Christos Mousas. Synthesizing affective virtual reality multicharacter experiences. Computer Animation and Virtual Worlds, 32(3-4):e2004:1–e2004:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cui:2021:IVI

[951] Dixuan Cui, David Whittinghill, Atsushi Fukada, Christos Mousas, and Nicoletta Adamo. Interacting with virtual instructors: the effect of gender and years of study on the perception of in-game instructors. Computer Animation and Virtual Worlds, 32(3-4): e2026:1–e2026:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Sun:2021:SVP

[952] Chao Sun, Srinivasan Ramachandran, Eric Paquette, and Won-Sook Lee. Single-view procedural braided hair modeling through braid unit identification. Computer Animation and Virtual Worlds, 32(3-4):e2007:1–e2007:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2021:FFE

[953] Fangyu Wu, Chaoyi Pang, and Bailing Zhang. FaceCaps for facial expression recognition. Computer Animation and Virtual Worlds, 32(3-4): e2021:1–e2021:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Lei:2021:IRB

[954] Songze Lei, Baihua Dong, Yonggang Li, Feng Xiao, and Feng Tian. Iris recognition based on few-shot learning. Computer Animation and Virtual Worlds, 32 (3-4):e2018:1–e2018:??, June 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Mao:2021:BRL

[955] Jiahui Mao, Tingting Li, Feiyu Zhang, Meili Wang, Jian Chang, and Xuequan Lu. Bas-relief layout arrangement via automatic method optimization. Computer Animation and Virtual Worlds, 32 (3-4):e2012:1–e2012:??, June 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Hu:2021:UMC

[956] Xiaoyan Hu, Xizhao Bao, Shunbo Xie, and Guoli Wei. Unsupervised motion capture data segmentation based on topic model. Computer Animation and Virtual Worlds, 32(3-4): e2005:1–e2005:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2021:LDM

[957] Chi Zhou, Zhangjiong Lai, Suzhen Wang, Lincheng Li, Xiaohan Sun, and Yu Ding. Learning a deep motion interpolation network for human skeleton animations. Computer Animation and Virtual Worlds, 32(3-4):e2003:1–e2003:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Victor:2021:LBP

[958] Léon Victor, Alexandre Meyer, and Saïda Bouakaz. Learning-based pose edition for efficient and interactive design. Computer Animation and Virtual Worlds, 32(3-4):e2013:1–e2013:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kim:2021:IMC

[959] Jongmin Kim, Yeongho Seol, and Taesoo Kwon. Interactive multi-character motion retargeting. Computer Animation and Virtual Worlds, 32(3-4): e2015:1–e2015:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Aman:2021:MLT

[960] Aytek Aman, Serkan Demirci, Ugur Güdükbay, and Ingo Wald. Multi-level tetrahedralization-based accelerator for ray-tracing animated scenes. Computer Animation and Virtual Worlds, 32(3-4): e2024:1–e2024:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Deng:2021:RBF

[961] Nianchen Deng, Zhenyi He, and Xubo Yang. Render-based factorization for additive light field display. Computer Animation and Virtual Worlds, 32(3-4): e2010:1–e2010:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2021:IDF

[962] Maolin Wu, Shiguang Liu, and Qing Xu. Improved divergence-free smoothed particle hydrodynamics via priority of divergence-free solver and SOR. Computer Animation and Virtual Worlds, 32 (3-4):e2006:1–e2006:??, June 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2021:PRV

[963] Qiaorui Chen, Shuai Zhang, and Yao Zheng. Parallel realistic visualization of particle-based fluid. Computer Animation and Virtual Worlds, 32(3-4): e2019:1–e2019:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wu:2021:DBA

[964] Huisi Wu, Yifan Li, Le Chen, Xueting Liu, and Ping Li. Deep boundary-aware semantic image segmentation. Computer Animation and Virtual Worlds, 32(3-4): e2023:1–e2023:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zha:2021:AAM

[965] Hengfeng Zha, Rui Liu, Xin Yang, Dongsheng Zhou, Qiang Zhang, and Xiaopeng Wei. ASFNet: Adaptive multiscale segmentation fusion network for real-time semantic segmentation. Computer Animation and Virtual Worlds, 32(3-4): e2022:1–e2022:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Xu:2021:SSG

[966] Zezi Xu and Shuangjiu Xiao. Spatial semantic graph enhanced monocular SLAM System. Computer Animation and Virtual Worlds, 32(3-4): e2025:1–e2025:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2021:ADA

[967] Yu Zhou, Zhihua Chen, Bin Sheng, Ping Li, Jinman Kim, and Enhua Wu. AFFdehazing: Attention-based feature fusion network for low-light image dehazing. Computer Animation and Virtual Worlds, 32(3-4):e2011:1–e2011:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bai:2021:EEE

[968] Zechen Bai, Naiming Yao, Nidhi Mishra, Hui Chen, Hongan Wang, and Nadia Magnenat Thalmann. Enhancing emotional experience by building emotional virtual characters in VR volleyball games. Computer Animation and Virtual Worlds, 32(3-4):e2008:1–e2008:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ferstl:2021:EEG

[969] Ylva Ferstl, Michael Neff, and Rachel McDonnell. ExpressGesture: Expressive gesture generation from speech through database matching. Computer Animation and Virtual Worlds, 32(3-4): e2016:1–e2016:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Noah:2021:EEA

[970] Naheem Noah and Sanchari Das. Exploring evolution of augmented and virtual reality education space in 2020 through systematic literature review. Computer Animation and Virtual Worlds, 32(3-4):e2020:1–e2020:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Liu:2021:BIM

[971] Xiaojun Liu, Changyan He, Hantao Zhao, Jinyuan Jia, and Chang Liu. Building information modeling indoor path planning: a lightweight approach for complex BIM building. Computer Animation and Virtual Worlds, 32(3-4): e2014:1–e2014:??, June 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2021:IIc

[972] Anonymous. Issue information. Computer Animation and Virtual Worlds, 32(5):e1935:1–e1935:??, September– October 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2021:EIc

[973] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 32.5. Computer Animation and Virtual Worlds, 32(5):e2034:1–e2034:??, September– October 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Wang:2021:RCD

[974] Monan Wang and Jiaqi Cao. A review of collision detection for deformable objects. Computer Animation and Virtual Worlds, 32(5):e1987:1– e1987:??, September–October 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Kumar:2021:HNF

[975] Priyadarshi B. Kumar, Dayal R. Parhi, and Mukesh Sethy. Humanoid navigation: a firefly based approach. Computer Animation and Virtual Worlds, 32(5): e1969:1–e1969:??, September–October 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Tastan:2021:HBR

[976] Oguzhan Tastan and Yusuf Sahillioglu.

Human body reconstruction from limited number of points. Computer Animation and Virtual Worlds, 32(5): e1995:1–e1995:??, September–October 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yoon:2021:ILC

[977] Jong-Chul Yoon and HyeongYeop Kang. Interactive learning in the classroom: a mobile augmented reality assistance application for learning. Computer Animation and Virtual Worlds, 32(5):e1989:1– e1989:??, September–October 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Cho:2021:XPA

[978] Yunsik Cho, Jiewon Kang, Jaekyung Jeon, Jongchan Park, Mingyu Kim, and Jinmo Kim. X-person asymmetric interaction in virtual and augmented realities. Computer Animation and Virtual Worlds, 32(5):e1985:1– e1985:??, September–October 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhuo:2021:MCE

[979] Lin Zhuo, Zhen Liu, Tingting Liu, Chih-Chieh Hung, and Yanjie Chai. Modeling crowd emotion from emergent event video. Computer Animation and Virtual Worlds, 32(5):e1988:1– e1988:??, September–October 2021. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Chen:2021:HPT

[980] Zhiyong Chen, Xiaowei Chen, Yong Ma, Shihui Guo, Yipeng Qin, and Minghong Liao. Human posture tracking with flexible sensors for motion recognition. Computer Animation and Virtual Worlds, 32(5):e1993:1–e1993:??, September–October 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2021:IId

[981] Anonymous. Issue information. Computer Animation and Virtual Worlds, 32 (6):e1936:1–e1936:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2021:EId

[982] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 32.6. Computer Animation and Virtual Worlds, 32 (6):e2037:1–e2037:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Qian:2021:SRA

[983] Kun Qian, Meili Wang, and Yaqing Cui. Simulation ready anatomy model generation pipeline for virtual surgery. Computer Animation and Virtual Worlds, 32(6):e1986:1–e1986:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Usman:2021:SSA

[984] Muhammad Usman, Brandon Haworth, Petros Faloutsos, and Mubbasir Kapadia. Simulation-as-a-service: Analyzing crowd movements in virtual environments. Computer Animation and Virtual Worlds, 32(6):e1990:1–e1990:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Bai:2021:EHF

[985] Junxuan Bai, Rong Dai, Ju Dai, and Junjun Pan. EmoDescriptor: a hybrid feature for emotional classification in dance movements. Computer Animation and Virtual Worlds, 32(6): e1996:1–e1996:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Ghafourzadeh:2021:LCE

[986] Donya Ghafourzadeh, Sahel Fallahdoust, Cyrus Rahgoshay, Andre Beauchamp, Adeline Aubame, Tiberiu Popa, and Eric Paquette. Local control editing paradigms for part-based 3D face morphable models. Computer Animation and Virtual Worlds, 32(6): e2028:1–e2028:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

GomezJauregui:2021:DEP

[987] David Antonio Gómez Jáuregui, Tom Giraud, Brice Isableu, and Jean-Claude Martin. Design and evaluation of postural interactions between users and a listening virtual agent during a simulated job interview. Computer Animation and Virtual Worlds, 32(6):e2029:1–e2029:??, November–December 2021. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Anonymous:2022:IIa

[988] Anonymous. Issue information. Computer Animation and Virtual Worlds, 33 (1):e1997:1–e1997:??, January–February 2022. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Thalmann:2022:EIa

[989] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial issue 33.1. Computer Animation and Virtual Worlds, 33 (1):e2042:1–e2042:??, January–February 2022. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Yan:2022:AVA

[990] Yuling Yan, Lijun Zhang, and Minye Chen. AGRMTS: a virtual aircraft maintenance training system using gesture recognition based on PSO-BPNN model. Computer Animation and Virtual Worlds, 33(1):e2031:1– e2031:??, January–February 2022. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2022:UGU

[991] Junsong Zhang, Shaoqiang Zhu, Kunxiang Liu, and Xiaoyu Liu. UGSC-GAN: User-guided sketch colorization with deep convolution generative adversarial networks. Computer Animation and Virtual Worlds, 33(1):e2032:1– e2032:??, January–February 2022. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Messaci:2022:ZFE

[992] Assia Messaci, Nadia Zenati, Mahmoud Belhocine, and Samir Otmane. Zoom-fwd: Efficient technique for 3D gestual interaction with distant and occluded objects in virtual reality. Computer Animation and Virtual Worlds, 33 (1):e2033:1–e2033:??, January–February 2022. CODEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhou:2022:MEF

[993] Wen Zhou, Wenying Jiang, Biao Jie, and Weixin Bian. Multiagent evacuation framework for a virtual fire emergency scenario based on generative adversarial imitation learning. Computer Animation and Virtual Worlds, 33(1):e2035:1– e2035:??, January–February 2022. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).

Zhang:2022:SIL

[994] Jingjing Zhang, Jingsheng Lei, Shengying Yang, and Xinqi Yang. Semantic interaction learning for fine-grained vehicle recognition. Computer Animation and Virtual Worlds, 33(1):e2036:1– e2036:??, January–February 2022. CO-DEN ???? ISSN 1546-4261 (print), 1546-427X (electronic).